A Complete Bibliography of Publications in Marine Policy (2000–2009)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

14 October 2023 Version 1.04

Title word cross-reference

111-198 [Ano06j]. 143-230 [Ano02d]. 173-250 [Ano01c]. 179-274 [Ano00c]. 17th [Dro04]. 183-434 [Ano09h]. 189-282 [Ano04i]. 189-290 [Ano05h]. 193-288 [Ano03i]. 199-298 [Ano06k].

2000 [PFS09]. 2006 [DE07]. 21st [Li03]. 229-358 [Ano07h]. 231-314 [Ano02e]. 24 [ABF00b]. 251-322 [Ano01d]. 255-534 [Ano08h]. 27 [LFL05a, SR05]. 275-360 [Ano00d]. 283-360 [Ano04j]. 291-384 [Ano05i]. 299-454 [Ano06l].

315-402 [Ano02f]. 323-388 [Ano01e]. 357-448 [Ano03j]. 359-584 [Ano07i]. 361-440 [Ano00e]. 361-450 [Ano04k]. 385-480 [Ano05j].

403-492 [Ano02g]. 435-536 [Ano09i]. 441-502 [Ano00f]. 449-556 [Ano03k]. 451-556 [Ano04l]. 455-604 [Ano06m]. 481-538 [Ano05k].

535-758 [Ano08i]. 537-736 [Ano09j]. 544 [SR05].

605-842 [Ano06n]. 645-746 [Ano07j].

737-868 [Ano09k]. 75-142 [Ano02h].

8 [DE07]. 81-228 [Ano07k]. 83-178 [Ano00g]. 845-1102 [Ano08j].

90 [HS08]. 91-172 [Ano01f]. 97-188 [Ano04m]. 97-192 [Ano03l].

abate [GDM06]. Aboriginal [McG03b]. absence [Fah08]. abstention [Tsu04]. abundance [Pol06]. acceptability [WP09]. Acceptable [Utn08a]. acceptance [KS09]. access [ANL09, BBB01, BT04, Elv09, FMPPF08, Gry03, Pet03]. accident [Cho07]. accidents [BBNB08, TK09]. Accord [FE09]. account [AS07, Bro05, Høn00]. accountability [AL02]. accumulation [SA02]. Achieving [NBR09]. ACIA [Sch07a]. acknowledgements [VA05a]. Acknowledgments [VA04]. across [CLyB06, MJ08]. act [KE03, FE09, Kim02]. Action [Fra08, Jon06]. active [GK05]. activities [DYB05, Hay05, Val05a, Ver07]. activity [OGC^{+}09, Ska05]. actors [dVMdC08]. adapting [Day08, Syd02, KTH^{+}07]. Adaptive [RSS^{+}07, GK05, MMR08, WASK06]. address [Bor01]. Addressing [Var08]. administration [SRW06]. adoption [CS07, Hal03, HH04]. advancing [CROP09, Dou08]. affect [SU07]. affecting [PKH01, Ser04]. Africa [Ano06g, BC06, Cla06, Har09, HS01, HK06a, HH00, KF02, NDV^{+}04, PS07, Sow06, vS02, vSBHS06, vSH06]. African [CvSP06, NH06, Wit06]. after [GGVL09, Gui04, Ho09b, JM00, Kim02, Sch05]. agency [DJ01, MB09]. agreement [Lev08]. agreements [Elv09, HJ02, KF02, Kim03b, Wit08]. aid [Pet03, SRVLGG02]. aims [Ano02a, Ano03d, Ano03e, Ano03f, Ano04b, Ano04a, Ano04c, Ano04d, Ano04e, Ano05a, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano06a, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano07a, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f]. air [LL06]. Al [AOPM00]. Al-Batinah [AOPM00]. Alaskan [HG01]. Albemarle [PBSS06]. Aleutian [Fin05]. algae [vdBND^{+}02]. alien [BWA^{+}03]. allocation [HK06b, SA02]. alternate [LS00]. alternative [BA05, Hea04, JHMC07, SA08]. America [GDI^{+}09, KTH^{+}07]. American [KTH^{+}07]. americanus [Lev09b]. among [BBNB08, HRP08]. Ampat [BRS08]. Analysis [BK09, TK09, EOU09, Fle01, Gez08, GM01a, Gui04, HS01, HCL^{+}09, KF07, KF08, KKR08, KYC05, LC06, LOS06, MPT09, MCO08, Mor06, RSJ^{+}03, SBF^{+}04, TRvAB05, Utn08b, vdBND^{+}02]. analytical [Nie03]. Anchovetta [Pon01]. anchovy [BRS08]. Andalusia [PSFE07, PNN08, dC08]. Angling [SJ03]. animal [KA06]. Antarctic [Gre01, Jab06]. ante [GM09]. anthropological [OJ08]. any [Sch07b]. apartheid [vSBHS06, vSH06]. APEC [Ho09b]. application [CGS08, HNMB07, Moe08, OGC^{+}09]. applications [Wes03]. applied [KTH^{+}07]. Appraising [CS07]. approach [AMS05, AE01, BA05, GH08, HY03, JBP09, JHMC07, KPN^{+}09, MFM09, MB09, Mor08, Par00, Pol02a, PS01, Ska04, Sta05, Sta08, SM06, Var08, WW05, AH06]. Approaches [TRF08, BM08, BWR08, BET^{+}07, Mur07, OB09]. Aquaculture [Han03, SLT08, AJ09, Bes06, EOU09, Gib09a, Hea08, HR04, KA06, Mar01, MN09, NSVW01, PPB06, WCB06, WP09]. architectures [Hin01]. Arctic [Hun09, Sta03, Sto07]. area [Jon08, Jon09, MMS09, NBR09, Ped08, QWJA09, SBF^{+}04, RTWJ05]. areas [BCM09a, Cli03, CCCD07, Fab08, Foc08, Ove06, RGB09, SR08, SBB^{+}09]. armed [Ho09a]. arrangement [KS04]. arrangements [Bee07, Gry03, LS00, WKT00, YM06]. arrest [Som01]. arrivals [ZYN09]. Artificial [KK08, Kai06]. Artisanal [PSFE07, PNN08, SR08, CVL02, FGA00, GGAD08, JGU09, WPBD03, dC08]. artisinal [AOPM00]. Asia [Ho09a, Kim03b, NDV^{+}04, PKH01, PPP^{+}07, WW06, WASK06]. Asian [Ber02, SGP^{+}08, TS05]. Assessing [AMS05, CFW^{+}07, KPN^{+}09, LCK04, Mar08, SDWF09]. Assessment [GGSRVL06, OGC^{+}09, FFM06, GP00, HHM^{+}09, Hun09, JHW09, LW07, Mal03, MS02, OB09, Por09, TS05, WF01, WBGC08, Yan08]. assimilation [Sch07c]. asymmetric [Jen08]. asymmetrical [Ebb03]. Atlantic [AL02, Bjø09, FE09, JM00, Lev08, BA05, BM06, BR08, DJ07, DB06, FPS05, GGVLIM03, GGVL05, Han07b, HR06, Lev09a, Lev09b, McG03b, MN09, Skj06, Sto09, Syd01]. attack [Ho09b]. attention [WAKP05]. attitudes [BH09, KKP00, Pet08]. attribute [EO09]. attributes [Utn08b]. auction [NZF09]. auctioning [Tro04]. auctions [AVH05, Arm01, Kap00, VEH02]. auratus [MS02]. Australia [MBJI08, Vin06, CCCD07, DPE^{+}08, DKM08, FHCS05, Jab06, Jab09, Kea01, Let00, MG06, MD08, Sco00]. Australian [BCM09b, CBB00, DCHM07, GM09, Gre01, Jab09, PG09, PPW^{+}09, RM00]. authorities [Hew07]. available [Eks09]. aware [FLT06]. awareness [FPHP09, JP07].

back [MD08]. balancing [SHQF06]. Bali [Val04c]. Baltic [Ugg07]. ban [WPBD03]. Banggi [TS07]. Bangladesh [BNA^{+}01]. Barbados [BC07b]. Barents [Høn00]. Barrier [Day08, NBR09]. barriers [Gib09a, GP00, Har01]. based [BA05, BWR08, CROP09, CFW^{+}07, CWRR09, Cla09, CN08, DPE^{+}08, Dou08, FF06, Fle01, FPS05, Gai08, GMN^{+}07, GDI^{+}09, GL08, GH08, GT01, KKI^{+}09, LOS06, LB04, MMS09, MPT09, PGM^{+}05, PKS^{+}09, Tho08, TG09, WP09, WBCK04, dVM04, dlM05, BET^{+}07]. Basel [Moe08]. baselines [Val03]. basin [HJ02]. basis [Fle03]. bass [Lev09a]. bathing [Hew07]. Batinah [AOPM00]. Bay [BRS08, MBJI08, KKI^{+}09]. be [Bra08, FSPKR09, KF07, dJDT09]. bear [MLRK08]. behavior [Val09]. behaviour [AEG^{+}08]. behaviours [Tho08]. behind [OJ08]. Beibu [YM06].

Belgian [DMVS07, Som01]. Belgium [BCM09a]. Belize [Per09]. Benchmarking [GKMR07, vSBHS06]. benefit [GTLC02]. benefits [GM09, KF02, Ove06, Ove07, SVL07, Ehl08]. benefitted [Har09]. benthic [TRF08]. Bering [Fin05, MLRK08]. better [BNA^{+}01, Isa06]. between [BC07b, CLT08, CW08, Kan03, Kim03a, KA06, Per09, RWB08, VGH^{+}08, Yan08, vHvT09]. Beyond [Jen06, Som01, Tsu04]. Bight [BM09]. bilateral [Wit08]. binary [KP04a]. biodiversity [BWA^{+}03, Kay04, McB04, QCB09, Sub07]. biofouling [RT08]. Biogeography [CSG^{+}09]. biological [BB06, FGA00, GGVL09, vdBND^{+}02]. Biosphere [PBB08]. bird [Hun06]. black [Lev09a, dVM02a]. block [AL02]. blue [Sta06]. Bluefin [Pol02a, CBB00]. bluefins [Ser04]. board [Ano04a, Ano02a, Ano03d, Ano03e, Ano03f, Ano03a, Ano03b, Ano03c, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano05a, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano06a, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano07a, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f]. Board/aims [Ano06a, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano07a, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f]. boat [Coo05, PG09]. boats [PC00]. bolts [Sta05]. border [Per09]. Bothkennar [MM04]. Bothnia [RC01]. bottom [Ove06, Ove07]. bottomfish [GL03]. boundaries [Gre01, MR01, dVM04]. Boundary [VVG03]. boycotts [HH06]. Brancaster [MBSLP02]. Brazil [KS04, SR08, dS04]. Brazilian [AS07]. Breaking [Moe08]. bridges [WdK02]. brief [NH07, WRD06]. British [Dro04, GMN^{+}07, Hea08, MG09, SVL07]. Building [MLRK08, FD03, Pot00, SP01, Val04a, VS01, Yan03, Yan08]. bulk [RM02]. buying [Kap00, MD08]. by-products [EY07]. bycatch [BJBC09, DCNP09, GDM06, MWL^{+}09, NZF09].

cables [CSH00]. Caird [Bor01]. Calamianes [Fab08]. California [BM09, Eks09, FF06, SBF^{+}04]. call [But07]. Calming [VH09]. campaign [SM06]. campaigns [JP07]. Can [Kar01, LS08, SU07, dJDT09, KF07]. Canada [BP08, Coo05, CWP06, DM09, DS09, FHCS05, FE09, Hea08, JM00, Lou07, McG03b]. Canadian [Bod02, Dav08, ET08, HR04, MN09]. Canal [Llá05a, Lla05b]. Cancellation [MG06]. cannot [vS02]. Cantabria [PL01]. capabilities [Mal03]. Capacity [IMP09, VS01, AEG^{+}08, FD03, Jen02, Joh05, PC00, Pot00, RSJ^{+}03, SGP^{+}08, SP01, SA02, Sta06, YY08]. Cape [PBB08]. capitalism [FL01]. caps [FW08]. capture [BdKvD^{+}08, Joh05, LW07, Sub07]. carbon [KS09]. Caribbean [BNA^{+}01, HMM^{+}04, LCK04, SRMM03, SM08]. CARICOM [SRMM03]. carriers [TS05]. Case [Lev08, AJ09, AMS05, AEG^{+}08, BCM09b, BL00, CFG06, CVL02, CW00, Dav08, DW08, DM09, DS09, DTFÁIM04, DMVS07, Dro04, DP00, EE03, FLT06, FH04, FMPPF08, Fra08, FGA00, GM09, HR06, HHR00, HK06b, How01, Lax06, LFI01, LOT05, MMS09, MPB^{+}02, Mat01, MGL09, PBSS06, PL01, SR08, Som03, Sub07, TWN^{+}09, Vee02, WPBD03, dJDT09, dC08, MLRK08]. case-study [GM09]. cases [AHS00]. Castellammare [WPBD03]. Casualties [RM02]. casualty [KF07]. casuistry [CCG^{+}07]. catalogue [NH07]. Catch [SHQF06, Agn00, CBB00, Gez08, MBJI08, Pet03, Pol06, BRS08]. Catch-quota [SHQF06]. catches [Han07b, Pol02a]. Catching [WAKP05]. Category [Joh06]. cause [VES06]. caused [FFM06]. causes [CFG05, FGA00]. CBCRM [MPT09]. CBM [BW09]. CBMs [Son05]. CCAMLR [Agn00, Par00, WBGC08]. Census [San05]. central [RSJ^{+}03, BC07a, Cam00, NSH01]. Centre [Ho09a]. centred [MWV02]. Centropristis [Lev09a]. century [Dro04, Li03]. cetaceans [JHW09]. chain [MN08]. challenge [Cla06, CSH00, DJ01, Li03, Yan03]. Challenges [QWJA09, BM08, Ehl08, HK06a, JC09b, KS04, Lan08, LTX08, Mil07, RWB08, dVM02b, BBNB08, Mit08, PBB08]. Change [HH00, AJ09, Cla06, FL01, Jen04, LS00, LCK04, LB04, Mal03, Sch07a, SB05, SS07]. changes [Sta07, Syd02]. changing [BC06, GHR^{+}08, Mar01, Vee02, Pet08]. characteristics [GM01b, Mit04]. Characterization [Lev09a]. Cheating [Han07a]. check [Joh02]. Cherai [TG09]. Chile [CGS08, Sch07c]. China [Key03, hS00, Son05, Kan03, Key01, QWJA09, Xue06, YY08]. Chinese [HY03, LTX08, XX05, Xiz04]. choice [HR04, KP04a, KPN^{+}09]. choices [SA02]. circle [JM04]. cities [Li03]. Civil [Mas03, MHJ07]. claim [Val03]. claims [Bes01, Til08]. classical [vdB00]. classifying [TRF08]. classroom [Fle01]. Climate [Cla06, Mil07, SB05, HNMB07, LCK04, Sch07a, SS07]. climatic [McG07]. closed [BdKvD^{+}08]. closure [Lev09b]. clubs [Ben01]. clusters [DS09]. Co [Bee07, CSG^{+}09, DTFÁIM04, MM05, Nob03, NBR09, Sch07c, CJ07, Cli03, Ebb02, Ebb09, EE03, FMPPF08, HS01, Jen05, KS04, Kap00, KKP00, KM04, Kar01, MMS09, NDV^{+}04, Nob00, PKH01, Tho08, TG09, WASK06, Yan03, Yan08, dS04]. co-governance [FMPPF08]. co-managed [Ebb09, Tho08]. Co-management [Bee07, CSG^{+}09, DTFÁIM04, MM05, NBR09, Sch07c, CJ07, Cli03, Ebb02, EE03, HS01, Jen05, KS04, Kap00, KKP00, KM04, Kar01, MMS09, NDV^{+}04, Nob00, PKH01, TG09, WASK06, Yan03, Yan08, dS04]. Co-operating [Nob03]. Coasean [Edw08]. Coast [BJBC09, NVPR09, AvdH08, CS07]. Coastal [CC09, FJH07, FP07, HS01, JBN07, Mas08, MCO08, OB09, BL08, BL00, Car07, Che08, DSH03, DM09, Edw09, EY07, Fle00, Fle03, GJGT04, Gib09a, GA07, HBB00, JHW09, Jen07, JC09a, KKI^{+}09, Lan08, LCK04, Li03, LOT05, MM05, MPT09, Mal03, OLC^{+}09, PRH^{+}06, Sag05, SMS06, SM07, SA02, SBB^{+}09, SM06, Tau07, Tza05, VES06, Wes03, TLHN07]. coasts [Smi03]. Cochin [TG09]. cod [Han07a, HHR00, SA02, SU07, Utn07]. cod-fisheries [Utn07]. Codes [hS00]. coherent [JC09b]. coins [Dro04]. cold [AvdH08]. cold-water [AvdH08]. collaboration [PPP^{+}07, SBB^{+}09]. Collaborative [dSK06, Dav08, KdSR07, RHCS05]. collapsed [Bra08, FSPKR09]. collapsing [JP07]. collection [Bal04]. Collective [Jon06, KE03]. collectors [FMPPF08]. Columbia [Hea08, SVL07]. combat [Sto09]. Combating [Ho09a]. combined [KPN^{+}09]. coming [Hun09]. Comment [Val00, FSPKR09, Haa00a]. Comments [LFL05a, LFL05b, SR05]. Commercial [HH06, AMW09, Bod02, EE03, Har07, HB07a, Kea01, LVH^{+}09, Lev09a, PLB04, Pon01, SS07, WJPN00]. Commission [Bjø09, Lev08, Hea04, Hea06, Ili08d]. commited [AL02]. commitment [SM08]. Common [DG05, HG01, MG09, HRP08, Jen02, RHKS09, SC04, dS04, FA06, GH03]. commons [McB04]. commonwealth [GM09, MG06, PPW^{+}09]. Commonwealth-managed [MG06]. communication [GDM06]. communities [Che03, FL01, PRH^{+}06, SHA08]. community [CWRR09, FÓ08, FFM06, How01, Jen00a, LOS06, LB04, MPT09, PS07, Tho08, TG09, WBCK04, WCKB09, Hat00]. community-based [CWRR09, LOS06, LB04, MPT09, Tho08, TG09, WBCK04]. companies [FLT06]. comparative [Fle01, Gez08, LOS06]. Comparing [MBJI08, vS02]. comparison [Kim03a, LDBB08, WNB^{+}08]. compensation [Kim03a, Mas03]. competition [HRP08]. Competitive [PLB04]. competitiveness [Bes06, CMCNL^{+}09]. complex [MMR08]. Compliance [Høn00, LL06, SW08, BCM09b, EE03, HK06a, Hau08, Nie03, NM03, Gez06]. complicated [GH08]. Compromise [Ili08a]. concept [Det06]. concepts [Gai08]. conceptual [MR01, RWB09]. Concern [Hun06]. concerning [Mur07]. concerns [MCK08]. Conclusions [Ehl08, Val04a, Val05b]. condemned [vS02]. Conditions [PKH01, PNN08]. conduct [hS00]. Cone [GDI^{+}09]. Conflict [BC07b, BL00, Bee07, BNA^{+}01, BL08, FTF00, Kea01, PPP^{+}07]. conflicts [BR07, Hil07, Maj07, RWB08]. conformity [Høn00]. confusion [CCG^{+}07]. Congressionally [Ber08]. consciousness [Sch07c]. consensus [Eyt00, FTF00, Val05a]. consequences [CFG05, Gib07, RHKS09]. Conservation [Lev08, AL02, BHJV09, BET^{+}07, DJ07, FMPPF08, Hun09, JBP09, Kay04, Lev09b, Ove07, RWB08, vZvDT02, CCG^{+}07]. considerations [Kay04]. Considering [FF06]. consolidation [Eyt00]. Consortium [HR06]. constrained [Ped08]. constraints [Hol00]. Constructs [GJGT04]. consulting [Mah06]. containerships [WF01]. contested [NBR09]. Context [LOS06, Foc08, Gai08, Mit08, KTH^{+}07]. contexts [CFW^{+}07, DS09]. continental [Jab06, Jab09, Key01]. continued [Joh05]. continuing [WDAT00]. contracts [Edw08, MN08]. Contractual [SMS06]. contribution [Dev00, Hea08]. control [CMW09, EE03, Erc06, KF07, McG03a, RT08, RNG06, dlM05]. controls [MBJI08]. controversies [BH09, Sta03]. Convention [Ili08c, Kan03, Kim03b, Moe08, Sto07, Ver07]. conventional [RNG06, MBJI08]. conventions [KF09]. conveyance [BM09]. cooperation [HRP08]. Cooperative [KM04, BA05, HR06]. Cooperatives [KE03]. coordinated [CWP06]. Coping [PRH^{+}06]. coral [AvdH08, JGU09, PPB06]. core [PS01, Pol02b]. COREPOINT [CC09]. Cork [OGC^{+}09]. corporate [FLT06, GL07]. corruption [HT09]. cost [CMCNL^{+}09, KM09, PBSS06].

Costa [Alp06]. Costs [AHS00, GM09, MM05, Ove06, Ove07, PE09, SW08]. Council [Gul09]. councils [Cor05, Oke03]. countries [Han08, PKS^{+}09, SGP^{+}08, WNB^{+}08, Wit08]. country [Som03, Van04]. course [Fle01]. crab [Fin05, Lou07]. crafting [KS04]. Creating [PJ08]. creation [JBN07]. Crew [Har09, WW05]. crisis [dVM02b]. criteria [AMS05, Nob00]. critical [CGHA08, MCO08]. critique [Cor05, DKM08]. Cross [DSH03, Son05, WASK06, DW08]. Cross-scale [WASK06, DW08]. Crossing [JC09b, MR01]. cruise [But07, Joh02]. cruising [Mar04]. cultural [Cla09, Ove06, Ove07]. culture [DCHM07, KPN^{+}09, MGL09, SLT08]. Current [Eks09, BET^{+}07, Mal03, PGM^{+}05, FF06]. currents [DSH03]. cycles [BK09].

damage [FW08, Mas03]. damages [FFM06, GGSRVL06]. Danish [NC06]. data [BHJV09, BBNB08, CROP09, Mas08, SRW06, ZFP05]. data-poor [BHJV09]. dataset [Eks09]. date [VA05b]. Days [RS03]. de-regulation [AW02]. Death [NVPR09]. debate [Mat01, WDAT00]. decade [Eyt00, HK06a, Wit06, vSBHS06]. decades [Hun09]. Decentralization [SM04]. decision [Nau06]. decisions [PL04]. decline [KHG08, Pol06, Sta06, dVM02b]. decommissioning [Ham03]. decrease [KF09]. deep [HF02]. deep-sea [HF02]. defence [OLC^{+}09]. Defining [DCNP09, Hil07, dJDT09, Foc08]. definition [Hay05, PGP08]. deliberation [Var08]. delivery [Fle01]. Demand [McG03a]. Demand-side [McG03a]. demise [SWM06]. democratic [Fle03]. demolition [KKR08]. Denmark [MPB^{+}02, NM03]. dependency [DJ01, HHR00]. depletion [HB07b]. Design [TMY06a, Fle01]. designated [Ber08]. Designation [BCM09a]. Designing [AMW09, JC09b]. desirable [Ili08a]. detect [Mal03]. deterioration [TG09]. determines [Fra08]. Deterring [Erc06]. Developing [CLyB06, HESM08, RWB09, Bee07, GL08, QWJA09, Syd02, TRvAB05, Wit08, Yan03, Gai08, KKI^{+}09]. Development [SM06, WD07, AH06, Ben05, BH00, Fle01, Han08, HR04, HH04, JHW09, KKI^{+}09, LOT05, Mah06, NSVW01, Pet02, RBT04, Sah01, SJ02, TRvAB05, Var08]. developmental [HESM08]. developments [Ber02, Kea01, Key01, Pot00, Wit06]. devices [SDWF09]. devolution [dVMdC08]. dialogue [Val04c]. Did [Pol06]. differ [Arm01]. differences [Bat05, PC00]. Different [CLT08]. differentiate [KF08]. differentiation [LOS06]. dilemma [Pol02b]. dimension [Jen06, KM04]. dimensional [JBP09]. dimensions [Ebb03, Ebb09, dV07]. Dingle [Kar01]. diplomacy [Coo05, Jab06, Jab09]. direct [CLyB06]. directed [Lev09b]. direction [Che08]. Directions [FP04, NC06, dSK06]. directive [Hew07]. disagreement [Val05a]. disagreements [Dja04]. disappearing [Van05]. disappointment [Jen00b]. disaster [PRH^{+}06]. discard [CFG06]. discarding [BRH06]. Discards [CFG05]. disciplinary [RSS^{+}07]. disciplines [Hum03, Pon03]. discovery [CROP09]. Disentangling [KHG08]. dispute [BC07b, Mor06, Pol02a, VVG03]. distant [HB00, HB01, Xue06]. distribution [BRS08, EY07, Han07b, Mor00, Til08]. Distributional [BM06, GT01]. distributive [Bee07]. disturbance [CGHA08, dJDT09]. Dive [Fab08]. divergent [Eva09]. diverse [DS09]. divide [JC09b]. Do [PG09, Gle08, Jen04, KF09]. documentation [Agn00]. Does [KF09, Gib09b, Gui04, Nie09]. dolphin [Bro05]. dolphin-safe [Bro05]. domain [Nau06]. domestic [RBT04]. Draft [DYB05]. drawing [SBF^{+}04]. dreams [Isa06]. driven [HR06]. drivers [LC06]. dry [RM02]. Duplication [FBH^{+}07]. durations [GM01a]. during [CGHA08, TK09]. dynamics [JM04].

East [Bjø09, CS07, DJ07, Kim03b, NDV^{+}04, Syd01, Vin06, FHCS05, AVH05, Ber02, Key03, Skj06, TS05]. eastern [FHCS05, LCK04, RHCS05]. eclectic [Sta08]. ecological [BCM09a, CN08, SR08, SDWF09, TTSS09, WRD06, dJDT09]. ecologically [JC09b]. Econometric [KF07, KF08, KKR08]. Economic [ANL09, GTLC02, Hun03, Pet02, BC07a, BB06, BK09, FH04, GGSRVL06, Gui04, Han08, HH06, IMP09, Key01, LDBB08, LC06, MGL09, NVPR09, NSVW01, Nøs08, Ove06, Ove07, PFS09, Sch08, Sub07, SVL07, TTSS09, Val05a, VA06, Van04, Van05, WJPN00, WPBD03, WCB06, dVM04, vdBND^{+}02]. Economics [LFI01, FA06, SLCS05, SAK06, vdB00]. economies [McG07]. economy [CvSP06, HB01, KYC05, VES06]. Ecosystem [FPS05, TLHN07, BM08, BWR08, CROP09, CFW^{+}07, Cla09, CN08, DPE^{+}08, Dou08, FMM^{+}07, FF06, Gai08, GDI^{+}09, GL08, GH08, HCL^{+}09, MFM09, MMS09, MB09, Mor08, Mur07, PS07, PKS^{+}09, Var08, dlM05, Eks09]. Ecosystem-based [FPS05, BWR08, CROP09, CFW^{+}07, Cla09, CN08, DPE^{+}08, Dou08, FF06, Gai08, GDI^{+}09, GL08, GH08, MMS09, PKS^{+}09, dlM05]. ecosystems [CVL02, Det06, Gai08, Nau06, vdBND^{+}02]. Ed [Ano04a]. Editorial [Ano02a, Ano03d, Ano03e, Ano03f, Ano03a, Ano03b, Ano03c, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano05a, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano06a, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano07a, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f]. Education [KKI^{+}09, How01, KTH^{+}07, Vee02]. EEM [CWP06]. EEZ [Bat05, DYB05, Foc08, Hay05, Ska05]. EEZs [Bal04]. effect [KF07]. effective [BT04, Haa00b, Haa00a, NBR09, Sto09, Val00]. Effectiveness [LFL05a, SRVLIM03, SR05, BET^{+}07, Gul09, LLJ07, PBSS06, LFL05b]. Effects [McG07, BM06, Cho07, CWP06, HH04, KF08, RHKS09, WBGC08]. efficiency [DTFÁIM04, IMP09, PC00]. effort [OJ08]. eight [Oke03]. elements [GL08, Moe08]. elephant [PE09]. eliciting [WP09]. Emergence [HR06, Gul09]. Emerging [Smi03, DJ07, dSK06]. emission [LL06]. empirical [HNMB07]. Employment [GM01b, Mit08, GM01a, PL01, PGM^{+}05]. Empowering [Fle01]. empowerment [Jen05, WCKB09]. enactment [Kim02]. Endangered [Ber08, RWB08]. energy [Mar08, PF02, Por09, SDWF09]. enforcement [DYB05, Fah08, Nøs08]. engagement [PD08]. engineering [Utn06]. engines [HH04]. England [DP00, CS07, JBN07, Kap00, Mar08]. English [CFG06]. enhance [FD03, Roa04]. Enhanced [Ebb02]. enhancement [MBJI08, PBSS06]. Enhancing [WCKB09, Skj06]. enlargement [dVM06]. entrepreneurs [Bee07]. entry [PL04]. environment [BR07, FD03, VGH^{+}08, Ver07]. Environmental [EOU09, HH04, Ugg07, Ben01, CWP06, FPHP09, Fra08, FE09, FFM06, Mas03, Sch08, Ver07]. Environmentally [Joh02]. Equity [Jon09, Kap00, Mar01]. era [JP07]. Erik [SR05]. ESD [KKI^{+}09]. Essential [CN08, BHJV09]. ESSIM [RHCS05, FHCS05]. establishing [Gib09a, Syd01]. Establishment [HMM^{+}04, Lev09b]. estimates [SWC09, ZYN09]. Estimating [NVPR09, DCNP09]. Estonian [VEH02, VES06]. estuaries [Tow02]. Estuary [FBH^{+}07, MM04, TG09, CVL02, KS04, SR08]. ethical [Ebb09]. ethics [Mau02, Sch08]. EU [Gez08, IMGGVL02, MG09, Mit08, PLB04, QCB09, dVM06, dV07, vHvT09]. Eubalaena [Lev09a, Lev09b]. Europe [HS06, MJ08, OB09, PGP08, QCB09, dVM06]. Europe. [CC09]. European [Det06, LFL05a, RTWJ05, SR05, ANL09, BWR08, BBB01, BM07, CW00, DG05, FLT06, FGA00, FA06, GP00, Gui04, Hat00, HBB00, KF02, LFI01, LFL05b, MN06, MN08, MCO08, Mor00, OB09, SRVLGG02, SRVLIM03, Wit08, dV07]. Evaluating [Dav08, Day08, Ber08]. Evaluation [Cho06, PPP^{+}08, GM09, LW07, NH07, NZF09, PKS^{+}09, SU07, Utn07]. everywhere [dJDT09]. Evidence [CMW09, BNA^{+}01, IMP09]. evolution [Car07, Fle00, HR04, Joh05, KdSR07]. evolving [Mas03, QCB09]. Ex [PGM^{+}05, GM09]. Ex-seafarers [PGM^{+}05]. Examination [SM08, CGHA08]. examining [Mas03, BH09]. example [PS07, Wit08]. excess [SGP^{+}08]. exchange [MN06]. exclusive [Han08, Key01, PFS09, Sch08, Val05a, VA06, Van04, Van05]. exit [PL04]. exiters [SWM06, SW08]. Exotic [vdBND^{+}02]. expand [Joh05]. Expanding [Bes05, MMS09, RNG06]. expansion [Sta06]. Expectations [Isa06]. experience [CBB00, Hea06, JHW09, Yan03, Lou07, PG09, WK09]. experiences [BBB01, NC06]. experiment [AVH05, KPN^{+}09, Rep01]. Experimental [Pol02a, Ver07]. exploration [WD07]. exploratory [RGB09]. export [Bes06]. extended [Jab06, Mar04]. extended-length [Mar04]. extension [HHM^{+}09, KTH^{+}07]. extractive [dS04].

face [BC06]. faced [KS04]. facilities [Por09]. facing [Bod02]. factors [CMW09, CLT08, NM03, OJ08, dVM04, dVMdC09]. factory [Sta08]. Faeroe [Gez08]. failed [AVH05]. failing [TWN^{+}09]. fails [RC01]. failure [DG05, GMN^{+}07, Kim02, Lou07, Pon03]. failures [FGA00, NH07]. fall [Sta08]. Far [AVH05]. farm [ENS04, MCK08]. farmed [EOU09, LS08]. Farming [PPB06, Gui04, KA06, RWB09, SU07, SVL07]. feasibility [CFW^{+}07]. features [MM05]. federal [Eks09, MB09]. fee [Mat01]. feeder [CLT08]. fees [Edw09]. Fiji [Lan08, TTSS09]. fin [Ser04]. final [Sta07]. Financial [FW08, MM02]. Financing [RGB09, Edw09]. fins [Ser04]. firm [Bes06]. firms [Mit04]. first [Edw08, Sco00, dS04, vSBHS06, Hea08]. Firth [SDWF09]. Fish [AVH05, BC06, JHMC07, MR01, PPP^{+}07, AL02, Arm01, BKV^{+}09, BC07b, Gui04, Han07b, Hun06, JGU09, Kap00, Mil07, Nie09, PPP^{+}08, SLT08, SS07, TRvAB05, TMY06b]. fish-producing [TRvAB05]. fished [dJDT09]. Fisher [AEG^{+}08, Mau02, SWM06, SW08]. Fisheries [ÅB02, Ano06g, Bjø09, DG05, Foc08, FA06, GH03, HK06a, HMM^{+}04, HB07b, Jen05, JC09a, Kea01, MG09, NDV^{+}04, Nøs08, NSH01, Per09, TWN^{+}09, AS07, AMS05, AL02, AE01, AH06, Alp06, AHS00, BdKvD^{+}08, BNA^{+}01, Ben05, BWR08, Ber08, Bes01, BR07, BB06, BCM09b, BBB01, BT04, BW09, BC06, Bra08, BL08, Cam00, CSG^{+}09, CFG05, CFW^{+}07, CJ07, Cla06, CVL02, CW00, CM00, CvSP06, DCNP09, Dav08, DGH^{+}06, DTFÁIM04, Ebb09, EE03, Eyt00, FP04, FF06, Fin05, FGA00, FPS05, FSPKR09, Gez06, Gib07, Gib08, GDM06, GCB^{+}08, GL05, GKMR07, GHR^{+}08, GM09, GH08, Gry03, GL03, GT01, HT09, Han03, Har01, Har07, HB07a, HESM08, Hat00, HS01, Hau08, HB00, HB01, HS06, HH00, HHR00, Hil07, HK06b, HG01, Hol00, Høn00, HJ02]. fisheries [HCK00, Hum03, IMP09, IMGGVL02, Isa06, JP07, Jen02, Jen00a, Jen00b, Jen04, JM04, Jen06, Jen07, Joh05, Joh06, JHMC07, KF02, KS04, KKP00, KM04, Kar01, Kim03b, KE03, KA06, LS00, LDBB08, LW07, Lev08, LFL05a, LFL05b, LB04, Lou07, Mah06, MMR08, MM05, MMS09, MPB^{+}02, MG06, McG03b, McG07, MJ01, MJ03, MHJ07, MJ08, MBJI08, MWL^{+}09, Mor00, Mor08, MM02, NVPR09, Nie03, NM03, NC06, Nie06, NH06, NH07, Nob03, OJ08, Par00, PG09, PPW^{+}09, PFS09, Pet02, PKH02, PKS^{+}09, Pol02b, PBB08, PKH01, PPP^{+}07, PNT09, PS01, Pon01, Pon03, PP04, RSS^{+}07, RWB08, Rep01, RNG06, SGP^{+}08, SM04, SMS06, SR08, Sch07a, Sch07b, SM07, SH03, Som03, Sow06, Sta05, Sta06, SC04, SRVLIM03, SR05, Syd01, Syd02, TTSS09]. fisheries [Tho08, TG09, TRvAB05, Utn06, Utn07, VL02, VGH^{+}08, VEH02, VES06, WD07, WKT00, WK09, WJPN00, WPBD03, WCKB09, WRD06, Wit06, Wit08, Xue06, Yan03, Yan08, YM06, ZFP05, dSK06, dVMdC08, vHvT09, vS02, vSBHS06, vSH06, vZvDT02, vdBL07, vdB00, WASK06]. Fishermen [Bod02, DKM08, Høn00, KKP00, KE03, SBF^{+}04, Val09]. fishers [BR07, BBNB08, PL04, Sch07c, LOS06, RS03]. Fishery [EE03, KHG08, Oke03, Agn00, AOPM00, BRS08, Bod02, BM06, BJBC09, CBB00, CFG06, CGS08, Cor05, Daw06, ENS04, EOU09, Fah08, FL01, Gai08, GGVLIM03, Kan03, Key03, KdSR07, LWH^{+}09, LFI01, Lev09a, Lev09b, LOT05, Lou07, MS02, McG03a, NZF09, PKH02, Pol02b, PS01, RSJ^{+}03, RC01, Sch05, Sch08, SA02, SWM06, SW08, SJ03, Syd02, TG09, WBCK04, DP00]. Fishing [FH04, Jon08, Mau02, PL08, PBB08, Sta03, WNB^{+}08, YY08, AMW09, ANL09, BM09, BM07, CGS08, Che03, DJ01, Ebb09, Erc06, Fab08, Gez06, Gez08, GL07, Hat00, HFPR02, HS06, Jen02, JGU09, Lax06, LC06, Let00, LLJ07, Mat01, McG07, MD08, OJ08, Ove06, Ove07, PC00, PGP08, Pet03, PSFE07, PNN08, PE09, Pol06, PPB06, RSJ^{+}03, SHQF06, SR08, SU07, SWC09, Sto09, SAK06, SRVLGG02, Utn08a, Utn08b, WBGC08]. fit [Ebb02, KTH^{+}07]. Fitting [CVL02]. Five [AEG^{+}08, TS05]. fixes [DGH^{+}06]. Fixing [BW09]. flag [Xue06]. flagging [HY03]. Fleet [GDM06, OJ08, DTFÁIM04, GGVL05, GGAD08, GGVL09, PSFE07, Sta07, SRVLGG02, Tza05, Utn08a]. fleets [Hat00, dC08]. Flexible [HS06]. floating [MGL09]. floor [DGH^{+}06]. Florida [Lev09b]. flying [BC07b]. Focusing [dC08, EOU09]. food [BKV^{+}09, PPP^{+}08, SWC09]. footing [Glo03]. force [Let00]. foreign [HFPR02, Pet03, Ska05]. Foreword [VA04, VA05a]. Formal [WF01]. formalisation [Kar01]. formation [AvdH08]. forms [McG03a]. Forth [MM04]. Forum [FBH^{+}07]. forward [Val04a, Val05b]. Fostering [Tho08]. found [BR08]. foundations [MGL09]. fragmented [EYG^{+}09]. framework [Cla09, FMM^{+}07, KKI^{+}09, Mae08, Nie03, QCB09, DPE^{+}08]. frameworks [Elv09, MN08]. Framing [NH07]. France [MPB^{+}02]. free [Elv09]. freedom [Ugg07, Van05]. frontier [Sta03]. Frost [SR05]. Frozen [AL02]. fuel [WDAT00]. Funding [Har07]. Further [LFL05b]. future [CvSP06, Ehl08, Llá03, NC06, SJ02, Sta06, dSK06]. fuzzy [HY03].

gains [ANL09, Nie09]. Galicia [FMPPF08, FGA00]. Galician [GGAD08, MGL09]. gap [Hin03]. gaps [WdK02]. gas [FE09]. gathering [DYB05, Hay05, Val05a]. GATT [JM00]. GBRWHA [NBR09]. Gender [Ben05, DSH03]. generated [But07]. genetic [LVH^{+}09]. Geo [SHA08]. Geo-technologies [SHA08]. Geographical [Han07b, Val01, dVMdC09]. Geopolitical [dVMdC09, Kay04, dVM06]. geopolitics [CSG^{+}09]. German [Foc08, MG09, PFS09]. get [SRMM03]. Ghana [BNA^{+}01, PS07]. Ghost [BM07]. GIS [BET^{+}07]. GIS-based [BET^{+}07]. glacialis [Lev09a, Lev09b]. glass [BR08, vSH06]. Global [SAK06, BWA^{+}03, CvSP06, DCNP09, Kay04, LTX08, TS05, WW05, Mal03]. Globalisation [AW02, FL01, Hun03]. globalization [dVM02b]. goals [Var08]. gold [Dro04]. GOOS [Mal03]. governability [Jen07]. Governance [BC07a, Cho06, Cla09, EYG^{+}09, Eva09, FMM^{+}07, FMPPF08, Gib08, GKMR07, GH03, Haa00b, Haa00a, Har09, Hin03, Jen07, JC09a, Joh06, KHG08, Lan08, MFM09, MHJ07, OB09, SP01, Sto09, TG09, Val00, Val01, VH09, dVM04, dVMdC08]. Governing [MMR08, EY07, MCK08, Ser04]. government [Dro04, MB09, MM02, Pet08]. Grand [DTFÁIM04]. Grant [KTH^{+}07]. Great [Day08, NBR09]. Greek [FLT06, Tza05, TK09]. Greek-owned [FLT06]. Green [dV07]. Greenland [GGVL09]. grey [Coo05, TLHN07]. Groundfish [DP00, HG01]. groundwork [MN09]. group [Hol00]. grouper [Hal03]. groups [HCK00, MJ01, MHJ07]. growing [LS08]. guarding [BdKvD^{+}08]. Guatemala [Per09]. Guidelines [VA06, CGHA08, DYB05]. Gulf [JHMC07, Lou07, BP08, KP04a, KP04b, KM09, RC01, WPBD03, YM06]. Gungliau [LOT05].

habitat [JBN07, TRF08]. halibut [Daw06, GGVL09]. Hamble [FBH^{+}07]. hammer [DGH^{+}06]. Hampshire [JBN07]. Hans [SR05]. Harbour [JBP09, OGC^{+}09]. harmful [vdBND^{+}02]. harvest [Joh05, JHMC07]. Hatchery [PBSS06]. Hawaii [NZF09, PL04]. Hawaiian [GL03]. hazard [DCHM07]. hazards [MN09]. health [JGU09, MN09, WNB^{+}08]. help [Kar01]. Heritage [GA07, NBR09, Sub07, Cla09, MMS09]. Herring [BM06, BP08, SB05]. hidden [OJ08, PE09]. hierarchical [dlM05]. High [Jab06, AGPT08, Jab09, Kay04, McG07, Ove07, SLT08, Sta03, Ugg07, Sag05]. high-latitude [Jab09, McG07]. high-value [SLT08]. highly [Mil07, Ser04]. Historic [FTF00]. Historical [MG09, AS07, FW08]. history [MJ03]. hoki [BJBC09]. home [But07]. Honduras [Per09]. Hong [JHW09]. hope [Sch07b]. hopes [Isa06]. hopping [Mor00, HFPR02]. Horn [PBB08]. human [PS07]. humane [GLP08]. hurricane [KK08]. Hydrographic [Bat05]. hydrothermal [Glo03].

Iceland [AHS00]. Icelandic [Eyt00, Mat01]. ICES [SC04]. ICM [Fle01, FD03]. ICZM [MCO08, CS07, HCL^{+}09, QCB09]. Identification [Sag05]. ideology [dVM02b]. IFC [Ano04f]. IFQ [CM00, Daw06]. Iglesias [LFL05a, SR05, LFL05b]. Iglesias-Malvido [LFL05a, SR05, LFL05b]. II [PNN08, Val05a]. illegal [Agn00, Bod02, CGS08, LC06, SAK06]. illustration [Pon01]. imbroglio [VL02]. impact [BK09, But07, HS08, HHM^{+}09, JGU09, KK08, NVPR09, PL01, Por09, JHW09]. Impacts [SS07, Hun03, LCK04, SDWF09, vdBND^{+}02, BM07]. impediments [LS00]. imperilled [Det06]. Implementation [FE09, Gib09a, CJ07, DYB05, GDI^{+}09, Hin03, KS04, MB09, RSS^{+}07, CMCNL^{+}09]. Implementing [FHCS05, HJ02, Kay04, Nau06, Pla08, PKS^{+}09, Vin06]. Implications [Hew07, PC00, RBT04, Bal04, BB06, BRH06, DTFÁIM04, GMN^{+}07, Gry03, Hin01, Jen07, Kea01, Lev09a, Lev09b, LTX08, RSJ^{+}03, Sch08, SH03, SB05, SHA08, SWC09, ZYN09, McG07, Ver07]. importance [CROP09, Dou08, TLHN07]. Important [NM03]. imports [GP00, Mar08]. improve [KF09, LLJ07]. improved [KP04b, KF07, vdB00]. Improvement [TG09]. Improving [Elv09, Wes03, vZvDT02]. incidental [Gez08]. income [BRS08, SWC09]. Index [Ano00a, Ano01a, Ano02b, Ano03g, Ano04g, Ano06h, Ano06i, CMCNL^{+}09]. India [BdKvD^{+}08, LB04, TG09]. Indian [CWRR09, BB06, KPN^{+}09, Pol06, Sha05]. indicator [LOT05]. indicators [RSS^{+}07, dJDT09]. indigenous [Bes01, DJ01, Hin01]. Individual [CBB00, AEG^{+}08, BRH06, Gib07, GM09, HHR00, PG09, PE09, SHQF06, TMY06a, Tro04]. Indonesia [BRS08, Cli03, Hal03, HB07b, NSH01, SM04]. industrial [HR04, NH06, Ove06, Ove07, Pol06]. industries [BC07a]. Industry [MS02, AW02, BH00, CWP06, DM09, EOU09, FLT06, FE09, Gui04, Har09, Har01, HS06, Jon08, KYC05, LTX08, MD08, MM02, RBT04]. industry-led [Har01]. influence [SJ02, TRvAB05]. influencing [NM03]. Information [Ho09a, BM09, Jen08, MN06, Mas08, Nau06, SLCS05, WJPN00]. infrastructure [LCK04, SRW06]. infusion [Wit08]. initial [FFM06]. initiative [FHCS05, RHCS05]. Initiatives [Roa04, dSK06]. injuries [BBNB08]. Innovation [DM09, LB04, NDV^{+}04]. Innovation-support [DM09]. innovations [HH04]. input [KYC05, McG03a]. inshore [CWRR09, Fah08, Kar01, TTSS09]. insights [CN08]. inspections [BK09, CMW09, KF07, KF08]. instability [VES06]. installations [Ham03]. Institutional [Ebb09, LS00, Nob00, RWB08, Sta07, AJ09, Ebb02, Ebb03, Elv09, Fra08, HH00, Hin01, LB04, Lou07, MM05, MGL09, NDV^{+}04, Sch08, Skj06, Yan08, YM06, vdB00, Jen07, Sto09]. Institutionalism [MG09]. Institutions [Jen04, CWRR09, CVL02, HG01, Pet02]. instrument [HR04, MN08]. instruments [Ser04]. Insularity [Pon03]. insurance [Ben01]. Integrated [QCB09, RHCS05, DSH03, FHCS05, GA07, LW07, Mas08, OGC^{+}09, Tau07, WCB06, vdBND^{+}02, MCO08]. integrating [BW09, McG03a, SJ03]. integration [CC09, CvSP06, Daw06, FBH^{+}07, Lan08, dVM02a, MCK08]. Intelligence [Bal04, DYB05, Hay05, Val05a]. intentionally [Ver07]. Interaction [VGH^{+}08]. interactions [BB06, GCB^{+}08, Son05]. Interdependence [KA06]. Interdisciplinary [WdK02]. interest [LVH^{+}09, MJ01, PKH02]. interests [BET^{+}07, Elv09, Oke03, WW06]. Intergovernmentalism [MG09]. interlocal [MLRK08]. internal [Som01]. International [DE07, Ham03, HRP08, Lev08, RT08, DG05, Det06, FTF00, HJ02, Ili08b, Kim03a, LL06, Mae08, MN08, Mas03, Syd02, Hea04, Hea06, Ili08c, Ili08d, Lev08]. interplay [Ebb02, Skj06, SC04, Sto09, Sto07]. interpretation [BET^{+}07]. Interviewing [CS07]. introducing [GGVL09]. Introduction [DE08, Han07c, Val04b, Val05a, VGH^{+}08, vSH06, FJH07, Kar01, Mat01, RSS^{+}07]. invasive [BWA^{+}03]. Investigation [LLJ07, Ove06, Ove07, WCB06]. involve [Som03]. involved [SRMM03]. Involving [Por09]. IODE [Mas08]. Ireland [FÓ08, Kar01, OGC^{+}09]. Irish [BET^{+}07, BR08]. irradiated [WDAT00]. ISC [Ho09a]. Island [AMS05, RBT04, Han08, HB07a, BJBC09]. Islands [Fab08, BPM09, Smi03, Fin05, Gez08, GL03, HT09, VS01]. ISM [TK09]. ISQ [Tro04]. Israel [Fle00]. issue [ABF00a, ABF00b, Bro05, CCG^{+}07, Mor06, SP01]. issues [FPHP09, GT01, Jon08, Jon09, Lan08, Smi03, Son05]. ITQ [AMW09, Eyt00, SA08]. ITQ-management [Eyt00]. ITQs [BW09, CGS08, VEH02]. IUU [Erc06, GL07, LC06, Sto09]. IVQ [SA08]. IWC [Ili08a].

Jamaica [Edw09, RGB09]. January [Ano00b, Ano01b, Ano02c, Ano03h, Ano04h, Ano05g, Ano08g, Ano09g]. Japan [MBJI08, AMS05, BH09, BP08, EY07, GLP08, Kan03, MMS09, Sco00, VL02]. Japanese [BP08, HB01, Key03, MM05]. jobs [GMN^{+}07]. joint [Key03]. Jørgen [SR05]. judicial [Dev00]. July [Ano00d, Ano01d, Ano02e, Ano04j, Ano05i, Ano06l, Ano07i, Ano08i, Ano09j]. jurisdiction [Dev00]. justice [Jon09].

Kabui [BRS08]. Kagoshima [AMS05]. keep [LS08]. Kenya [Eva09]. Kerala [TG09, vdBL07]. Kerry [Kar01]. Key [GL08, BC07a, Hay05, LC06]. keys [Pon03, BR08]. Kingdom [Pet08]. know [Gle08]. knowledge [Hea08, HHM^{+}09, Mau02, SR08, SBF^{+}04, WRD06]. known [Gib07]. Kong [JHW09]. Korea [VL02, VVG03, Che03, Cho06, Cho07, CW08, Kan03, Kim04, MM02]. Korean [kK05, KYC05, RNG06].

labour [Gle08, WW05]. Labrador [BBNB08]. Lafuente [LFL05a, SR05, LFL05b]. Lagoon [KS04, JBN07, CVL02, SR08]. land [GMN^{+}07]. land-based [GMN^{+}07]. Landlords [Mar01]. Landscape [HCL^{+}09]. Landscape-scale [HCL^{+}09]. Large [Eks09, FMM^{+}07, MFM09, Pol06, SRMM03]. largest [LWH^{+}09]. last [Sta03]. Latin [KTH^{+}07]. latitude [Jab06, Jab09, McG07]. Law [Wit06, Kan03, Kim03b, Nøs08, Ska05, Som01, Syd02, Sto07, Ver07]. Lawrence [BP08, Lou07]. laws [Eks09]. layer [SHA08]. Laying [MN09]. Leaning [MHJ07]. learned [Ehl08]. learning [BBB01, Fle01, FD03, NH07]. leaseholders [Mar01]. leasing [PE09]. led [Har01, OLC^{+}09]. legal [BCM09a, Mae08, RT08, Ser04, Sto07, vdBND^{+}02]. legislation [Bes06, BET^{+}07]. Legislative [Wit06]. Legitimacy [Jen00b, Nie03, PJ08]. length [Mar04]. Leone [TWN^{+}09]. Lesser [Gib07]. Lesser-known [Gib07]. Lessons [BBNB08, Hea06, BM08, Day08, Ehl08, GDI^{+}09, HJ02, HBB00, Kap00, McB04, NM03, NBR09, PKH01, VG03, dS04, Har07, NDV^{+}04]. level [CW00]. liability [GL07, Kim03a, Mas03]. Liberal [MG09]. liberalisation [Nie09]. liberalising [ANL09]. liberalization [Chi07]. Liberum [RZ03]. licences [PG09]. licensing [Bod02]. life [BK09, Isa06, San05]. Limitations [RNG06]. Limits [Jen07, MJ03, BRH06]. Lindebo [SR05]. line [Ove06, Ove07]. Liner [MN06]. liners [CLT08]. lines [CLT08]. link [FH04, Jen00a]. linkages [DW08, DSH03, WASK06]. linked [KTH^{+}07]. linking [BBNB08]. Lion [Ber08, GTLC02]. literature [Nøs08, VG03, TLHN07]. litter [SM06]. live [PPP^{+}08]. livelihood [SLT08]. Livelihoods [AH06, AE01, PRH^{+}06]. Lobster [Coo05, Bod02, PKH02]. Local [How01, WRD06, AMS05, CW00, FH04, FL01, Hew07, LB04, PJ08]. logistics [MN08]. Løkkergaard [SR05]. long [LDBB08]. long-term [LDBB08]. longline [GCB^{+}08, NZF09, Ove06, Ove07, Pol06, PL04, WBGC08]. looking [BR08, vSH06]. looking-glass [BR08]. looming [Cla06]. lopsided [SA02]. losing [ZFP05]. losses [Nie09]. Louisiana [Kai06, KK08]. Low [CMCNL^{+}09]. Low-cost [CMCNL^{+}09].

M [Cho07]. M/V [Cho07]. Mabini [Maj07]. machinery [Joh05]. Made [BP08]. Mainport [HHM^{+}09]. maintenance [MWV02]. major [KF09, TRvAB05]. making [Hea04, JM04, TRvAB05, MJ08]. Malacca [Val03]. Malaysia [Val03]. Malaysian [Sah01]. Malthusian [TS07]. Maluku [NSH01]. Malvido [LFL05a, SR05, LFL05b]. mammal [MWL^{+}09]. mammals [CGHA08, Hun09]. manage [CW00]. managed [Ebb09, MG06, MM04, MBSLP02, PG09, PPW^{+}09, Tho08]. Management [Gez08, HNMB07, IMGGVL02, Li03, McB04, MCO08, Oke03, ÅB02, AL02, AE01, Alp06, AHS00, BM08, BA05, Bee07, BNA^{+}01, BWR08, Bes05, Bes06, BT04, BW09, BC06, BR08, BM07, BL08, BP08, CSG^{+}09, Car07, CROP09, Che08, Cho07, CFW^{+}07, CJ07, CWRR09, Cla09, Cla06, Cli03, CSH00, CVL02, CN08, Dav08, DPE^{+}08, Day08, DGH^{+}06, DSH03, DTFÁIM04, DMVS07, Dou08, Ebb02, Ebb03, Edw09, EE03, FF06, Fle00, FJH07, FP07, FHCS05, FE09, FGA00, FFM06, FPS05, Gai08, GJGT04, GDI^{+}09, GL05, GA07, GK05, GH08, Gry03, GL03, GT01, HF02, Har07, HB07a, HS01, Hea06, HS06, Hol00, HCK00, HBB00, Hum03, JBP09, Jen00a, Jen00b, Jen05, Jen06, JHMC07, KS04, Kap00, KKP00, KM04, Kar01, Key03, KdSR07, LWH^{+}09, Lev09a]. management [Lev09b, Llá04, LB04, MM05, MMS09, MPT09, MPB^{+}02, Mas08, MB09, McG03a, MLRK08, MCK08, MM04, MJ01, MJ03, MBJI08, Mit04, MWL^{+}09, Mor08, Mur07, Nie03, NDV^{+}04, NC06, NH07, Nob00, Nob03, NSH01, NBR09, OLC^{+}09, OGC^{+}09, Par00, Per09, PKH02, PJ08, PKS^{+}09, Pol02b, PBB08, PKH01, PPP^{+}07, PS01, Pon01, Pon03, QCB09, RSS^{+}07, RBT04, RSJ^{+}03, Rep01, RHCS05, RNG06, SM04, Sch07b, Sch08, Sch07c, SH03, SJ02, SRMM03, SM08, SB05, Som03, Sta05, SM06, SJ03, Syd01, Tau07, TG09, Tow02, TMY06b, Utn06, WKT00, WK09, Wes03, WBCK04, WASK06, WRD06, Yan03, Yan08, YM06, YY08, dS04, dSK06, dlM05, vZvDT02, vdB00, Llá05a, Mil07, TLHN07, WR07, Eyt00, Lev08]. Managing [Cam00, DKM08, HT09, MCO08, SGP^{+}08, Gib07]. manner [DYB05]. Mapping [PFS09]. March [Ano00g, Ano01f, Ano02h, Ano03l, Ano04m, Ano06j, Ano07k, Ano09h]. Mare [RZ03]. mariculture [Hal03, JM00, MCK08]. Marine [AGPT08, ABF00b, BM08, BPM09, BWA^{+}03, Eks09, FÓ08, GDI^{+}09, LVH^{+}09, LFL05a, LFL05b, PS07, San05, SRW06, SR05, Tow02, Vin06, WDAT00, dlM05, BHJV09, Bat05, BR07, BCM09a, CWRR09, Cli03, CCCD07, CGHA08, CN08, DCNP09, DPE^{+}08, Day08, DJ07, Det06, DM09, DMVS07, Dou08, DP00, EO09, Ehl08, Eva09, Fab08, FMM^{+}07, FPHP09, Foc08, FHCS05, Gib07, Gib09b, GL08, Glo03, GK05, Haa00b, Haa00a, Hin01, Hun03, Hun09, HH04, Jon08, Jon09, Mae08, MFM09, MWL^{+}09, Mur07, Nau06, Ove06, PGP08, Pla08, PD08, PNT09, QWJA09, QCB09, RGB09, RWB09, SBF^{+}04, SDWF09, SJ02, SHA08, Sub07, Val00, VS01, Ver07, WW06, WJPN00, WP09, ZFP05, dVMdC09, vHvT09, vSH06, vdBND^{+}02, Ano07g, DPE^{+}08, DE07, Gul09, RSS^{+}07]. Maritime [Bod02, DS09, Pot00, QCB09, dVM06, AW02, Ben01, Cho07, Cla09, DCHM07, Gre01, Ho09b, KYC05, Llá04, MWV02, PL01, Roa04, Som01, dS04, dVM02a, dVM02b, dVMdC09, vHvT09, dV07]. market [BP08, Elv09, GMN^{+}07, Gle08, Kap00, KKR08, Tro04, TMY06b, VES06, WW05]. market-oriented [TMY06b]. marketing [KE03]. markets [Bes06]. Marsh [MBSLP02]. matrix [Mor06]. matter [KF09]. Mauritius [Hol00]. May [Ano00c, Ano01c, Ano02d, Ano03i, Ano04i, Ano05h, Ano06k, Ano07h, Ano08h, Ano09i]. mean [Gib09b]. means [DYB05]. MEAs [SM08]. measurement [FP04, PC00]. measures [Sto09, WDAT00]. Measuring [WK09]. Mechanism [HMM^{+}04]. Mechanisms [Sub07]. Mechanization [BB06]. Medal [Bor01]. Media [CCCD07]. Mediterranean [Fra08, IMP09, Kap00, dJDT09, dVM02a]. meeting [VA05b]. mega [Li03]. mega-cities [Li03]. members [Har09]. Membership [Oke03]. Menticirrhus [Lev09b]. merchant [Kin05, LL06]. messages [MD08]. meta [MPT09]. meta-analysis [MPT09]. methodological [WW05]. methods [WP09]. Mexico [JHMC07, KP04a, KP04b, KM09]. migrant [BRS08]. migratory [Mil07, Ser04, SB05]. Military [Hay05, Van04, DYB05, Val05a]. minimise [WBGC08]. minimising [CGHA08]. mining [HF02]. mislabeling [JP08]. missing [HHR00, Jen00a, SHA08]. mistake [FW08]. mitigation [BL08, JHW09]. Mixed [MD08]. Mobility [SH03]. Modalities’ [Xiz04]. model [HNMB07, KP04a, KTH^{+}07, RWB09, TMY06b]. Modeling [PL04, WdK02]. Modelling [Nie09, KPN^{+}09]. modern [Val01]. Modernisation [Ili08c]. module [Mal03]. Monitoring [Con07, Gez06, Hea08, VG03, CWP06, Day08]. moratorium [Sch05]. mortality [Gez06, WBGC08]. most [Utn08b]. mouths [RS03]. move [vHvT09]. MPAs [BW09, JC09b]. Mueang [JGU09]. multi [AMS05, BB06, EO09, HR06, RSS^{+}07]. multi-attribute [EO09]. multi-criteria [AMS05]. multi-disciplinary [RSS^{+}07]. multi-species [BB06]. multi-stakeholder-driven [HR06]. Multiple [Mor06, BET^{+}07, Var08]. multiple-use [BET^{+}07]. multispecies [SHQF06]. Murky [CCCD07, Dro04, Mau02]. Museum [Pot00]. mussel [MGL09, WCB06]. Mutual [Ben01]. myths [Mur07].

Nadu [BdKvD^{+}08]. NAFTA [JM00]. narrative [Joh06]. nation [HRP08, Kan03]. nation-states [HRP08]. National [Sah01, KYC05, Per09, TRvAB05, Pot00]. nationalism [dVM02a]. nationals [Erc06]. Nations [Ver07, TRvAB05, Hea08]. native [RM00]. Natura [PFS09]. Natural [MMS09, HK06b, MR01, Rep01, DP00]. nature [BET^{+}07, vZvDT02]. Nautical [Vee02]. naval [Ber02, Ska05]. navigate [EYG^{+}09]. navigating [DSH03]. navigation [VA06]. navigational [Van05]. NEAFC [Bjø09]. need [Day08, KKP00, Son05]. needed [WAKP05]. needs [WJPN00]. negotiations [Gry03, Sco00]. Neo [vdB00]. Neo-classical [vdB00]. Nephrops [CFG06]. net [KM09]. net-trawling [KM09]. Netherlands [Vee02]. nets [GL07]. Network [Gib08, JC09b]. networks [SBB^{+}09]. Newfoundland [FE09, AHS00, BBNB08, Sch05]. nga [JGU09]. Niño [TLHN07]. no [Jon06, Jon08, Jon09]. no-take [Jon06, Jon08, Jon09]. nomenclature [Ser04]. non [Mat01]. non-technical [Mat01]. Norfolk [MBSLP02]. Normalization [Ili08d]. North [MBSLP02, AL02, Bjø09, CFG05, CS07, DJ07, DW08, GGVLIM03, GGVL05, HRP08, Kim03b, Lev09a, Lev09b, McB04, OB09, Skj06, Tsu04, VVG03]. North-East [Skj06]. North/South [VVG03]. northeast [Han07b, HR06, KdSR07, Sto09, dSK06]. Northeastern [JM00, SWC09]. northern [HJ02, Val03]. Northwest [Haa00a, CC09, DB06, Ebb02, Ebb09, HR06]. Northwestern [GL03]. Norway [ÅB02, AJ09, Arm01, AvdH08, AHS00, Gez08, SA08]. Norwegian [EOU09, FH04, Gez06, HHR00, JM04, Joh05, MJ03, SB05, SA02, SS07, Til08, Utn07]. note [Pon08]. Nova [CWP06, Lou07]. November [Ano00f, Ano02g, Ano03k, Ano04l, Ano05k, Ano06n, Ano07j, Ano08j, DE07]. nuclear [WDAT00]. number [dVMdC08]. Nuts [Sta05]. NW [FGA00, NVPR09, Val00, WPBD03, Haa00b].

objective [PPW^{+}09]. Objectives [MPB^{+}02, Hil07]. obligations [FH04]. observation [KP04b]. Observations [AOPM00, Ham03]. Observing [Mal03]. Occupational [MN09, WNB^{+}08]. Ocean [DB06, Edw08, Mal03, Val01, Car07, Cho06, CW08, CSH00, Edw09, EYG^{+}09, KP04b, PF02, RHCS05, Sah01, SP01, dVM04, Agn00, BC07a, Cam00, CWRR09, HRP08, Hun06, Pol06, RSJ^{+}03]. oceanic [KS09]. Oceans [Hin03, Vin06, FHCS05, RM00, Smi03, VH09, dV07]. off [AvdH08, Utn08b, Jab09]. offences [GL07]. officers [GM01a, GM01b]. Offshore [DJ07, BBNB08, CWP06, FE09, Gez06, Ham03, KP04a, Ped08, PFS09, Por09, RM00]. oil [FW08, FE09, FFM06, GGSRVL06, HS08, Kim03a, LLJ07, Mas03, NVPR09, VG03, Kim02]. Oman [AOPM00]. One [Ser04]. Online [Ano07g]. OPA [HS08]. Open [Llá03, FMPPF08, Cor05]. operating [Nob03, Van04]. operation [GT01]. Operational [Xiz04]. operations [Bal04, KM09, Lla05b, MWV02]. Opportunities [Har01, BBNB08, DJ07, KS04, Mit08, Sah01, Tau07]. option [Hin01, SLT08]. Options [Dja04, PPP^{+}08, RT08]. Oregon [HESM08]. organisation [Syd01]. organisational [Mit04]. organisations [Syd02]. organization [Hin01]. organizations [DM09, LOS06, SJ03, Yan03]. orientation [Tro04]. oriented [TMY06b]. origins [Bod02]. other [Ser04]. outcomes [WNB^{+}08]. outcompete [ENS04]. output [KYC05, RNG06]. Over-capacity [AEG^{+}08]. over-represented [Oke03]. overfishing [Pon08, TS07]. overflight [VA06]. overlaps [Bat05, EYG^{+}09]. Overview [Bjø09, ÅB02, HS01, Jen08, TTSS09]. owned [FLT06]. ownership [HFPR02]. owning [Mit04].

P&I [Ben01]. Pacific [BC07a, Cam00, Haa00a, HRP08, Val00, VS01, BPM09, BKV^{+}09, Ebb02, Ebb09, Haa00b, HT09, Han08, Hun03, McB04, Ove06, Ove07, Pet02, PP04, RBT04, RSJ^{+}03, Tsu04, VS01, WW06]. padu} [LB04]. Pages [Ano00b, Ano00d, Ano00g, Ano00c, Ano00f, Ano00e, Ano01b, Ano01d, Ano01f, Ano01c, Ano01e, Ano02c, Ano02e, Ano02h, Ano02d, Ano02g, Ano02f, Ano03h, Ano03l, Ano03i, Ano03k, Ano03j, Ano04h, Ano04j, Ano04m, Ano04i, Ano04l, Ano04k, Ano05g, Ano05i, Ano05h, Ano05k, Ano05j, Ano06l, Ano06j, Ano06k, Ano06n, Ano06m, Ano07i, Ano07k, Ano07h, Ano07j, Ano08g, Ano08i, Ano08h, Ano08j, Ano09g, Ano09j, Ano09h, Ano09i, Ano09k]. Pagrus [MS02, MBJI08]. paid [Jab09]. Painting [DGH^{+}06]. Palawan [PPP^{+}08]. Panama [Llá05a, Lla05b, Lla06]. Panamanian [Llá04]. Paper [dV07]. paradox [dVMdC08]. parallel [Car07]. Paris [DE07]. Participation [Alp06, BWR08, GH08, Hol00, MS02, MJ03, VGH^{+}08, dVMdC08]. Participatory [MJ08, SBF^{+}04, WBCK04, AMS05, BL08, WCKB09]. particular [WJPN00]. partnership [CW08, SM06, FBH^{+}07]. partnerships [Fle03]. party [Mit04]. passenger [Tza05]. past [Llá03, vS02, CvSP06]. Path [HHR00]. Patos [CVL02, KS04, SR08]. patrolling [Let00]. pattern [PL01, SB05]. Patterns [Lou07]. pelagic [GCB^{+}08, SRMM03]. Pentland [SDWF09]. people [Bes01, CS07]. peoples’ [DJ01]. Perceived [MCK08]. perception [DCHM07, MBSLP02]. perceptions [GDI^{+}09, HHM^{+}09, MG09]. Performance [Fah08, MPT09, Ben01, Bjø09, Fra08, LDBB08, MGL09, NZF09]. perils [MR01, Yan08]. period [TK09]. permit [NZF09]. Personal [WR07, BL00]. perspective [FÓ08, FP07, Hea04, HH06, HH00, Kim04, kK05, LDBB08, MFM09, MN06, Pla08, San05, Sha05, SLCS05, Ugg07, Val01, WR07, Xiz04, dVM06, XX05]. perspectives [CLT08, Eva09, Jon08, NSVW01, TTSS09, YY08, Ebb09, QCB09, SGP^{+}08]. perturb [Ver07]. Peruvian [Pon01]. petroleum [CWP06]. Pew [Cor05]. Phang [JGU09]. Phang-nga [JGU09]. phenomenon [BPM09]. Philippines [Fab08, Maj07, MPT09, PPP^{+}08, Sub07]. Phocoena [JBP09]. Phronetic} [Jen06]. pilots [DCHM07]. piracy [Ho09a]. plan [FHCS05, GGVL09, Fra08, Vin06]. planes [Van04]. Planning [BKV^{+}09, DPE^{+}08, DE07, MM04, AGPT08, BHJV09, CN08, Dav08, Day08, DW08, DMVS07, Dou08, Ehl08, FÓ08, Foc08, GL08, Mae08, OGC^{+}09, Pla08, PD08, RHCS05, SBF^{+}04, SHA08, Tau07, Tow02, dVMdC09, vZvDT02]. plans [GA07]. plutonium [WDAT00]. polar [MLRK08]. Policies [EY07, Car07, DJ07, HH04, KF02, McG07, PLB04, SM08, VG03, dVMdC09, vdBND^{+}02]. Policy [ABF00b, Hin01, AS07, AH06, AvdH08, Bee07, Bod02, CC09, DG05, DPE^{+}08, Dro04, DP00, Fle00, GMN^{+}07, HY03, Hat00, Hea04, HH00, HR04, Kea01, Lev09a, LFL05a, LFL05b, LTX08, MR01, McG03b, MWL^{+}09, NSVW01, OLC^{+}09, PL08, Pet02, Pet08, PNN08, Pla08, PPP^{+}08, PNT09, PP04, RM00, Sah01, SM07, SWC09, SBB^{+}09, SC04, SLCS05, SRVLIM03, SR05, TRvAB05, Val09, WRD06, Wit06, ZYN09, vHvT09, Ano07g, DG05, FA06, GH03, LFL05b, MG09, QCB09, Vin06, dV07]. policy-led [OLC^{+}09]. policy-making [TRvAB05]. policymakers [Gib09b]. Political [AJ09, TS05, HB01, dV07]. politically [RTWJ05]. politics [Mau02, Ped08]. Pollution [Kim02, FW08, Kim03a, KF09, LL06, Mas03]. pool [HRP08, Jen02]. poor [BHJV09, SLT08]. population [DSH03]. populations [Hun06]. porpoise [JBP09]. Port [CLT08, Dev00, Lla06, CMW09, CMCNL^{+}09, KF07]. ports [But07, PLB04, Pet08, TS05]. Portugal [Car07]. posed [Jon06]. Positioning [GHR^{+}08]. possession [Edw08]. possible [Ili08a, LC06, MCK08]. post [Ili08b, Val01, vSBHS06, vSH06, Daw06, TK09]. post-apartheid [vSBHS06, vSH06]. post-IFQ [Daw06]. post-ISM [TK09]. post-modern [Val01]. pot [Lev09a]. Potential [SVL07, AEG^{+}08, Edw09, JBN07, KP04b, KE03, Pot00]. power [Jon09]. Poyil [TG09]. practical [WRD06, YY08]. Practice [OLC^{+}09, AH06, Alp06, Arm01, Day08, KKI^{+}09]. practices [Hea08, MJ08]. pre [TK09]. pre- [TK09]. Precaution [SC04, Hea06]. Precautionary [HF02, Par00, GL05, Hea06, Pol02a]. precedes [CJ07]. predatory [Pol06]. Predicting [ZYN09]. prefecture [AMS05]. preferences [PPW^{+}09, WP09]. preliminary [EOU09, Hun09]. present [CvSP06, Llá03]. pressure [PPB06]. Prestige [FFM06, GGSRVL06, NVPR09]. prevent [VG03]. prevention [MN09]. prey [HRP08]. prices [Gui04]. Prince [Cho07]. principal [Foc08]. principle [GL05, MCO08, Mor00, Tsu04]. principles [AH06, MCO08, Sch07c, TMY06a, Utn06]. priorities [SDWF09]. Private [PKH02, BT04, JM00]. privatization [Lla06]. Privatizing [Che03]. probability [KF07]. problem [CFG06, Gez08, HRP08, JC09a]. problems [BW09, Ebb03, Jon06, Key01, Ser04]. process [GL08, PD08, RHCS05, WBGC08, dVMdC08]. producing [TRvAB05]. production [BP08, KA06, LS08, MGL09, WCB06, vZvDT02]. productivity [FP04]. products [EY07, GP00]. profile [SWM06]. Profitability [SRVLGG02, BRS08, FH04, GGVL05, GGAD08, WJPN00]. profitable [GGVL09]. program [CWP06, GTLC02, HESM08, Kai06, PBSS06, Ber08, KK08]. Programme [Ho09b, VS01]. programs [AMW09, MPT09, PBSS06]. progress [FPS05, PKS^{+}09, VA05b]. project [CC09, MM04]. projects [BWR08]. promise [DJ01, Yan08]. promote [BBNB08]. property [BH00, BM06, HG01, Kea01, Mar01, PKH02, RHKS09, SMS06, dS04]. proposal [RTWJ05]. proposals [DTFÁIM04, Jon08, Jon09]. proposed [BET^{+}07, MM04]. Proposing [CWP06]. prospect [Hal03]. Prospects [Cli03, Haa00b, Haa00a, Val00, Key01, KPN^{+}09, WCB06]. protect [Det06]. protected [BCM09a, Cli03, CCCD07, Fab08, Jon08, Jon09, NBR09, Ove06, QWJA09, RGB09, SBF^{+}04]. Protecting [JHW09, Skj06]. protection [AvdH08, BR07, BET^{+}07, GTLC02, Ugg07, Ver07]. providers [CLT08]. province [PPP^{+}08, JGU09]. Provinces [Bod02]. provisions [MG06, Ver07]. PSSA [Det06, RTWJ05, Ugg07]. Public [FPHP09, JM00, KS09, MBSLP02, AS07, BT04, MJ01, Nau06, PKH02, Por09, WP09, dVMdC08]. Publicly [Eks09]. Pulau [TS07]. Putting [AH06, Glo03].

quality [EO09, TRF08]. Quantitative [NZF09, GP00]. Quebec [DM09]. quota [AVH05, Bes05, Bes06, HHR00, Mor00, SHQF06, VEH02, HFPR02]. quota-hopping [Mor00, HFPR02]. quotas [ANL09, AEG^{+}08, BRH06, CBB00, Gib07, GM09, PG09, PE09, SHQF06, TMY06a, Tro04]. radical [AJ09]. radioactive [WDAT00]. raft [MGL09]. raised [Jon08, Jon09]. Raja [BRS08]. range [MR01]. rapid [Pol06]. rates [Pol06]. ratification [KF09]. Rationalization [Fin05]. re [JBN07]. re-creation [JBN07]. realignment [MM04, MBSLP02]. realities [PBB08]. reality [GH03, Joh02]. realizing [KE03]. ReCAAP [Ho09a]. Reconciliation [RM00]. reconstruction [Hin01]. Recovering [Ho09b, ZFP05]. Recovery [Ho09b, GGVL09]. recreation [OGC^{+}09]. recreational [AMW09, JHMC07, Kea01, MBJI08, PGP08, PL01, SJ03]. recreational/commercial [Kea01]. recycling [LLJ07]. red [CGS08, Ser04]. redfeed [Til08]. redistribution [HH00]. reducing [HS08, NZF09, PPB06, LL06]. Reduction [Jen02, PC00]. Reef [KK08, NBR09, JGU09, Kai06, PPB06, PPP^{+}08, TTSS09, Day08]. Reefs [Sub07, PPB06]. reference [PFS09, WJPN00]. reform [GH03, HK06a, Isa06, MJ03, Wit06, vSBHS06, Ano06g]. Reforming [HB07a, SLCS05]. reforms [Har09, vS02]. regime [Ebb02, HB07b, HNMB07, Ili08b, Mas03, RBT04, RNG06, Ska05, SA08, SC04, Sto07, Val04a, Yan03, Yan08]. regimes [Ebb03, Fra08, Kim03a, Mar01, Nie06, Syd01, WNB^{+}08]. region [DM09, Haa00b, Haa00a, HT09, Val00, WW06, DP00]. Regional [HMM^{+}04, Syd02, Vin06, DS09, FHCS05, HB07b, Hin01, Syd01, WKT00, dVM02a, Oke03]. regions [BHJV09, McG07, Syd02, dVMdC09]. registers [Llá03]. Registration [WW06]. Regueiro [SR05, LFL05b]. Reguiero [LFL05a]. Regulating [BBB01, Kim03a]. Regulation [Ili08c, AJ09, AW02, EE03, HR04, RC01]. regulations [Eks09, LL06, RHKS09, Sta03, Sta06]. regulatory [Fah08, HB07b, MN08, WNB^{+}08, WDAT00]. Rehabilitating [PRH^{+}06]. related [Jon08]. relations [Kan03, MLRK08]. relationship [GH08, KA06]. relative [Mor00]. reliability [MWV02, Tza05]. reliability-centred [MWV02]. remember [vS02]. removal [KP04a]. Renaming [JP08]. Renewable [PF02, Por09, SDWF09]. rent [AEG^{+}08, SLCS05]. repeat [vS02]. Replacing [BRH06]. Reply [Haa00a, SR05]. Report [VA05b, Cor05]. reporting [CCCD07]. representation [Fle03]. representative [JC09b]. represented [Oke03]. Republic [Kim04]. requiem [CM00]. requirement [Tow02]. rescue [Ber08]. Research [Ber08, Bat05, BWR08, CC09, Glo03, Har01, HR06, KM04, KTH^{+}07, Mau02, PL08, SBB^{+}09, Ver07, WBCK04, WCKB09, TLHN07]. Reservarum [RZ03]. reserve [BPM09, dS04, PBB08]. reserves [GK05]. Resilience [Gib09b, Ebb09, MLRK08]. resolution [BL00]. Resolving [CFG06, Dja04, RWB08]. Resource [DP00, Maj07, BA05, MR01, MPT09, Mur07, NH07, SJ02, SLCS05]. resources [Bes01, CWRR09, HK06b, HRP08, Hun03, JM00, LVH^{+}09, Mor00]. response [Hum03, Xue06]. responses [BM07, Gez08, PS07]. responsibilities [NC06, Xue06]. responsibility [FLT06, Gos08]. responsible [MJ08]. restoration [BM09]. restoring [Nau06]. Restructuring [LTX08]. result [Nie09, Pol06]. Rethinking [Hau08, PP04, Tsu04]. returns [HCL^{+}09]. revenues [Edw09]. review [CW08, LVH^{+}09, MWL^{+}09, Sow06, WJPN00, Lev08]. revision [Hew07]. revolution [Gui04]. rhetoric [GH03]. Rica [Alp06]. Right [ET08, Lev09a, Lev09b, SMS06, Van05]. rights [BA05, BH00, BR07, BT04, BM06, DJ01, GT01, Hea08, HS06, Kea01, Lax06, RM00]. rights-based [BA05, GT01]. RIO [Hin01]. rise [JP07, Sta06, Sta08]. rising [dVMdC08]. Risk [Sag05, Ben01, BM09, DCHM07, HCL^{+}09, TS05, WBGC08]. River [PBSS06]. Roanoke [PBSS06]. robbery [Ho09a]. rock [PKH02]. roe [BP08, Ska04]. Role [FFM06, HHM^{+}09, BH00, Bes06, DMVS07, Elv09, KYC05, Mah06, Mas08, MHJ07, Mit04, MWL^{+}09, PG09, Ska05, SBB^{+}09, Tro04, VGH^{+}08, VS01]. roles [Bee07, Bjø09]. room [PE09]. Rotterdam [HHM^{+}09]. rough [Ber02]. rule [JM04, NM03]. rule-making [JM04]. rules [Ham03, Høn00]. Russia [VL02, HJ02]. Russian [AVH05, Ska04].

sablefish [SVL07]. safe [BdKvD^{+}08, Bro05]. safe-guarding [BdKvD^{+}08]. Safety [DCHM07, KKP00, Ben01, Cho07, KF08, KF09, MN09, PL08, WF01, WNB^{+}08, PNN08]. Sagami [MBJI08]. sailing [Ber02]. salmon [BR08, ENS04, Ebb02, Ebb03, EOU09, Elv09, Gui04, Hea08, HRP08, LS08, WCB06]. Saltwater [SWC09]. sardine [HNMB07]. scale [AE01, DW08, EY07, Hau08, HCL^{+}09, IMP09, Isa06, Joh06, OJ08, PNT09, SGP^{+}08, Sow06, WD07, WASK06]. scallop [BA05]. scarce [HK06b]. scenario [dC08]. scheme [Agn00, BET^{+}07]. Science [WKT00, CW08, DG05, DM09, Hea04, KM04, LFI01, SBB^{+}09, Tow02, Wes03, ZFP05, Ber08]. science-policy [SBB^{+}09]. Scientific [Hea04, Hum03, Bat05, EY07, FFM06, Glo03, LVH^{+}09, Nau06, Pon03, Ver07]. scientist [San05]. scope [Ano02a, Ano03d, Ano03e, Ano03f, Ano04b, Ano04a, Ano04c, Ano04d, Ano04e, Ano05a, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano06a, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano07a, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Mas03, SAK06, dVMdC09]. Scotia [CWP06, Lou07]. Scotian [RHCS05, FHCS05]. Scotland [SDWF09, JM00, MM04, Nob03]. Sea [Ber08, Cho07, KTH^{+}07, BA05, CSG^{+}09, DMVS07, Dou08, FLT06, GTLC02, HF02, Kan03, KKP00, Kim03b, Lev09a, MWL^{+}09, NZF09, PM02, PLB04, Roa04, RTWJ05, RS03, Val03, dVM02b, CFG05, DW08, Fin05, Høn00, IMP09, Key03, KTH^{+}07, SM08, hS00, Son05, Sto07, Ugg07, VVG03, Ver07, dVM02a]. Sea. [BET^{+}07]. seabird [MWL^{+}09]. seabirds [WBGC08]. seafarers [GMN^{+}07, Gle08, Mit08, PGM^{+}05, WW05]. seafaring [GM01a, GM01b]. Seafood [Kap00, BH00, Bes06, GP00, Gui04, JP07, JP08, LTX08]. Searching [WCB06]. seas [AGPT08, Ber02, Kay04, Ove07, Ugg07, dV07]. seaside [FÓ08]. season [BdKvD^{+}08, KK08]. seasonal [Tro04]. secrets [JP08]. sector [Gib09a, LFI01, LDBB08, Llá04, PL01, SJ03, VS01, VES06, dVM02b]. security [BKV^{+}09, Kin05, Roa04, dVM04]. security-based [dVM04]. seismic [CGHA08]. selection [CLT08, KP04a]. selling [Kap00]. sensitive [RTWJ05]. September [Ano00e, Ano01e, Ano02f, Ano03j, Ano04k, Ano05j, Ano06m, Ano09k]. sequestration [KS09]. SERMP [FHCS05]. Service [MN08, CLT08]. services [BM09]. set [HY03]. seven [GM09]. severance [KP04a]. shaping [MWL^{+}09]. shared [SRMM03]. Sharing [Ho09a, NC06, vdBL07]. Shark [GCB^{+}08, Lev08, DKM08, MBJI08]. Shelf [RHCS05, FHCS05, Jab06, Jab09, Key01]. shellfish [FMPPF08, HR04]. Shetland [CW00]. shifts [HNMB07]. ship [BK09, But07, KF08, KKR08, Mit04]. shipping [CLyB06, CLT08, Chi07, FLT06, Gos08, HY03, Kin05, KF09, MN06, PM02, RM02, TK09]. ships [KF07, Moe08, Som01, Van04]. Shiretoko [MMS09]. shore [PGM^{+}05]. shore-based [PGM^{+}05]. short [FLT06, LDBB08, PM02]. short- [LDBB08]. Shrimp [NSVW01, CGS08, KPN^{+}09, Sta03]. Sicily [WPBD03]. side [McG03a, Pon01]. SIDS [SM08]. Sierra [TWN^{+}09]. simulation [KPN^{+}09]. since [Han07b]. Sino [Key03]. Sino-Japanese [Key03]. Site [Sub07, NBR09]. sites [PFS09]. situation [PGM^{+}05]. six [WNB^{+}08]. sizes [OJ08]. Slow [LWH^{+}09]. Small [Isa06, PNT09, SRMM03, AE01, EY07, FÓ08, HB07a, Hau08, IMP09, JHW09, Joh06, OJ08, Pol02b, PS01, SGP^{+}08, Sow06, SWM06, SW08, WD07]. Small-scale [Isa06, PNT09, AE01, EY07, Hau08, IMP09, Joh06, OJ08, SGP^{+}08, Sow06, WD07]. smallest [Utn08b]. Smart [HR04]. SNA1 [MS02]. snapper [MS02]. snow [Lou07]. Social [Gos08, WP09, FLT06, KM04, Ove06, Ove07, PBB08, San05, dV07]. society [MHJ07, Val01, Cor05]. socio [TTSS09]. socio-economic [TTSS09]. sociocultural [Hol00]. Socioeconomic [FGA00, SBF^{+}04, BHJV09]. soft [SA02]. Sole [DTFÁIM04]. Solent [Con07, FJH07, FP07, FBH^{+}07, Hew07]. solution [PPB06, SMS06, Sta07]. solutions [CFG05, LC06]. solve [Pol02b]. some [Ham03, Ser04, WAKP05]. Sound [PBSS06]. Sound/Roanoke [PBSS06]. South [BJBC09, GDI^{+}09, FHCS05, LB04, Ano06g, BC06, CCCD07, CvSP06, DPE^{+}08, Har09, HS01, HK06a, HH00, NDV^{+}04, NH06, hS00, Son05, Sow06, Syd01, VL02, VVG03, VS01, Vin06, Wit06, vS02, vSBHS06, vSH06]. Southampton [But07]. Southeast [PPP^{+}07, SGP^{+}08]. southeastern [Lev09a]. Southern [JGU09, BP08, CBB00, Cla06, DKM08, How01, Agn00, BM09, GDI^{+}09, Hun06, Lou07, NDV^{+}04, Pol02a]. Soviet [VEH02, VES06]. Spain [FMPPF08, FGA00, NVPR09, dC08, GGVLIM03, MPB^{+}02, PL01, dVM02b]. Spanish [CMCNL^{+}09, FFM06, GGVL05, GGVL09, Kap00, Lax06, MGL09]. Spatial [BR07, DW08, DE07, Val09, AGPT08, CN08, DMVS07, Dou08, Ehl08, FÓ08, Foc08, GL08, Mae08, OGC^{+}09, Pla08, PD08, SHA08, SRW06, Tau07, dVMdC09]. spawning [SB05]. Special [SP01, Oke03, PFS09]. Specialist [RSS^{+}07]. species [BWA^{+}03, Ber08, BB06, Pol06, RWB08, Ser04, SRMM03]. Spending [PL01]. spill [FFM06, GGSRVL06, NVPR09]. spillover [RHKS09]. spills [HS08, VG03]. sport [SWC09]. spring [SB05]. spring-spawning [SB05]. squid [Har09]. St [BP08, Lou07]. stability [Mor00]. stable [SA08]. Stakeholder [BWR08, Fle03, HCK00, PPW^{+}09, FP07, HR06, VGH^{+}08, WP09, Yan03, BCM09b, GDI^{+}09, GH08]. stakeholders [MJ01, MJ08, PD08, SGP^{+}08, Som03]. stand [Gui04, hS00]. state [CMW09, Con07, Dev00, Eks09, Erc06, HB07b, KF07, Xue06, Fle00, dVMdC09]. States [JM00, HRP08, SRMM03, TWN^{+}09, AL02, BM08, MN08, Oke03, RBT04, WK09]. statistical [PSFE07, SWC09]. statistics [OJ08, vZvDT02]. stay [PL04]. Steller [Ber08, GTLC02]. Step [CJ07]. steps [GL08, LWH^{+}09]. Stewardship [Gul09]. stock [MS02, MBJI08, PBSS06]. stocking [PBSS06]. stocks [BC06, FSPKR09, Hun06, Mil07, SS07, TMY06b]. stocktake [OB09]. straddling [TMY06b]. strait [Son05, MLRK08, Val03]. Straits [CLyB06, Pla00]. Strategic [Lan08, SDWF09, Tow02]. strategies [Bod02, LLJ07, MCK08, TRvAB05]. strategy [Cam00]. Strengths [PM02]. striata [Lev09a]. Structural [Hol00, Hat00, LFL05a, LFL05b, SRVLGG02, SRVLIM03, SR05]. structure [CWP06, MJ03, Sta07]. structures [KP04a, LS00]. Studies [VS01, AEG^{+}08, MPB^{+}02, MLRK08]. study [Arm01, BCM09b, But07, CFG06, CW00, Dav08, DG05, DP00, EOU09, Fra08, GM09, How01, Lev08, MWV02, OJ08, PBSS06, PSFE07, RGB09, SR08, Som03, SW08, SC04, WW05, dJDT09]. studying [Nie03]. Submarine [CSH00]. submission [Ano07g]. Subsidies [Hat00, AS07, Gry03]. Subsistence [SM07, Sow06]. substandard [KF07]. success [BCM09a, Hil07, Kim02, Lou07, PKH01, WK09]. successful [Yan03]. Sultanate [AOPM00]. Summary [Val04c, DE07]. Sundarbans [KPN^{+}09]. Supply [Nie06, Pon01, MN08, Pon03]. support [DM09, Nau06, RSS^{+}07]. supporting [WBCK04]. Surís [SR05, LFL05a, LFL05b]. Surís-Regueiro [SR05, LFL05b]. Suris-Reguiero [LFL05a]. survey [Nøs08, WP09]. survey-based [WP09]. surveying [Bat05]. surveys [CGHA08]. sustainability [AMS05, BC07a, BA05, BB06, DG05, Gai08, GJGT04, SU07, Utn07, Utn08a, Utn08b, WPBD03, WCB06, Wit08]. Sustainable [BR08, Edw09, LOT05, Gib09a, Glo03, Gry03, HS06, Joh02, KKI^{+}09, Ove06, PBB08, SM08, Sta05, Tho08, Utn08b, AH06]. Svalbard [Ped08]. sweat [Mar01]. Swedish [BL08, EE03, Ugg07]. Swimming [Ebb03]. swordfish [GGVLIM03, GGVL05, NZF09]. synergies [Skj06]. synthesis [BWR08]. System [Mal03, Utn07, AMS05, Bes05, Bes06, CMCNL^{+}09, Cho06, GH03, HHR00, LOT05, LB04, NZF09, QWJA09, RNG06, SU07, Til08, VEH02, VES06, WCB06, dlM05]. Systematic [BHJV09]. Systems [Utn06, Ebb09, GT01, IMGGVL02, KP04b, LW07, MMR08, RNG06].

T. [HHR00]. TAC [RNG06]. TACs [Mor00]. tags [JHMC07]. Taiwan [CLyB06, Chi07, HB00, LOT05, hS00]. take [Jon06, Jon08, Jon09]. Tamil [BdKvD^{+}08]. tanker [Pla00]. Tanzania [How01]. target [CMW09]. targeting [KF07]. Tasmanian [PKH02]. taxonomy [Ser04]. Team [RSS^{+}07]. Technical [Tza05, BB06, Mat01, PC00, DGH^{+}06]. technological [Sta05]. Technologies [SJ02, SHA08]. Technology [GL03, Bal04, CW08, DM09, Var08]. temperature [Han07b]. Ten [Kim02, Mur07, Sch05, Sow06, vSH06]. ten-year [Sow06, vSH06]. tendencies [Che03]. tension [vHvT09]. tensions [DJ07]. term [LDBB08]. terminology [Ska05]. terms [Hay05]. territorial [Val03]. territory [HB07a, Gre01]. terrorist [Ho09b]. Testing [Det06]. Thai [JGU09]. Thailand [JGU09]. their [Bes01, BC06, DYB05, DTFÁIM04, Hun09, Pot00, SJ02]. Theoretic [YY08]. theoretical [VG03]. Theory [Arm01, Alp06, HY03]. there [CW00, KTH^{+}07, KA06, Sch07b]. third [Mit04]. Those [vS02]. threat [BWA^{+}03]. threatened [Ber08]. threats [Hun09]. three [McG07, SGP^{+}08]. tilefish [KdSR07]. Time [BdKvD^{+}08]. Time-zoning [BdKvD^{+}08]. Tingloy [Maj07]. title [RM00]. Tobago [BC07b, Som03]. Tokyo [KKI^{+}09, VA05b]. tonnes [Dro04]. tool [EYG^{+}09, HCL^{+}09]. toothfish [Agn00]. top [Bee07]. tourism [Che03, Fab08, HH06, Joh02]. tourist [Edw09]. toxic [Moe08]. Trade [GP00, Ho09b, JP08, Sto09, Elv09, Nie09, PPP^{+}08, Utn08b]. trade-off [Utn08b]. trading [Pet03]. traditional [SR08]. traffic [Lla05b, Pla00]. tragedy [SMS06, SA02]. training [How01]. transaction [MM05]. Transferability [Lax06, BT04]. transferable [BRH06, CBB00, Gib07, GM09, PE09, TMY06a]. transfers [MM02]. Transformation [NH06, Har09]. transition [VES06]. Transitions [BM06]. Transparency [FE09]. transportation [WDAT00]. trawl [WPBD03]. trawlers [FH04, Sta08]. trawling [KM09, Sta03, dJDT09]. Trends [Mar04]. tri [Per09]. tri-national [Per09]. trials [Nob03]. tribulations [Nob03]. Trinidad [BC07b, Som03]. trip [BRH06]. tropical [BNA^{+}01, CFW^{+}07, Mil07, SH03]. tropics [WRD06]. troubled [RM00]. trunk [CLT08]. tsunami [ZYN09]. Tubbataha [Sub07]. Tuna [Pol06, BC07a, Bro05, CBB00, Cam00, HB00, HB01, LWH^{+}09, Mil07, RBT04, RSJ^{+}03, Sco00, Ser04, WKT00, Pol02a]. Tunas [Lev08, SH03]. turbulent [VH09]. Turkish [Pla00]. turtle [CSG^{+}09, MWL^{+}09]. turtles [NZF09]. two [Arm01, Ebb03, LOS06, McG03a, Pon03, Ser04].

UK [SDWF09, Bro05, Fle03, FPHP09, GM01a, GM01b, Gle08, HFPR02, JC09b, MPB^{+}02, MBSLP02, OJ08, PGM^{+}05, Pot00, SJ02, WR07, WJPN00]. Uncertainty [GK05, Jen08, Hea04, Pon01, Pon03]. Understanding [Eva09, TLHN07, BNA^{+}01, BBNB08, Mor06]. underutilized [HESM08]. UNESCO [DE07, Sub07]. Unintended [RHKS09]. Union [DG05, LFL05a, LFL05b, SR05, ANL09, CW00, FA06, HBB00, MN08, SRVLGG02, SRVLIM03, Wit08]. Union. [Mor00]. unique [Det06]. United [JM00, MN08, BM08, Kan03, Oke03, Pet08, Ver07, WK09]. university [KKI^{+}09, VS01]. university-based [KKI^{+}09]. unravelling [GL07]. unregulated [Agn00, LC06]. Unreported [BJBC09, LC06]. update [KM09, Pla00]. upon [AOPM00]. upstream [Ebb03]. USA [Cla09, KdSR07, MWL^{+}09]. usage [vZvDT02]. Use [BT04, BKV^{+}09, BET^{+}07, DMVS07, Dou08, FD03, Hea04, Let00, Maj07, Mit04, Ove06, RSS^{+}07, WP09]. used [CMW09]. user [Edw09, Hol00, MJ01, MHJ07, WR07]. user-group [Hol00]. user-groups [MJ01, MHJ07]. using [Gib07]. Utila [KHG08].

V [Cho07]. Validity [Val03]. value [Joh06, KP04b, SLT08]. values [Sub07]. Valuing [EO09]. Varela [LFL05a, SR05, LFL05b]. Varela-Lafuente [LFL05a, SR05, LFL05b]. variability [McG07, Mil07, PS07]. variations [Han07b]. various [KF08]. Veil [GL07]. vendace [RC01]. vents [Glo03]. versus [CSG^{+}09, MBJI08, OLC^{+}09]. Vertical [Daw06]. vessel [AEG^{+}08, FH04, HHR00]. vessels [ET08, HFPR02, LL06, LLJ07, RT08, Utn08b, Sag05]. viability [BC07a]. viable [SLT08]. vicious [JM04]. Vietnam [PNT09, RSS^{+}07, vZvDT02]. view [Gle08]. viewpoints [TS05]. views [BCM09b]. virtual [FD03]. vision [dV07]. volume [Ano06i, Ano01a, Ano02b, Ano03g, Ano04g, Ano06h]. voluntary [SM06].

walrus [MLRK08]. wars [PPP^{+}07]. warship [Dro04]. waste [But07, Hea08, LLJ07]. wastes [WDAT00]. water [AvdH08, EO09, HB00, HB01, Hew07, Xue06]. waterborne [Mar08]. Watercraft [WR07, BL00]. waters [BM07, BL00, CCCD07, Dro04, ET08, FTF00, Mau02, RM00, Sag05, Som01, SS07, VH09]. way [Val04a, Val05b]. weaknesses [PM02]. web [Fle01, KHG08]. web-based [Fle01]. welfare [Nie09]. West [BJBC09, MBSLP02, OB09, PS07, VVG03, KF02]. Western [PPP^{+}08, Lev09a, RSJ^{+}03, BC07a, Cam00, CWRR09, Det06, RTWJ05]. whale [Lev09a, Lev09b]. whales [ET08]. Whaling [BH09, Hea04, Hea06, Ili08d, ABF00a, ABF00b, CCG^{+}07, EY07, GLP08, HH06, Ili08b, Mor06, Ili08c]. whether [Har09]. whiting [Lev09b, Sta06]. who [KF02, vS02]. wicked [JC09a]. wider [CS07]. wild [BR08]. Will [Pol02b, Bra08, CW00, FSPKR09]. wind [MCK08]. within [ANL09, HHR00, MB09]. without [Eyt00, HH00]. women [FMPPF08]. word [SM07]. Work [vdBL07]. working [PNN08]. Workshop [DE07]. World [MMS09, NBR09, Sub07, GHR^{+}08, LWH^{+}09, Mar04, TRvAB05, Vee02]. worldwide [CGHA08]. wreck [FTF00]. WSSD [Nau06]. WTO [Gry03].

year [Sow06, vSH06]. years [Kim02, Sch05]. Yellow [VVG03]. Yemen [WD07]. Yoron [AMS05]. youth [BH09].

Zealand [BJBC09, MS02, BH00, Bes01, Bes05, Bes06, BR07, DCHM07, Dev00, Har07, Kea01, RWB09, SWM06, SW08, VH09, Yan03, Yan08]. zero [CJ07]. Zone [MCO08, Coo05, Han08, Key01, PFS09, Sch08, VA06, Van04, Van05]. zones [Jon06, Let00, Som01, Val05a]. zoning [BdKvD^{+}08, BET^{+}07, DPE^{+}08, DB06, Edw08].

References

Aarland:2002:FMN

[ÅB02] Kristin Årland and Trond Bjørndal. Fisheries management in Norway — an overview. Marine Policy, 26(4):307–313, July 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000131.

Aron:2000:WI

[ABF00a] William Aron, William Burke, and Milton M. R. Freeman. The whaling issue. Marine Policy, 24(3):179–191, May 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X99000317.

Aron:2000:WIM

[ABF00b] William Aron, William Burke, and Milton M. R. Freeman. The whaling issue: [Marine Policy 2000; 24 (3):179– 91]. Marine Policy, 24(6):501–502, November 2000. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X00000300.

Allison:2001:LAM

[AE01] Edward H. Allison and Frank Ellis. The livelihoods approach and management of small-scale fisheries. Marine Policy, 25(5):377– 388, September 2001. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X01000239.

Asche:2008:FBI

[AEG^{+}08] Frank Asche, Håkan Eggert, Eyjolfur Gudmundsson, Ayoe Hoff, and Sean Pascoe. Fisher’s behaviour with individual vessel quotas — over-capacity and potential rent: Five case studies. Marine Policy, 32(6):920–927, November 2008.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000225.

Agnew:2000:IUF

[Agn00] D. J. Agnew. The illegal and unregulated fishery for toothfish in the Southern Ocean, and the CCAMLR catch documentation scheme. Marine Policy, 24(5):361–374, September 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X00000129.

Ardron:2008:MSP

[AGPT08] Jeff Ardron, Kristina Gjerde, Sian Pullen, and Virginie Tilot. Marine spatial planning in the high seas. Marine Policy, 32(5):832–839, September 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X08000729.

Allison:2006:PPS

[AH06] Edward H. Allison and Benoit Horemans. Putting the principles of the Sustainable Livelihoods Approach into fisheries development policy and practice. Marine Policy, 30(6):757–766, November 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000042.

Arnason:2000:CFM

[AHS00] Ragnar Arnason, Rögnvaldur Hannesson, and William E. Schrank. Costs of fisheries management: the cases of Iceland, Norway and Newfoundland. Marine Policy, 24(3):233–243, May 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X99000299.

Aarset:2009:PRR

[AJ09]  Bernt Aarset and Stig-Erik Jakobsen.  Political regula-

tion and radical institutional change: the case of aquaculture in Norway. Marine Policy, 33(2):280–287, March 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0800122X.

Alder:2002:FFB

[AL02] Jacqueline Alder and Gail Lugten. Frozen fish block: how commited are North Atlantic States to accountability, conservation and management of fisheries? Marine Policy, 26(5):345– 357, September 2002. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X02000167.

Alpizar:2006:PFM

[Alp06] Marco Antonio Quesada Alpízar. Participation and fisheries management in Costa Rica: From theory to practice. Marine Policy, 30(6):641–650, November 2006. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000710.

Adrianto:2005:ALS

[AMS05] Luky Adrianto, Yoshiaki Matsuda, and Yoshiaki Sakuma. Assessing local sustainability of fisheries system: a multicriteria participatory approach with the case of Yoron Island, Kagoshima prefecture, Japan. Marine Policy, 29(1):9–23, January 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000144.

Abbott:2009:DIP

[AMW09] Joshua Abbott, Vishwanie Maharaj, and James E. Wilen. Designing ITQ programs for commercial recreational fishing. Marine Policy, 33(5):766–774, September 2009. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X0900027X.

Andersen:2009:EGL

[ANL09] Jesper L. Andersen, Max Nielsen, and Erik Lindebo. Economic gains of liberalising access to fishing quotas within the European Union. Marine Policy, 33(3):497–503, May 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001577.

Anonymous:2000:I

[Ano00a] Anonymous. Index. Marine Policy, 24(6):I–II, November 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X00000312.

Anonymous:2000:PJa

[Ano00b] Anonymous. Pages 1–82 (January 2000). Marine Policy, 24(1): ??, January 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2000:PMb

[Ano00c] Anonymous. Pages 179-274 (May 2000). Marine Policy, 24(3): ??, May 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2000:PJb

[Ano00d] Anonymous. Pages 275-360 (July 2000). Marine Policy, 24(4): ??, July 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2000:PS

[Ano00e] Anonymous. Pages 361-440 (September 2000). Marine Policy, 24 (5):??, September 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2000:PN

[Ano00f] Anonymous. Pages 441-502 (November 2000). Marine Policy, 24 (6):??, November 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2000:PMa

[Ano00g] Anonymous. Pages 83-178 (March 2000). Marine Policy, 24(2): ??, March 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2001:IV

[Ano01a] Anonymous. Index to Volume 25 (2001). Marine Policy, 25(6): I–II, November 2001. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X01000380.

Anonymous:2001:PJa

[Ano01b] Anonymous. Pages 1–90 (January 2001). Marine Policy, 25(1): ??, January 2001. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2001:PMb

[Ano01c] Anonymous. Pages 173-250 (May 2001). Marine Policy, 25(3): ??, May 2001. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2001:PJb

[Ano01d] Anonymous. Pages 251-322 (July 2001). Marine Policy, 25(4): ??, July 2001. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2001:PS

[Ano01e] Anonymous. Pages 323-388 (September 2001). Marine Policy, 25 (5):??, September 2001. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2001:PMa

[Ano01f] Anonymous. Pages 91-172 (March 2001). Marine Policy, 25(2): ??, March 2001. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2002:EBA

[Ano02a] Anonymous. Editorial Board and aims and scope. Marine Policy, 26(6):ifc, November 2002. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X02000611.

Anonymous:2002:IV

[Ano02b] Anonymous. Index to Volume 26 (2202). Marine Policy, 26(6): I–II, November 2002. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X02000672.

Anonymous:2002:PJa

[Ano02c] Anonymous. Pages 1–74 (January 2002). Marine Policy, 26(1): ??, January 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2002:PMb

[Ano02d] Anonymous. Pages 143-230 (May 2002). Marine Policy, 26(3): ??, May 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2002:PJb

[Ano02e] Anonymous. Pages 231-314 (July 2002). Marine Policy, 26(4): ??, July 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2002:PS

[Ano02f] Anonymous. Pages 315-402 (September 2002). Marine Policy, 26 (5):??, September 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2002:PN

[Ano02g] Anonymous. Pages 403-492 (November 2002). Marine Policy, 26 (6):??, November 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2002:PMa

[Ano02h] Anonymous. Pages 75-142 (March 2002). Marine Policy, 26(2): ??, March 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2003:EBa

[Ano03a] Anonymous. Editorial Board. Marine Policy, 27(1):ifc, January 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0200074X.

Anonymous:2003:EBb

[Ano03b] Anonymous. Editorial Board. Marine Policy, 27(2):ifc, March 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000095.

Anonymous:2003:EBc

[Ano03c] Anonymous. Editorial Board. Marine Policy, 27(3):ifc, May 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000289.

Anonymous:2003:EBAa

[Ano03d] Anonymous. Editorial Board and aims and scope. Marine Policy, 27(4):ifc, July 2003. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X03000526.

Anonymous:2003:EBAb

[Ano03e] Anonymous. Editorial Board and aims and scope. Marine Policy, 27(5):ifc, September 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X0300068X.

Anonymous:2003:EBAc

[Ano03f] Anonymous. Editorial Board and aims and scope. Marine Policy, 27(6):ifc, November 2003. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X03000927.

Anonymous:2003:IV

[Ano03g] Anonymous. Index to Volume 27 (2003). Marine Policy, 27(6): I–III, November 2003. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X03000988.

Anonymous:2003:PJ

[Ano03h] Anonymous. Pages 1–96 (January 2003). Marine Policy, 27(1): ??, January 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2003:PMb

[Ano03i] Anonymous. Pages 193-288 (May 2003). Marine Policy, 27(3): ??, May 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2003:PS

[Ano03j] Anonymous. Pages 357-448 (September 2003). Marine Policy, 27 (5):??, September 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2003:PN

[Ano03k] Anonymous. Pages 449-556 (November 2003). Marine Policy, 27 (6):??, November 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2003:PMa

[Ano03l] Anonymous. Pages 97-192 (March 2003). Marine Policy, 27(2): ??, March 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2004:EBAb

[Ano04a] Anonymous. Ed board and aims and scope. Marine Policy, 28 (2):ifc, March 2004. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X0300160X.

Anonymous:2004:EBAa

[Ano04b] Anonymous. Editorial Board and aims and scope. Marine Policy, 28(1):ifc, January 2004. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X03001398.

Anonymous:2004:EBAc

[Ano04c] Anonymous. Editorial Board and aims and scope. Marine Policy, 28(3):ifc, May 2004. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X04000041.

Anonymous:2004:EBAd

[Ano04d] Anonymous. Editorial Board and aims and scope. Marine Policy, 28(4):ifc, July 2004. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X04000223.

Anonymous:2004:EBAe

[Ano04e] Anonymous. Editorial Board and aims and scope. Marine Policy, 28(6):ifc, November 2004. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X04000673.

Anonymous:2004:IEB

[Ano04f] Anonymous. IFC: Editorial Board. Marine Policy, 28(5):ifc, September 2004. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X04000417.

Anonymous:2004:IV

[Ano04g] Anonymous. Index to Volume 28. Marine Policy, 28(6):I– III, November 2004. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X04000727.

Anonymous:2004:PJa

[Ano04h] Anonymous. Pages 1–96 (January 2004). Marine Policy, 28(1): ??, January 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2004:PMb

[Ano04i] Anonymous. Pages 189-282 (May 2004). Marine Policy, 28(3): ??, May 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2004:PJb

[Ano04j] Anonymous. Pages 283-360 (July 2004). Marine Policy, 28(4): ??, July 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2004:PS

[Ano04k] Anonymous. Pages 361-450 (September 2004). Marine Policy, 28 (5):??, September 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2004:PN

[Ano04l] Anonymous. Pages 451-556 (November 2004). Marine Policy, 28 (6):??, November 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2004:PMa

[Ano04m] Anonymous. Pages 97-188 (March 2004). Marine Policy, 28(2): ??, March 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2005:EBAa

[Ano05a] Anonymous. Editorial Board and aims and scope. Marine Policy, 29(1):ifc, January 2005. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X04000910.

Anonymous:2005:EBAb

[Ano05b] Anonymous. Editorial Board and aims and scope. Marine Policy, 29(2):ifc, March 2005. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X04001095.

Anonymous:2005:EBAc

[Ano05c] Anonymous. Editorial Board and aims and scope. Marine Policy, 29(3):ifc, May 2005. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X04001174.

Anonymous:2005:EBAd

[Ano05d] Anonymous. Editorial Board and aims and scope. Marine Policy, 29(4):ifc, July 2005. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X0400123X.

Anonymous:2005:EBAe

[Ano05e] Anonymous. Editorial Board and aims and scope. Marine Policy, 29(5):CO2, September 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X05000163.

Anonymous:2005:EBAf

[Ano05f] Anonymous. Editorial Board and aims and scope. Marine Policy, 29(6):CO2, November 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X0500062X.

Anonymous:2005:PJa

[Ano05g] Anonymous. Pages 1–94 (January 2005). Marine Policy, 29(1): ??, January 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2005:PM

[Ano05h] Anonymous. Pages 189-290 (May 2005). Marine Policy, 29(3): ??, May 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2005:PJb

[Ano05i] Anonymous. Pages 291-384 (July 2005). Marine Policy, 29(4): ??, July 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2005:PS

[Ano05j] Anonymous. Pages 385-480 (September 2005). Marine Policy, 29 (5):??, September 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2005:PN

[Ano05k] Anonymous. Pages 481-538 (November 2005). Marine Policy, 29 (6):??, November 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2006:EBAa

[Ano06a] Anonymous. Editorial Board/aims and scope. Marine Policy, 30 (1):CO2, January 2006. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X05000850.

Anonymous:2006:EBAb

[Ano06b] Anonymous. Editorial Board/aims and scope. Marine Policy, 30 (2):CO2, March 2006. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X05001004.

Anonymous:2006:EBAc

[Ano06c] Anonymous. Editorial Board/aims and scope. Marine Policy, 30(3):CO2, May 2006. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X06000108.

Anonymous:2006:EBAd

[Ano06d] Anonymous. Editorial Board/aims and scope. Marine Policy, 30(4):CO2, July 2006. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X06000285.

Anonymous:2006:EBAe

[Ano06e] Anonymous. Editorial Board/aims and scope. Marine Policy, 30(5):CO2, September 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X06000546.

Anonymous:2006:EBAf

[Ano06f] Anonymous. Editorial Board/aims and scope. Marine Policy, 30 (6):CO2, November 2006. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X06000637.

Anonymous:2006:FRS

[Ano06g] Anonymous. Fisheries Reform in South Africa 1994–2004. Marine Policy, 30(1):??, January 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2006:IVa

[Ano06h] Anonymous. Index to Volume 29. Marine Policy, 30(1):I–III, January 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000837.

Anonymous:2006:IVb

[Ano06i] Anonymous. Index to volume 30. Marine Policy, 30(6):I– IV, November 2006. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X06000728.

Anonymous:2006:PMa

[Ano06j] Anonymous. Pages 111-198 (March 2006). Marine Policy, 30(2): ??, March 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2006:PMb

[Ano06k] Anonymous. Pages 199-298 (May 2006). Marine Policy, 30(3): ??, May 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2006:PJ

[Ano06l] Anonymous. Pages 299-454 (July 2006). Marine Policy, 30(4): ??, July 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2006:PS

[Ano06m] Anonymous. Pages 455-604 (September 2006). Marine Policy, 30 (5):??, September 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2006:PN

[Ano06n] Anonymous. Pages 605-842 (November 2006). Marine Policy, 30 (6):??, November 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2007:EBAa

[Ano07a] Anonymous. Editorial Board/aims and scope. Marine Policy, 31 (1):CO2, January 2007. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X0600087X.

Anonymous:2007:EBAb

[Ano07b] Anonymous. Editorial Board/aims and scope. Marine Policy, 31 (2):CO2, March 2007. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X06001102.

Anonymous:2007:EBAc

[Ano07c] Anonymous. Editorial Board/aims and scope. Marine Policy, 31(3):CO2, May 2007. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X0600128X.

Anonymous:2007:EBAd

[Ano07d] Anonymous. Editorial Board/aims and scope. Marine Policy, 31(4):ifc, July 2007. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X07000310.

Anonymous:2007:EBAe

[Ano07e] Anonymous. Editorial Board/aims and scope. Marine Policy, 31 (5):ifc, September 2007. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X07000632.

Anonymous:2007:EBAf

[Ano07f] Anonymous. Editorial Board/aims and scope. Marine Policy, 31 (6):ifc, November 2007. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X07000826.

Anonymous:2007:OSM

[Ano07g] Anonymous. Online submission to Marine Policy. Marine Policy, 31(4):359, July 2007. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X07000164.

Anonymous:2007:PMb

[Ano07h] Anonymous. Pages 229-358 (May 2007). Marine Policy, 31(3): ??, May 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2007:PJ

[Ano07i] Anonymous. Pages 359-584 (July 2007). Marine Policy, 31(4): ??, July 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2007:PN

[Ano07j] Anonymous. Pages 645-746 (November 2007). Marine Policy, 31 (6):??, November 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2007:PMa

[Ano07k] Anonymous. Pages 81-228 (March 2007). Marine Policy, 31(2): ??, March 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2008:EBAa

[Ano08a] Anonymous. Editorial Board/aims and scope. Marine Policy, 32 (1):ifc, January 2008. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X0700139X.

Anonymous:2008:EBAb

[Ano08b] Anonymous. Editorial Board/aims and scope. Marine Policy, 32(2):ifc, March 2008. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08000109.

Anonymous:2008:EBAc

[Ano08c] Anonymous. Editorial Board/aims and scope. Marine Policy, 32(3):ifc, May 2008. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08000420.

Anonymous:2008:EBAd

[Ano08d] Anonymous. Editorial Board/aims and scope. Marine Policy, 32(4):ifc, July 2008. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08000821.

Anonymous:2008:EBAe

[Ano08e] Anonymous. Editorial Board/aims and scope. Marine Policy, 32 (5):ifc, September 2008. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08001097.

Anonymous:2008:EBAf

[Ano08f] Anonymous. Editorial Board/aims and scope. Marine Policy, 32 (6):ifc, November 2008. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08001334.

Anonymous:2008:PJa

[Ano08g] Anonymous. Pages 1–146 (January 2008). Marine Policy, 32 (1):??, January 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2008:PM

[Ano08h] Anonymous. Pages 255-534 (May 2008). Marine Policy, 32(3): ??, May 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2008:PJb

[Ano08i] Anonymous. Pages 535-758 (July 2008). Marine Policy, 32(4): ??, July 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2008:PN

[Ano08j] Anonymous. Pages 845-1102 (November 2008). Marine Policy, 32(6):??, November 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2009:EBAa

[Ano09a] Anonymous. Editorial Board/aims and scope. Marine Policy, 33 (1):ifc, January 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08001619.

Anonymous:2009:EBAb

[Ano09b] Anonymous. Editorial Board/aims and scope. Marine Policy, 33(2):ifc, March 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X0800170X.

Anonymous:2009:EBAc

[Ano09c] Anonymous. Editorial Board/aims and scope. Marine Policy, 33(3):ifc, May 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X09000104.

Anonymous:2009:EBAd

[Ano09d] Anonymous. Editorial Board/aims and scope. Marine Policy, 33(4):ifc, July 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X09000359.

Anonymous:2009:EBAe

[Ano09e] Anonymous. Editorial Board/aims and scope. Marine Policy, 33 (5):ifc, September 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X09000761.

Anonymous:2009:EBAf

[Ano09f] Anonymous. Editorial Board/aims and scope. Marine Policy, 33 (6):ifc, November 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X09001146.

Anonymous:2009:PJa

[Ano09g] Anonymous. Pages 1–182 (January 2009). Marine Policy, 33 (1):??, January 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2009:PMa

[Ano09h] Anonymous. Pages 183-434 (March 2009). Marine Policy, 33(2): ??, March 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2009:PMb

[Ano09i] Anonymous. Pages 435-536 (May 2009). Marine Policy, 33(3): ??, May 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2009:PJb

[Ano09j] Anonymous. Pages 537-736 (July 2009). Marine Policy, 33(4): ??, July 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2009:PS

[Ano09k] Anonymous. Pages 737-868 (September 2009). Marine Policy, 33 (5):??, September 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Al-Oufi:2000:OUA

[AOPM00] Hamed Al-Oufi, Andrew Palfreman, and Ewen McLean. Observations upon the Al-Batinah artisinal fishery, the Sultanate of Oman. Marine Policy, 24(5):423–429, September 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X00000178.

Armstrong:2001:TPW

[Arm01] Claire W. Armstrong. Theory and practice of why auctions differ — a study of two fish auctions in Norway. Marine Policy, 25(3):

209–214, May 2001. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X01000112.

Abdallah:2007:HAB

[AS07] Patrízia R. Abdallah and Ussif R. Sumaila. An historical account of Brazilian public policy on fisheries subsidies. Marine Policy, 31(4):444–450, July 2007. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X07000073.

Armstrong:2008:FPP

[AvdH08] Claire W. Armstrong and Sybille van den Hove. The formation of policy for protection of cold-water coral off the coast of Norway. Marine Policy, 32(1):66–73, January 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000516.

Anferova:2005:FQA

[AVH05] Elena Anferova, Markus Vetemaa, and Rögnvaldur Hannesson. Fish quota auctions in the Russian Far East: a failed experiment. Marine Policy, 29(1):47–56, January 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0400017X.

Alderton:2002:GRM

[AW02] Tony Alderton and Nik Winchester. Globalisation and deregulation in the maritime industry. Marine Policy, 26(1):35–43, January 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X01000343.

Baskaran:2005:ASS

[BA05] R. Baskaran and J. L. Anderson. Atlantic sea scallop management: an alternative rights-based cooperative approach to resource sustainability. Marine Policy, 29(4):357–369, July 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000569.

Ball:2004:ICO

[Bal04] Desmond Ball. Intelligence collection operations and EEZs: the implications of new technology. Marine Policy, 28(1):67–82, January 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03001271.

Bateman:2005:HSE

[Bat05] Sam Bateman. Hydrographic surveying in the EEZ: differences and overlaps with marine scientific research. Marine Policy, 29 (2):163–174, March 2005. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X04000855.

Bhat:2006:MTI

[BB06] Mahadev G. Bhat and Ramachandra Bhatta. Mechanization and technical interactions in multi-species Indian fisheries: implications for economic and biological sustainability. Marine Policy, 30(3):237–248, May 2006. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X05000096.

Boude:2001:RAF

[BBB01] Jean-Pierre Boude, Jean Boncoeur, and Denis Bailly. Regulating the access to fisheries: learning from European experiences. Marine Policy, 25(4):313–322, July 2001. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X01000033.

Binkley:2008:LOC

[BBNB08] M. Binkley, K. Bigney, B. Neis, and S. Bornstein. Lessons from offshore: Challenges and opportunities in linking data to promote understanding of accidents and injuries among Newfoundland and Labrador fishers 1989–2001. Marine Policy, 32(6):905– 912, November 2008. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X0800016X.

Branch:2006:FST

[BC06] George M. Branch and Barry M. Clark. Fish stocks and their management: the changing face of fisheries in South Africa. Marine Policy, 30(1):3–17, January 2006. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000436.

Barclay:2007:GTI

[BC07a] Kate Barclay and Ian Cartwright. Governance of tuna industries: the key to economic viability and sustainability in the Western and Central Pacific Ocean. Marine Policy, 31(3):348–358, May 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06001023.

Blake:2007:CFF

[BC07b] Andon Blake and Gary A. Campbell. Conflict over flying fish: the dispute between Trinidad & Tobago and Barbados. Marine Policy, 31(3):327–335, May 2007. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X0600100X.

Bogaert:2009:DMP

[BCM09a] Dirk Bogaert, An Cliquet, and Frank Maes. Designation of marine protected areas in Belgium: a legal and ecological success? Marine Policy, 33(6):878–886, November 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000591.

Bose:2009:SVF

[BCM09b] Shekar Bose and Adnan Crees-Morris. Stakeholder’s views on fisheries compliance: an Australian case study. Marine Policy, 33 (2):248–253, March 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X0800105X.

Bavinck:2008:TZS

[BdKvD^{+}08] M. Bavinck, L. de Klerk, D. van Dijk, J. V. Rothuizen, A. N. Blok, J. R. Bokhorst, E. K. van Haastrecht, T. J. C. van de Loo, J. G. J. Quaedvlieg, and J. Scholtens. Timezoning for the safe-guarding of capture fisheries: a closed season in Tamil Nadu, India. Marine Policy, 32(3):369–378, May 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000930.

Beem:2007:CMT

[Bee07] Betsi Beem. Co-management from the top? The roles of policy entrepreneurs and distributive conflict in developing comanagement arrangements. Marine Policy, 31(4):540–549, July 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06001333.

Bennett:2001:MRP

[Ben01] Paul Bennett. Mutual risk: P&I insurance clubs and maritime safety and environmental performance. Marine Policy, 25(1): 13–21, January 2001. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X00000294.

Bennett:2005:GFD

[Ben05] Elizabeth Bennett. Gender, fisheries and development. Marine Policy, 29(5):451–459, September 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X04000648.

Bergin:2002:EAN

[Ber02] Anthony Bergin. East Asian naval developments — sailing into rough seas. Marine Policy, 26(2):121–131, March 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X01000434.

Berman:2008:EST

[Ber08] Matthew D. Berman. Endangered species, threatened fisheries: Science to the rescue! Evaluating the Congressionally designated Steller Sea Lion Research Program. Marine Policy, 32(4):580– 591, July 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001285.

Bess:2001:NZI

[Bes01] Randall Bess. New Zealand’s indigenous people and their claims to fisheries resources. Marine Policy, 25(1):23–32, January 2001. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X00000324.

Bess:2005:ENZ

[Bes05] Randall Bess. Expanding New Zealand’s quota management system. Marine Policy, 29(4):339–347, July 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000533.

Bess:2006:NZS

[Bes06] Randall Bess. New Zealand seafood firm competitiveness in export markets: the role of the quota management system and aquaculture legislation. Marine Policy, 30(4):367–378, July 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000400.

Boyes:2007:PMU

[BET^{+}07] Suzanne J. Boyes, Michael Elliott, Shona M. Thomson, Stephen Atkins, and Paul Gilliland. A proposed multiple-use zoning scheme for the Irish Sea.: an interpretation of current legislation through the use of GIS-based zoning approaches and effectiveness for the protection of nature conservation interests. Marine Policy, 31(3):287–298, May 2007. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X06000856.

Bess:2000:RPR

[BH00] Randall Bess and Michael Harte. The role of property rights in the development of New Zealand’s seafood industry. Marine Policy, 24(4):331–339, July 2000. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X00000099.

Bowett:2009:WCE

[BH09] Julia Bowett and Pete Hay. Whaling and its controversies: Examining the attitudes of Japan’s youth. Marine Policy, 33(5):775– 783, September 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X09000281.

Ban:2009:SMC

[BHJV09] Natalie C. Ban, Gretchen J. A. Hansen, Michael Jones, and Amanda C. J. Vincent. Systematic marine conservation planning in data-poor regions: Socioeconomic data is essential. Marine Policy, 33(5):794–800, September 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X0900030X.

Bremner:2009:UBN

[BJBC09] Graeme Bremner, Peter Johnstone, Tracy Bateson, and Philip Clarke. Unreported bycatch in the New Zealand West Coast South Island hoki fishery. Marine Policy, 33(3):504–512, May 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001590.

Bjorndal:2009:ORP

[Bjø09] Trond Bjørndal. Overview, roles, and performance of the North East Atlantic Fisheries Commission (NEAFC). Marine Policy, 33(4):685–697, July 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X09000086.

Bijwaard:2009:ASL

[BK09] Govert E. Bijwaard and Sabine Knapp. Analysis of ship life cycles — The impact of economic cycles and ship inspections. Marine Policy, 33(2):350–369, March 2009. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001310.

Bell:2009:PUF

[BKV^{+}09] Johann D. Bell, Mecki Kronen, Aliti Vunisea, Warwick J. Nash, Gregory Keeble, Andreas Demmke, Scott Pontifex, and Serge Andréfouët. Planning the use of fish for food security in the Pacific. Marine Policy, 33(1):64–76, January 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000778.

Burger:2000:CRC

[BL00] Joanna Burger and Justin Leonard. Conflict resolution in coastal waters: the case of personal watercraft. Marine Policy, 24(1): 61–67, January 2000. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X99000135.

Bruckmeier:2008:SCF

[BL08] Karl Bruckmeier and Christina Höj Larsen. Swedish coastal fisheries — From conflict mitigation to participatory management. Marine Policy, 32(2):201–211, March 2008. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001133.

Brandt:2006:DEP

[BM06] Sylvia Brandt and David McEvoy. Distributional effects of property rights: Transitions in the Atlantic herring fishery. Marine Policy, 30(6):659–670, November 2006. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000758.

Brown:2007:GFE

[BM07] James Brown and Graeme Macfadyen. Ghost fishing in European waters: Impacts and management responses. Marine Policy, 31 (4):488–504, July 2007. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X07000024.

Barnes:2008:MEA

[BM08] Cassandra Barnes and Katherine W. McFadden. Marine ecosystem approaches to management: challenges and lessons in the United States. Marine Policy, 32(3):387–392, May 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000954.

Breffle:2009:RFS

[BM09] William S. Breffle and Kristen K. Maroney. The restoration of fishing services and the conveyance of risk information in the Southern California Bight. Marine Policy, 33(4):561–570, July 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001929.

Bennett:2001:TBU

[BNA^{+}01] Elizabeth Bennett, Arthur Neiland, Emilia Anang, Paul Bannerman, A. Atiq Rahman, Saleemul Huq, Shajahan Bhuiya, Mark Day, Michelle Fulford-Gardiner, and Wesley Clerveaux. Towards a better understanding of conflict management in tropical fisheries: evidence from Ghana, Bangladesh and the Caribbean. Marine Policy, 25(5):365–376, September 2001. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X01000227.

Bodiguel:2002:FFC

[Bod02] Clotilde Bodiguel. Fishermen facing the commercial lobster fishery licensing policy in the Canadian Maritime Provinces: origins of illegal strategies, 1960–2000. Marine Policy, 26(4):271–281, July 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0200009X.

Borgese:2001:CMA

[Bor01] E. M. Borgese. Caird Medal address. Marine Policy, 25(6):391– 397, November 2001. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X01000252.

Burke:2008:MJJ

[BP08] Lenore Burke and John Phyne. Made in Japan: the Japanese market and herring roe production and management in Canada’s southern Gulf of St. Lawrence. Marine Policy, 32(1):79–88, January 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0700053X.

Bartlett:2009:MRP

[BPM09] C. Y. Bartlett, K. Pakoa, and C. Manua. Marine reserve phenomenon in the Pacific islands. Marine Policy, 33(4):673–678, July 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000062.

Bess:2007:SCN

[BR07] Randall Bess and Ramana Rallapudi. Spatial conflicts in New Zealand fisheries: the rights of fishers and protection of the marine environment. Marine Policy, 31(6):719–729, November 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000152.

Brennan:2008:SMW

[BR08] Ruth E. Brennan and Lynda D. Rodwell. Sustainable management of wild Irish Atlantic salmon: Keys found through the looking-glass. Marine Policy, 32(6):1072–1079, November 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000389.

Branch:2008:AFW

[Bra08] Trevor A. Branch. Not all fisheries will be collapsed in 2048. Marine Policy, 32(1):38–39, January 2008. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000474.

Branch:2006:RTL

[BRH06] Trevor A. Branch, Kate Rutherford, and Ray Hilborn. Replacing trip limits with individual transferable quotas: implications for discarding. Marine Policy, 30(3):281–292, May 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000059.

Brown:2005:ADS

[Bro05] James Brown. An account of the dolphin-safe tuna issue in the UK. Marine Policy, 29(1):39–46, January 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000168.

Bailey:2008:MAF

[BRS08] Megan Bailey, Christovel Rotinsulu, and U. Rashid Sumaila. The migrant anchovy fishery in Kabui Bay, Raja Ampat, Indonesia: Catch, profitability, and income distribution. Marine Policy, 32 (3):483–488, May 2008. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X07001066.

Bradshaw:2004:UPA

[BT04] Matt Bradshaw and Oliver Tully. Use of private access rights in fisheries: effective management through public transferability? Marine Policy, 28(5):411–418, September 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03001325.

Butt:2007:ICS

[But07] Nickie Butt. The impact of cruise ship generated waste on home ports and ports of call: a study of Southampton. Marine Policy, 31(5):591–598, September 2007. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000218.

Brady:2009:FPF

[BW09] Mark Brady and Staffan Waldo. Fixing problems in fisheries — integrating ITQs, CBM and MPAs in management. Marine Policy, 33(2):258–263, March 2009. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001073.

Bax:2003:MIA

[BWA^{+}03] Nicholas Bax, Angela Williamson, Max Aguero, Exequiel Gonzalez, and Warren Geeves. Marine invasive alien species: a threat to global biodiversity. Marine Policy, 27(4):313–323, July 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000411.

Berghofer:2008:SPE

[BWR08] Augustin Berghöfer, Heidi Wittmer, and Felix Rauschmayer. Stakeholder participation in ecosystem-based approaches to fisheries management: a synthesis from European research projects. Marine Policy, 32(2):243–253, March 2008. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001169.

Campbell:2000:MTF

[Cam00] H. F. Campbell. Managing tuna fisheries: a new strategy for the Western and Central Pacific Ocean. Marine Policy, 24(2): 159–163, March 2000. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X99000275.

Carneiro:2007:PEO

[Car07] Gonçalo Carneiro. The parallel evolution of ocean and coastal management policies in Portugal. Marine Policy, 31(4):421–433, July 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000139.

Campbell:2000:ITC

[CBB00] David Campbell, Debbie Brown, and Tony Battaglene. Individual transferable catch quotas: Australian experience in the southern bluefin tuna fishery. Marine Policy, 24(2):109–117, March 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X99000172.

Cooper:2009:CRP

[CC09]  J. A. G. Cooper and V. Cummins.  Coastal research

and policy integration in northwest Europe. The CORE-POINT project. Marine Policy, 33(6):869–870, November 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0900058X.

Compas:2007:MWM

[CCCD07] Eric Compas, Beverley Clarke, Cecile Cutler, and Kathy Daish. Murky waters: Media reporting of marine protected areas in South Australia. Marine Policy, 31(6):691–697, November 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0700019X.

Clapham:2007:WIC

[CCG^{+}07] Phillip J. Clapham, Simon Childerhouse, Nicolas J. Gales, Lorenzo Rojas-Bracho, Michael F. Tillman, and Robert L. Brownell. The whaling issue: Conservation, confusion, and casuistry. Marine Policy, 31(3):314–319, May 2007. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000984.

Catchpole:2005:DNS

[CFG05] T. L. Catchpole, C. L. J. Frid, and T. S. Gray. Discards in North Sea fisheries: causes, consequences and solutions. Marine Policy, 29(5):421–430, September 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000612.

Catchpole:2006:RDP

[CFG06] T. L. Catchpole, C. L. J. Frid, and T. S. Gray. Resolving the discard problem — a case study of the English, Nephrops fishery. Marine Policy, 30(6):821–831, November 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000212.

Christie:2007:AFE

[CFW^{+}07] Patrick Christie, David L. Fluharty, Alan T. White, Liza Eisma-Osorio, and William Jatulan. Assessing the feasibility of ecosystem-based fisheries management in tropical contexts. Marine Policy, 31(3):239–250, May 2007. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000807.

Compton:2008:CEW

[CGHA08] Ross Compton, Lissa Goodwin, Richard Handy, and Victor Abbott. A critical examination of worldwide guidelines for minimising the disturbance to marine mammals during seismic surveys. Marine Policy, 32(3):255–262, May 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000607.

Chavez:2008:IUI

[CGS08] Carlos Chávez, Nuria González, and Hugo Salgado. ITQs under illegal fishing: an application to the red shrimp fishery in Chile. Marine Policy, 32(4):570–579, July 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001261.

Cheong:2003:PTF

[Che03] So-Min Cheong. Privatizing tendencies: fishing communities and tourism in Korea. Marine Policy, 27(1):23–29, January 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000477.

Cheong:2008:NDC

[Che08] So-Min Cheong. A new direction in coastal management. Marine Policy, 32(6):1090–1093, November 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X08000407.

Chiu:2007:LST

[Chi07] Rong-Her Chiu. The liberalization of shipping in Taiwan. Marine Policy, 31(3):258–265, May 2007. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000820.

Cho:2006:EOG

[Cho06] Dong Oh Cho. Evaluation of the ocean governance system in Korea. Marine Policy, 30(5):570–579, September 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000692.

Cho:2007:EMV

[Cho07] Dong-Oh Cho. The effects of the M/V Sea Prince accident on maritime safety management in Korea. Marine Policy, 31(6): 730–735, November 2007. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X07000127.

Chuenpagdee:2007:SZF

[CJ07] Ratana Chuenpagdee and Svein Jentoft. Step zero for fisheries co-management: What precedes implementation. Marine Policy, 31(6):657–668, November 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X07000437.

Clark:2006:CCL

[Cla06] Barry M. Clark. Climate change: a looming challenge for fisheries management in southern Africa. Marine Policy, 30(1):84–95, January 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000448.

Claesson:2009:EBF

[Cla09] Stefan Claesson. An ecosystem-based framework for governance and management of maritime cultural heritage in the USA. Marine Policy, 33(4):698–706, July 2009. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000189.

Clifton:2003:PCM

[Cli03] Julian Clifton. Prospects for co-management in Indonesia’s marine protected areas. Marine Policy, 27(5):389–395, September 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000265.

Chang:2008:PSF

[CLT08] Young-Tae Chang, Sang-Yoon Lee, and Jose L. Tongzon. Port selection factors by shipping lines: Different perspectives between trunk liners and feeder service providers. Marine Policy, 32(6): 877–885, November 2008. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08000079.

Chang:2006:DDS

[CLyB06] Chih-Ching Chang, Kuang Lin, and Jya yuan Bao. Developing direct shipping across the Taiwan Straits. Marine Policy, 30(4): 392–399, July 2006. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X05000266.

Criddle:2000:RIU

[CM00] Keith R. Criddle and Seth Macinko. A requiem for the IFQ in US fisheries? Marine Policy, 24(6):461–469, November 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X00000257.

Castillo-Manzano:2009:LCP

[CMCNL^{+}09] José I. Castillo-Manzano, Mercedes Castro-Nuño, Fernando González Laxe, Lourdes López-Valpuesta, and M. Teresa Arévalo-Quijada. Low-cost port competitiveness index: Implementation in the Spanish port system. Marine Policy, 33(4):591–598, July 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001942.

Cariou:2009:ETF

[CMW09] Pierre Cariou, Maximo Q. Mejia, and Francois-Charles Wolff. Evidence on target factors used for port state control inspections. Marine Policy, 33(5):847–859, September 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000487.

Crowder:2008:EEI

[CN08] Larry Crowder and Elliott Norse. Essential ecological insights for marine ecosystem-based management and marine spatial planning. Marine Policy, 32(5):772–778, September 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000651.

Conway:2007:MSS

[Con07] Gemma Conway. Monitoring the state of the Solent. Marine Policy, 31(5):632–637, September 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X07000279.

Cook:2005:LBD

[Coo05] Beverly Cook. Lobster boat diplomacy: the Canada–US grey zone. Marine Policy, 29(5):385–390, September 2005.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000582.

Corkett:2005:PRU

[Cor05] Christopher J. Corkett. The Pew Report on US fishery councils: a critique from the Open Society. Marine Policy, 29(3):247–253, May 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000351.

Carollo:2009:IDD

[CROP09] Cristina Carollo, Dave J. Reed, John C. Ogden, and David Palandro. The importance of data discovery and management in advancing ecosystem-based management. Marine Policy, 33(4): 651–653, July 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X09000025.

Chaniotis:2007:IPA

[CS07] Peter Chaniotis and Selina Stead. Interviewing people about the coast on the coast: Appraising the wider adoption of ICZM in North East England. Marine Policy, 31(4):517–526, July 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06001369.

Campbell:2009:CMS

[CSG^{+}09] Lisa M. Campbell, Jennifer J. Silver, Noella J. Gray, Sue Ranger, Annette Broderick, Tatum Fisher, Matthew H. Godfrey, Shannon Gore, John Jeffers, Corrine Martin, Andrew Mc-Gowan, Peter Richardson, Carlos Sasso, Lorna Slade, and Brendan Godley. Co-management of sea turtle fisheries: Biogeography versus geopolitics. Marine Policy, 33(1):137–145, January 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000936.

Coffen-Smout:2000:SCC

[CSH00] Scott Coffen-Smout and Glen J. Herbert. Submarine cables: a challenge for ocean management. Marine Policy, 24(6):441– 448, November 2000. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X00000270.

CoswigKalikoski:2002:FIE

[CVL02] Daniela Coswig Kalikoski, Marcelo Vasconcellos, and Les Lavkulich. Fitting institutions to ecosystems: the case of artisanal fisheries management in the estuary of Patos Lagoon. Marine Policy, 26(3):179–196, May 2002. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X01000483.

Crosoer:2006:ISA

[CvSP06] David Crosoer, Lance van Sittert, and Stefano Ponte. The integration of South African fisheries into the global economy: Past, present and future. Marine Policy, 30(1):18–29, January 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000461.

Crean:2000:TWM

[CW00] Kevin Crean and Steve J. Wisher. Is there the will to manage fisheries at a local level in the European Union? A case study from Shetland. Marine Policy, 24(6):471–481, November 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X00000208.

Cho:2008:ROS

[CW08] Dong-Oh Cho and Mary Anne Whitcomb. A review of the ocean science and technology partnership between US and Korea. Marine Policy, 32(3):502–513, May 2008. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001194.

Curran:2006:PCE

[CWP06] K. J. Curran, P. G. Wells, and A. J. Potter. Proposing a coordinated environmental effects monitoring (EEM) program structure for the offshore petroleum industry, Nova Scotia, Canada. Marine Policy, 30(4):400–411, July 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X0500028X.

Cinner:2009:TIC

[CWRR09] Joshua E. Cinner, Andrew Wamukota, Herilala Randriamahazo, and Ando Rabearisoa. Toward institutions for communitybased management of inshore marine resources in the Western Indian Ocean. Marine Policy, 33(3):489–496, May 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001565.

Davis:2008:ECF

[Dav08] Neil A. Davis. Evaluating collaborative fisheries management planning: a Canadian case study. Marine Policy, 32(6):867– 876, November 2008. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08000067.

Dawson:2006:VIP

[Daw06] Robert Dawson. Vertical integration in the post-IFQ halibut fishery. Marine Policy, 30(4):341–346, July 2006. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000357.

Day:2008:NPM

[Day08] Jon Day. The need and practice of monitoring, evaluating and adapting marine planning and management — lessons from the Great Barrier Reef. Marine Policy, 32(5):823–831, September 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000717.

Doherty:2006:OZN

[DB06] Penny A. Doherty and Mark Butler. Ocean zoning in the Northwest Atlantic. Marine Policy, 30(4):389–391, July 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000254.

delCorral:2008:FAF

[dC08] David Florido del Corral. Focusing on artisanal fleets in a new scenario: the case of Andalusia (Spain). Marine Policy, 32(6):1004–1012, November 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X08000304.

Darbra:2007:SCH

[DCHM07] R. M. Darbra, J. F. E. Crawford, C. W. Haley, and R. J. Morrison. Safety culture and hazard risk perception of Australian and New Zealand maritime pilots. Marine Policy, 31(6):736– 745, November 2007. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X07000115.

Davies:2009:DEG

[DCNP09] R. W. D. Davies, S. J. Cripps, A. Nickson, and G. Porter. Defining and estimating global marine fisheries bycatch. Marine Policy, 33(4):661–672, July 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X09000050.

Douvere:2007:IWM

[DE07] Fanny Douvere and Charles N. Ehler. International Workshop on Marine Spatial Planning, UNESCO, Paris, 8–10 November 2006: a summary. Marine Policy, 31(4):582–583, July 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000103.

Douvere:2008:I

[DE08] Fanny Douvere and Charles Ehler. Introduction. Marine Policy, 32(5):759–761, September 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X08000638.

Detjen:2006:WEP

[Det06] Markus Detjen. The Western European PSSA — Testing a unique international concept to protect imperilled marine ecosystems. Marine Policy, 30(4):442–453, July 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000369.

Devine:2000:PSJ

[Dev00] Dermott Devine. Port state jurisdiction: a judicial contribution from New Zealand. Marine Policy, 24(3):215–219, May 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X99000305.

Daw:2005:FSS

[DG05] Tim Daw and Tim Gray. Fisheries science and sustainability in international policy: a study of failure in the European Union’s Common Fisheries Policy. Marine Policy, 29(3):189–197, May 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000284.

Degnbol:2006:PFH

[DGH^{+}06] Poul Degnbol, Henrik Gislason, Susan Hanna, Svein Jentoft, Jesper Raakjær Nielsen, Sten Sverdrup-Jensen, and Douglas Clyde Wilson. Painting the floor with a hammer: Technical fixes in fisheries management. Marine Policy, 30(5):534–543, September 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000588.

Davis:2001:CPI

[DJ01] Anthony Davis and Svein Jentoft. The challenge and the promise of indigenous peoples’ fishing rights — from dependency to agency. Marine Policy, 25(3):223–237, May 2001. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X01000148.

DeSanto:2007:OMC

[DJ07] Elizabeth M. De Santo and Peter J. S. Jones. Offshore marine conservation policies in the North East Atlantic: Emerging tensions and opportunities. Marine Policy, 31(3):336–347, May 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06001011.

Djalal:2004:ORD

[Dja04] Hasjim Djalal. Options for resolving disagreements. Marine Policy, 28(1):83–87, January 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X03001283.

deJuan:2009:DEI

[dJDT09] S. de Juan, M. Demestre, and S. Thrush. Defining ecological indicators of trawling disturbance when everywhere that can be fished is fished: a Mediterranean case study. Marine Policy, 33 (3):472–478, May 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08001541.

Dwyer:2008:MSF

[DKM08] Peter D. Dwyer, Tanya J. King, and Monica Minnegal. Managing shark fishermen in southern Australia: a critique. Marine Policy, 32(3):263–273, May 2008. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X07000711.

delaMare:2005:MEB

[dlM05] William K. de la Mare. Marine ecosystem-based management as a hierarchical control system. Marine Policy, 29(1):57–68, January 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000181.

Doloreux:2009:ISO

[DM09] David Doloreux and Yannik Melançon. Innovation-support organizations in the marine science and technology industry: the case of Quebec’s coastal region in Canada. Marine Policy, 33(1): 90–100, January 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08000808.

Douvere:2007:RMS

[DMVS07] F. Douvere, F. Maes, A. Vanhulle, and J. Schrijvers. The role of marine spatial planning in sea use management: the Belgian case. Marine Policy, 31(2):182–191, March 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000765.

Douvere:2008:IMS

[Dou08] Fanny Douvere. The importance of marine spatial planning in advancing ecosystem-based sea use management. Marine Policy, 32(5):762–771, September 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X0800064X.

Dyer:2000:NRR

[DP00] Christopher L. Dyer and John J. Poggie. The Natural Resource Region and marine policy: a case study from the New England Groundfish Fishery. Marine Policy, 24(3):245–255, May 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X99000214.

Day:2008:MPF

[DPE^{+}08] Val Day, Rosemary Paxinos, Jon Emmett, Alison Wright, and Meg Goecker. The Marine Planning Framework for South Australia: a new ecosystem-based zoning policy for marine management. Marine Policy, 32(4):535–543, July 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001212.

Dromgoole:2004:MWG

[Dro04] Sarah Dromgoole. Murky waters for government policy: the case of a 17th century British warship and 10 tonnes of gold coins. Marine Policy, 28(3):189–198, May 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X03001039.

daSilva:2004:CPC

[dS04] Patricia Pinto da Silva. From common property to comanagement: lessons from Brazil’s first maritime extractive reserve. Marine Policy, 28(5):419–428, September 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X03001337.

Doloreux:2009:MCD

[DS09] David Doloreux and Richard Shearmur. Maritime clusters in diverse regional contexts: the case of Canada. Marine Policy, 33 (3):520–527, May 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08001887.

Diamond:2003:CCN

[DSH03] Nancy K. Diamond, Lesley Squillante, and Lynne Z. Hale. Cross currents: navigating gender and population linkages for integrated coastal management. Marine Policy, 27(4):325–331, July 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000447.

daSilva:2006:CFM

[dSK06] Patricia Pinto da Silva and Andrew Kitts. Collaborative fisheries management in the Northeast US: Emerging initiatives and future directions. Marine Policy, 30(6):832–841, November 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000224.

Dominguez-Torreiro:2004:CMP

[DTFÁIM04] Marcos Domínguez-Torreiro, Ana B. Freijeiro-Álvarez, and Carlos Iglesias-Malvido. Co-management proposals and their efficiency implications in fisheries management: the case of the Grand Sole fleet. Marine Policy, 28(3):213–219, May 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03001106.

deVivero:2007:EVO

[dV07] Juan Luis Suárez de Vivero. The European vision for oceans and seas — Social and political dimensions of the Green Paper on Maritime Policy for the EU. Marine Policy, 31(4):409–414, July 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06001060.

deVivero:2002:MBS

[dVM02a] Juan. L. Suárez de Vivero and Juan. C. Rodríguez Mateos. The Mediterranean and Black Sea: regional integration and maritime nationalism. Marine Policy, 26(5):383–401, September 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000209.

deVivero:2002:SSD

[dVM02b] Juan L. Suárez de Vivero and Juan C. Rodríguez Mateos. Spain and the sea. The decline of an ideology, crisis in the maritime sector and the challenges of globalization. Marine Policy, 26(3): 143–153, May 2002. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X01000446.

deVivero:2004:NFO

[dVM04] Juan Luis Suárez de Vivero and Juan Carlos Rodríguez Mateos. New factors in ocean governance. From economic to security-based boundaries. Marine Policy, 28(2):185–188, March 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000630.

deVivero:2006:MEE

[dVM06] Juan Luis Suárez de Vivero and Juan Carlos Rodríguez Mateos. Maritime Europe and EU enlargement. A geopolitical perspective. Marine Policy, 30(2):167–172, March 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04001071.

deVivero:2008:PPP

[dVMdC08] Juan L. Suárez de Vivero, Juan C. Rodríguez Mateos, and D. Florido del Corral. The paradox of public participation in fisheries governance. The rising number of actors and the devolution process. Marine Policy, 32(3):319–325, May 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000772.

deVivero:2009:GFM

[dVMdC09] Juan L. Suárez de Vivero, Juan C. Rodríguez Mateos, and David Florido del Corral. Geopolitical factors of maritime policies and marine spatial planning: State, regions, and geographical planning scope. Marine Policy, 33(4):624–634, July 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0800198X.

Degnbol:2008:SPN

[DW08] Ditte Degnbol and Douglas Clyde Wilson. Spatial planning on the North Sea: a case of cross-scale linkages. Marine Policy, 32 (2):189–200, March 2008. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X0700111X.

Djalal:2005:DGM

[DYB05] Hasjim Djalal, Alexander Yankov, and Anthony Bergin. Draft guidelines for military and intelligence gathering activities in the EEZ and their means and manner of implementation and enforcement. Marine Policy, 29(2):175–183, March 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000867.

Ebbin:2002:EFT

[Ebb02] Syma A. Ebbin. Enhanced fit through institutional interplay in the Pacific Northwest salmon co-management regime. Marine Policy, 26(4):253–259, July 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X02000076.

Ebbin:2003:SUI

[Ebb03] Syma A. Ebbin. Swimming upstream: institutional dimensions of asymmetrical problems in two salmon management regimes. Marine Policy, 27(5):441–448, September 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000071.

Ebbin:2009:IED

[Ebb09] Syma A. Ebbin. Institutional and ethical dimensions of resilience in fishing systems: Perspectives from co-managed fisheries in the Pacific Northwest. Marine Policy, 33(2):264–270, March 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0800119X.

Edwards:2008:OZF

[Edw08] Steven Edwards. Ocean zoning, first possession and Coasean contracts. Marine Policy, 32(1):46–54, January 2008. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000498.

Edwards:2009:SFO

[Edw09] Peter E. T. Edwards. Sustainable financing for ocean and coastal management in Jamaica: the potential for revenues from tourist user fees. Marine Policy, 33(2):376–385, March 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001413.

Eggert:2003:FCR

[EE03] Håkan Eggert and Anders Ellegård. Fishery control and regulation compliance: a case for co-management in Swedish commercial fisheries. Marine Policy, 27(6):525–533, November 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000782.

Ehler:2008:CBL

[Ehl08] Charles Ehler. Conclusions: Benefits, lessons learned, and future challenges of marine spatial planning. Marine Policy, 32(5):840– 843, September 2008. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08000730.

Ekstrom:2009:CCL

[Eks09] Julia A. Ekstrom. California Current Large Marine Ecosystem: Publicly available dataset of state and federal laws and regulations. Marine Policy, 33(3):528–531, May 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X0800153X.

Elvestad:2009:IMA

[Elv09] Christel Elvestad. Improving market access for salmon through free trade agreements: the role of interests and institutional frameworks. Marine Policy, 33(4):606–612, July 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001966.

Eagle:2004:WFS

[ENS04] Josh Eagle, Rosamond Naylor, and Whitney Smith. Why farm salmon outcompete fishery salmon. Marine Policy, 28(3):259– 270, May 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000885.

Eggert:2009:VMA

[EO09] Håkan Eggert and Björn Olsson. Valuing multi-attribute marine water quality. Marine Policy, 33(2):201–206, March 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001000.

Ellingsen:2009:EAN

[EOU09] H. Ellingsen, J. O. Olaussen, and I. B. Utne. Environmental analysis of the Norwegian fishery and aquaculture industry — a preliminary study focusing on farmed salmon. Marine Policy, 33 (3):479–488, May 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08001553.

Erceg:2006:DIF

[Erc06] Diane Erceg. Deterring IUU fishing through state control over nationals. Marine Policy, 30(2):173–179, March 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000023.

Elvin:2008:RWV

[ET08] S. S. Elvin and C. T. Taggart. Right whales and vessels in Canadian waters. Marine Policy, 32(3):379–386, May 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000942.

Evans:2009:UDP

[Eva09] Louisa S. Evans. Understanding divergent perspectives in marine governance in Kenya. Marine Policy, 33(5):784–793, September 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000293.

Endo:2007:PGD

[EY07] Aiko Endo and Masahiro Yamao. Policies governing the distribution of by-products from scientific and small-scale coastal whaling in Japan. Marine Policy, 31(2):169–181, March 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000753.

Ekstrom:2009:TNO

[EYG^{+}09] Julia A. Ekstrom, Oran R. Young, Steve D. Gaines, Maria Gordon, and Bonnie J. McCay. A tool to navigate overlaps in fragmented ocean governance. Marine Policy, 33(3):532–535, May 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001589.

Eythorsson:2000:DIM

[Eyt00] Einar Eythórsson. A decade of ITQ-management in Icelandic fisheries: consolidation without consensus. Marine Policy, 24(6): 483–492, November 2000. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X0000021X.

Frost:2006:CFP

[FA06] Hans Frost and Peder Andersen. The Common Fisheries Policy of the European Union and fisheries economics. Marine Policy, 30(6):737–746, November 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X06000029.

Fabinyi:2008:DTF

[Fab08] Michael Fabinyi. Dive tourism, fishing and marine protected areas in the Calamianes Islands, Philippines. Marine Policy, 32(6): 898–904, November 2008. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08000158.

Fahy:2008:PIF

[Fah08] Edward Fahy. Performance of an inshore fishery in the absence of regulatory enforcement. Marine Policy, 32(6):1037– 1042, November 2008. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X0800033X.

Fletcher:2007:HEP

[FBH^{+}07] Stephen Fletcher, Emma Beagley, Tracey Hewett, Alan Williams, and Karen McHugh. The Hamble Estuary Partnership and Solent Forum: Duplication or integration? Marine Policy, 31(5):619– 627, September 2007. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X07000255.

Fletcher:2003:UVL

[FD03] Stephen Fletcher and Wendy Dodds. The use of a virtual learning environment to enhance ICM capacity building. Marine Policy, 27(3):241–247, May 2003. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X03000034.

Fraser:2009:CNA

[FE09] Gail S. Fraser and Joanne Ellis. The Canada–Newfoundland Atlantic Accord Implementation Act: Transparency of the environmental management of the offshore oil and gas industry. Marine Policy, 33(2):312–316, March 2009. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001267.

Field:2006:CEB

[FF06] John C. Field and Robert C. Francis. Considering ecosystembased fisheries management in the California Current. Marine Policy, 30(5):552–569, September 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X05000606.

Freire:2006:RSS

[FFM06] Juan Freire, Luis Fernández, and Ramón Muiño. Role of the Spanish scientific community in the initial assessment and management of the environmental damages caused by the Prestige oil spill. Marine Policy, 30(4):308–314, July 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000291.

Freire:2000:SBC

[FGA00] Juan Freire and Antonio García-Allut. Socioeconomic and biological causes of management failures in European artisanal fisheries: the case of Galicia (NW Spain). Marine Policy, 24(5):375– 384, September 2000. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X00000130.

Flaaten:2004:FVP

[FH04] Ola Flaaten and Knut Heen. Fishing vessel profitability and local economic link obligations — the case of Norwegian trawlers. Marine Policy, 28(6):451–457, November 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03001350.

Foster:2005:IIO

[FHCS05] Elizabeth Foster, Marcus Haward, and Scott Coffen-Smout. Implementing integrated oceans management: Australia’s south east regional marine plan (SERMP) and Canada’s eastern Scotian shelf integrated management (ESSIM) initiative. Marine Policy, 29(5):391–405, September 2005. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000594.

Fina:2005:RBS

[Fin05] Mark Fina. Rationalization of the Bering Sea and Aleutian Islands crab fisheries. Marine Policy, 29(4):311–322, July 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000508.

Fletcher:2007:CMSa

[FJH07] Stephen Fletcher, David Johnson, and Tracey Hewett. Coastal management in the Solent: an introduction. Marine Policy, 31(5):585–590, September 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X07000206.

Floysand:2001:GLC

[FL01] Arnt Fløysand and Knut Bjørn Lindkvist. Globalisation, local capitalism and fishery communities in change. Marine Policy, 25 (2):113–121, March 2001. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X00000427.

Fletcher:2000:ECM

[Fle00] Stephen Fletcher. The evolution of coastal management policy in the State of Israel. Marine Policy, 24(5):395–405, September 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X00000154.

Fletcher:2001:ELT

[Fle01] Stephen Fletcher. Empowering learning through design: a comparative analysis of ICM course development for web-based and classroom delivery. Marine Policy, 25(6):457–466, November 2001. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X01000306.

Fletcher:2003:SRD

[Fle03] Stephen Fletcher. Stakeholder representation and the democratic basis of coastal partnerships in the UK. Marine Policy, 27(3): 229–240, May 2003. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X02000854.

Fafaliou:2006:ESI

[FLT06] Irene Fafaliou, Maria Lekakou, and Ioannis Theotokas. Is the European shipping industry aware of corporate social responsibility? The case of the Greek-owned short sea shipping companies. Marine Policy, 30(4):412–419, July 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000308.

Fanning:2007:LME

[FMM^{+}07] Lucia Fanning, Robin Mahon, Patrick McConney, Jorge Angulo, Felicity Burrows, Bisessar Chakalall, Diego Gil, Milton Haughton, Sherry Heileman, Sergio Martínez, L’ouverture Ostine, Adrian Oviedo, Scott Parsons, Terrence Phillips, Claudia Santizo Arroya, Bertha Simmons, and Cesar Toro. A large marine ecosystem governance framework. Marine Policy, 31(4): 434–443, July 2007. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X07000097.

Frangoudes:2008:OAC

[FMPPF08] Katia Frangoudes, Begoña Marugán-Pintos, and José J. Pascual-Fernández. From open access to co-governance and conservation: the case of women shellfish collectors in Galicia (Spain). Marine Policy, 32(2):223–232, March 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001145.

Flannery:2008:MSP

[FÓ08] Wesley Flannery and Micheál Ó Cinnéide. Marine spatial planning from the perspective of a small seaside community in Ireland. Marine Policy, 32(6):980–987, November 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000274.

Fock:2008:FCM

[Foc08] Heino O. Fock. Fisheries in the context of marine spatial planning: Defining principal areas for fisheries in the German EEZ. Marine Policy, 32(4):728–739, July 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X0800002X.

Felthoven:2004:DPM

[FP04] Ronald G. Felthoven and Catherine J. Morrison Paul. Directions for productivity measurement in fisheries. Marine Policy, 28(2): 161–169, March 2004. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X03000848.

Fletcher:2007:CMSb

[FP07] Stephen Fletcher and Kate Pike. Coastal management in the Solent: the stakeholder perspective. Marine Policy, 31(5):638– 644, September 2007. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X07000280.

Fletcher:2009:PAM

[FPHP09] Stephen Fletcher, Jonathan S. Potts, Carolyn Heeps, and Kate Pike. Public awareness of marine environmental issues in the UK. Marine Policy, 33(2):370–375, March 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001401.

Frid:2005:EBF

[FPS05] Chris Frid, Odette Paramor, and Catherine Scott. Ecosystembased fisheries management: progress in the NE Atlantic. Marine Policy, 29(5):461–469, September 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X0400065X.

Frantzi:2008:WDI

[Fra08] Sofia Frantzi. What determines the institutional performance of environmental regimes?: a case study of the Mediterranean Action Plan. Marine Policy, 32(4):618–629, July 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001315.

Froese:2009:NSC

[FSPKR09] Rainer Froese, Amanda Stern-Pirlot, and Kathleen Kesner-Reyes. Out of new stocks in 2020: a comment on “Not all fisheries will be collapsed in 2048”. Marine Policy, 33(1):180–181, January 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000596.

Fletcher-Tomenius:2000:HWI

[FTF00] Paul Fletcher-Tomenius and Craig Forrest. Historic wreck in international waters: conflict or consensus? Marine Policy, 24 (1):1–10, January 2000. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X99000196.

Faure:2008:FCO

[FW08] Michael Faure and Hui Wang. Financial caps for oil pollution damage: a historical mistake? Marine Policy, 32(4):592–606, July 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001297.

Goodhead:2007:HMP

[GA07] Tim Goodhead and Zeynep Aygen. Heritage management plans and integrated coastal management. Marine Policy, 31(5):607– 610, September 2007. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X07000231.

Gaichas:2008:CEB

[Gai08] Sarah K. Gaichas. A context for ecosystem-based fishery management: Developing concepts of ecosystems and sustainability. Marine Policy, 32(3):393–401, May 2008. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000966.

Gilman:2008:SIP

[GCB^{+}08] Eric Gilman, Shelley Clarke, Nigel Brothers, Joanna Alfaro-Shigueto, John Mandelman, Jeff Mangel, Samantha Petersen, Susanna Piovano, Nicola Thomson, Paul Dalzell, Miguel Donoso, Meidad Goren, and Tim Werner. Shark interactions in pelagic longline fisheries. Marine Policy, 32(1):1–18, January 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000565.

Gelcich:2009:MEB

[GDI^{+}09] Stefan Gelcich, Omar Defeo, Oscar Iribarne, Graciano Del Carpio, Random DuBois, Sebastian Horta, Juan Pablo Isacch, Natalio Godoy, Pastor Coayla Peñaloza, and Juan Carlos Castilla. Marine ecosystem-based management in the Southern Cone of South America: Stakeholder perceptions and lessons for implementation. Marine Policy, 33(5):801–806, September 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000311.

Gilman:2006:FCA

[GDM06] Eric L. Gilman, Paul Dalzell, and Sean Martin. Fleet communication to abate fisheries bycatch. Marine Policy, 30(4):360–366, July 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000394.

Gezelius:2006:MFM

[Gez06] Stig S. Gezelius. Monitoring fishing mortality: Compliance in Norwegian offshore fisheries. Marine Policy, 30(5):462–469, September 2006. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X05000412.

Gezelius:2008:MRP

[Gez08] Stig S. Gezelius. Management responses to the problem of incidental catch in fishing: a comparative analysis of the EU, Norway, and the Faeroe Islands. Marine Policy, 32(3):360–368, May 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000929.

Garza-Gil:2008:PAG

[GGAD08] M. Dolores Garza-Gil and Lucy Amigo-Dobaño. The profitability of the artisanal Galician fleet. Marine Policy, 32(1):74–78, January 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000528.

Garza-Gil:2006:AED

[GGSRVL06] M. Dolores Garza-Gil, Juan C. Surís-Regueiro, and Manuel M. Varela-Lafuente. Assessment of economic damages from the Prestige oil spill. Marine Policy, 30(5):544–551, September 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0500059X.

Garza-Gil:2005:PSS

[GGVL05] Ma. Dolores Garza-Gil and Manuel M. Varela-Lafuente. The profitability of the Spanish swordfish fleet in the North Atlantic. Marine Policy, 29(6):533–538, November 2005. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X04001058.

Garza-Gil:2009:SGH

[GGVL09] M. Dolores Garza-Gil and Manuel M. Varela-Lafuente. Is the Spanish Greenland halibut fleet profitable after introducing the biological recovery plan? Marine Policy, 33(2):254–257, March 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001061.

Garza-Gil:2003:SNA

[GGVLIM03] M. Dolores Garza-Gil, Manuel M. Varela-Lafuente, and Carlos Iglesias-Malvido. Spain’s North Atlantic swordfish fishery. Marine Policy, 27(1):31–37, January 2003. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000489.

Gray:2003:RCF

[GH03] Tim Gray and Jenny Hatchard. The 2002 reform of the Common Fisheries Policy’s system of governance — rhetoric or reality? Marine Policy, 27(6):545–554, November 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000666.

Gray:2008:CRS

[GH08] Tim Gray and Jenny Hatchard. A complicated relationship: Stakeholder participation and the ecosystem-based approach to fisheries management. Marine Policy, 32(2):158–168, March 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0700108X.

Grafton:2008:PFC

[GHR^{+}08] R. Quentin Grafton, Ray Hilborn, Lori Ridgeway, Dale Squires, Meryl Williams, Serge Garcia, Theodore Groves, James Joseph, Kieran Kelleher, Tom Kompas, Gary Libecap, Carl Gustaf Lundin, Mitsutaku Makino, Thorolfur Matthiasson, Richard McLoughlin, Ana Parma, Gustavo San Martin, Ben Satia, Carl-Christian Schmidt, Maree Tait, and Lin Xiu Zhang. Positioning fisheries in a changing world. Marine Policy, 32(4):630–634, July 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001327.

Gibbs:2007:LKC

[Gib07] Mark T. Gibbs. Lesser-known consequences of managing marine fisheries using individual transferable quotas. Marine Policy, 31 (2):112–116, March 2007. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X06000479.

Gibbs:2008:NGF

[Gib08] Mark T. Gibbs. Network governance in fisheries. Marine Policy, 32(1):113–119, January 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X07000577.

Gibbs:2009:IBE

[Gib09a] Mark T. Gibbs. Implementation barriers to establishing a sustainable coastal aquaculture sector. Marine Policy, 33(1):83–89, January 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000791.

Gibbs:2009:RWI

[Gib09b] Mark T. Gibbs. Resilience: What is it and what does it mean for marine policymakers? Marine Policy, 33(2):322–331, March 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001280.

Gallagher:2004:CSC

[GJGT04] Anthony Gallagher, David Johnson, Gillian Glegg, and Colin Trier. Constructs of sustainability in coastal management. Marine Policy, 28(3):249–255, May 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X03001003.

Grafton:2005:UAA

[GK05] R. Quentin Grafton and Tom Kompas. Uncertainty and the active adaptive management of marine reserves. Marine Policy, 29(5):471–479, September 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X04000880.

Grafton:2007:BFG

[GKMR07] R. Quentin Grafton, Tom Kompas, Richard McLoughlin, and Nick Rayns. Benchmarking for fisheries governance. Marine Policy, 31(4):470–479, July 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X07000048.

Gunatilake:2003:TMB

[GL03] H. M. Gunatilake and Pin Sun Leung. Technology and management of bottomfish fisheries in Northwestern Hawaiian Islands. Marine Policy, 27(1):59–67, January 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000520.

Gonzalez-Laxe:2005:PPF

[GL05] Fernando González-Laxe. The precautionary principle in fisheries management. Marine Policy, 29(6):495–505, November 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000995.

Griggs:2007:VNU

[GL07] L. Griggs and G. Lugten. Veil over the nets (unravelling corporate liability for IUU fishing offences). Marine Policy, 31(2):159–168, March 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000741.

Gilliland:2008:KES

[GL08] Paul M. Gilliland and Dan Laffoley. Key elements and steps in the process of developing ecosystem-based marine spatial planning. Marine Policy, 32(5):787–796, September 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000675.

Glen:2008:WDW

[Gle08] David Glen. What do we know about the labour market for seafarers?: a view from the UK. Marine Policy, 32(6):845– 855, November 2008. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X07001522.

Glowka:2003:PMS

[Glo03] Lyle Glowka. Putting marine scientific research on a sustainable footing at hydrothermal vents. Marine Policy, 27(4):303–312, July 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000423.

Gales:2008:JWH

[GLP08] Nick Gales, Russell Leaper, and Vassili Papastavrou. Is Japan’s whaling humane? Marine Policy, 32(3):408–412, May 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0700098X.

Glen:2001:AED

[GM01a] David Glen and James McConville. An analysis of the employment durations of UK seafaring officers 1999. Marine Policy, 25 (4):293–301, July 2001. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X01000136.

Glen:2001:ECU

[GM01b] David Glen and James McConville. Employment characteristics of UK seafaring officers in 1999. Marine Policy, 25(3):215–221, May 2001. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X01000124.

Grafton:2009:EAE

[GM09] R. Q. Grafton and A. McIlgorm. Ex ante evaluation of the costs and benefits of individual transferable quotas: a case-study of seven Australian commonwealth fisheries. Marine Policy, 33(4): 714–719, July 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X09000219.

Gardner:2007:PIM

[GMN^{+}07] B. M. Gardner, P. B. Marlow, M. M. Naim, R. Nair, and S. J. Pettit. The policy implications of market failure for the landbased jobs market for British seafarers. Marine Policy, 31(2): 117–124, March 2007. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X06000480.

Goss:2008:SRS

[Gos08] Richard Goss. Social responsibility in shipping. Marine Policy, 32(1):142–146, January 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X07000619.

Guillotreau:2000:TBE

[GP00] P. Guillotreau and N. Péridy. Trade barriers and European imports of seafood products: a quantitative assessment. Marine Policy, 24(5):431–437, September 2000. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X0000018X.

Green:2001:AMB

[Gre01] Julia Green. Australian maritime boundaries: the Australian Antarctic Territory. Marine Policy, 25(1):1–11, January 2001. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X00000282.

Grynberg:2003:WFS

[Gry03] Roman Grynberg. WTO fisheries subsidies negotiations: implications for fisheries access arrangements and sustainable management. Marine Policy, 27(6):499–511, November 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000642.

Guyader:2001:DIO

[GT01] Olivier Guyader and Olivier Thébaud. Distributional issues in the operation of rights-based fisheries management systems. Marine Policy, 25(2):103–112, March 2001. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X00000415.

Giraud:2002:EBP

[GTLC02] Kelly Giraud, Branka Turcin, John Loomis, and Joseph Cooper. Economic benefit of the protection program for the Steller sea lion. Marine Policy, 26(6):451–458, November 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000258.

Guillotreau:2004:HDE

[Gui04] Patrice Guillotreau. How does the European seafood industry stand after the revolution of salmon farming: an economic analysis of fish prices. Marine Policy, 28(3):227–233, May 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03001088.

Gulbrandsen:2009:EEM

[Gul09] Lars H. Gulbrandsen. The emergence and effectiveness of the Marine Stewardship Council. Marine Policy, 33(4):654–660, July 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000049.

Haas:2000:RCP

[Haa00a] P. M. Haas. Reply to comment on “Prospects for effective marine governance in the Northwest Pacific region”. Marine Policy, 24(6):499–500, November 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X00000245.

Haas:2000:PEM

[Haa00b] Peter M. Haas. Prospects for effective marine governance in the NW Pacific region. Marine Policy, 24(4):341–348, July 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X00000105.

Halim:2003:PAG

[Hal03] Abdul Halim. A prospect for adoption of grouper mariculture in Indonesia. Marine Policy, 27(2):159–167, March 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000878.

Hamzah:2003:IRD

[Ham03] B. A. Hamzah. International rules on decommissioning of offshore installations: some observations. Marine Policy, 27(4):339–348, July 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0300040X.

Hannesson:2003:AF

[Han03] Rögnvaldur Hannesson. Aquaculture and fisheries. Marine Policy, 27(2):169–178, March 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X02000830.

Hannesson:2007:CAC

[Han07a] Rögnvaldur Hannesson. Cheating about the cod. Marine Policy, 31(6):698–705, November 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X07000188.

Hannesson:2007:GDF

[Han07b] Rögnvaldur Hannesson. Geographical distribution of fish catches and temperature variations in the northeast Atlantic since 1945. Marine Policy, 31(1):32–39, January 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000431.

Hannesson:2007:I

[Han07c] Rögnvaldur Hannesson. Introduction. Marine Policy, 31(1):1–4, January 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000388.

Hannesson:2008:EEZ

[Han08] Rögnvaldur Hannesson. The exclusive economic zone and economic development in the Pacific island countries. Marine Policy, 32(6):886–897, November 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X08000080.

Harte:2001:OBI

[Har01] Michael Harte. Opportunities and barriers for industryled fisheries research. Marine Policy, 25(2):159–167, March 2001. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X01000069.

Harte:2007:FCF

[Har07] Michael Harte. Funding commercial fisheries management: Lessons from New Zealand. Marine Policy, 31(4):379–389, July 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06001254.

Hara:2009:CMS

[Har09] Mafaniso Hara. Crew members in South Africa’s squid industry; whether they have benefitted from transformation and governance reforms. Marine Policy, 33(3):513–519, May 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001681.

Hatcher:2000:SEF

[Hat00] Aaron Hatcher. Subsidies for European fishing fleets: the European Community’s structural policy for fisheries 1971– 1999. Marine Policy, 24(2):129–140, March 2000. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X9900024X.

Hauck:2008:RSS

[Hau08] Maria Hauck. Rethinking small-scale fisheries compliance. Marine Policy, 32(4):635–642, July 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X07001339.

Hayashi:2005:MIG

[Hay05] Moritaka Hayashi. Military and intelligence gathering activities in the EEZ: definition of key terms. Marine Policy, 29(2): 123–137, March 2005. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X04000806.

Haward:2000:TDW

[HB00] Marcus Haward and Anthony Bergin. Taiwan’s distant water tuna fisheries. Marine Policy, 24(1):33–43, January 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X99000056.

Haward:2001:PEJ

[HB01] Marcus Haward and Anthony Bergin. The political economy of Japanese distant water tuna fisheries. Marine Policy, 25(2): 91–101, March 2001. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X00000385.

Harte:2007:RMC

[HB07a] Michael Harte and John Barton. Reforming management of commercial fisheries in a small island territory. Marine Policy, 31(4): 371–378, July 2007. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X06001242.

Heazle:2007:FDS

[HB07b] Michael Heazle and John G. Butcher. Fisheries depletion and the state in Indonesia: Towards a regional regulatory regime. Marine Policy, 31(3):276–286, May 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000844.

Humphrey:2000:ULC

[HBB00] Sarah Humphrey, Peter Burbridge, and Caroline Blatch. US lessons for coastal management in the European Union. Marine Policy, 24(4):275–286, July 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X00000038.

Hughey:2000:SGF

[HCK00] Kenneth F. D. Hughey, Ross Cullen, and Geoffrey N. Kerr. Stakeholder groups in fisheries management. Marine Policy, 24(2): 119–127, March 2000. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X99000184.

Hills:2009:LSA

[HCL^{+}09] Jeremy Hills, Margaret Carlisle, Martin Le Tissier, David Muir, and Caroline Robinson. Landscape-scale analysis of ecosystem risk and returns: a new tool for ICZM. Marine Policy, 33(6): 887–900, November 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X09000633.

Heazle:2004:SUI

[Hea04] Michael Heazle. Scientific uncertainty and the International Whaling Commission: an alternative perspective on the use of science in policy making. Marine Policy, 28(5):361–374, September 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03001179.

Heazle:2006:LPI

[Hea06] Michael Heazle. Lessons in precaution: the International Whaling Commission experience with precautionary management. Marine Policy, 30(5):496–509, September 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000552.

Heaslip:2008:MSA

[Hea08] Robyn Heaslip. Monitoring salmon aquaculture waste: the contribution of First Nations’ rights, knowledge, and practices in British Columbia, Canada. Marine Policy, 32(6):988–996, November 2008. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08000286.

Harte:2008:DUF

[HESM08] Michael Harte, Polly Endreny, Gil Sylvia, and Heather Munro Mann. Developing underutilized fisheries: Oregon’s developmental fisheries program. Marine Policy, 32(4):643–652, July 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001340.

Hewett:2007:IRB

[Hew07] Tracey Hewett. Implications of the revision to the bathing water directive for local authorities in the Solent. Marine Policy, 31(5):628–631, September 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X07000267.

Halfar:2002:PMD

[HF02] Jochen Halfar and Rodney M. Fujita. Precautionary management of deep-sea mining. Marine Policy, 26(2):103–106, March 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X01000410.

Hatcher:2002:QHF

[HFPR02] Aaron Hatcher, Julian Frere, Sean Pascoe, and Kate Robinson. “Quota-hopping’’ and the foreign ownership of UK fishing vessels. Marine Policy, 26(1):1–11, January 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X01000355.

Holland:2001:CPI

[HG01] Daniel S. Holland and Jay J. C. Ginter. Common property institutions in the Alaskan groundfish fisheries. Marine Policy, 25 (1):33–42, January 2001. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X00000336.

Hersoug:2000:CRI

[HH00] Bjorn Hersoug and Petter Holm. Change without redistribution: an institutional perspective on South Africa’s new fisheries policy. Marine Policy, 24(3):221–231, May 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X99000238.

Hyvattinen:2004:EPM

[HH04] Heli Hyvättinen and Mikael Hildén. Environmental policies and marine engines — effects on the development and adoption of innovations. Marine Policy, 28(6):491–502, November 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000028.

Herrera:2006:CWT

[HH06] Guillermo E. Herrera and Porter Hoagland. Commercial whaling, tourism, and boycotts: an economic perspective. Marine Policy, 30(3):261–269, May 2006. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X05000072.

Hommes:2009:RPK

[HHM^{+}09] Saskia Hommes, Suzanne J. M. H. Hulscher, Jan P. M. Mulder, Henriëtte S. Otter, and Hans Th. A. Bressers. Role of perceptions and knowledge in the impact assessment for the extension of Mainport Rotterdam. Marine Policy, 33(1):146–155, January 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000948.

Hersoug:2000:MPD

[HHR00] Bjørn Hersoug, Petter Holm, and Stein Arne Rånes. The missing T. Path dependency within an individual vessel quota system — the case of Norwegian cod fisheries. Marine Policy, 24(4): 319–330, July 2000. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X00000087.

Hilborn:2007:DSF

[Hil07] Ray Hilborn. Defining success in fisheries and conflicts in objectives. Marine Policy, 31(2):153–158, March 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000613.

Hinds:2001:PIR

[Hin01] Lennox Hinds. Policy implications and reconstruction of marine institutional architectures: the regional indigenous organization (RIO) option. Marine Policy, 25(6):415–426, November 2001. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X01000264.

Hinds:2003:OGI

[Hin03] Lennox Hinds. Oceans governance and the implementation gap. Marine Policy, 27(4):349–356, July 2003. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000393.

Honneland:2002:IIF

[HJ02] Geir Hønneland and Anne-Kristin Jørgensen. Implementing international fisheries agreements in Russia — lessons from the northern basin. Marine Policy, 26(5):359–367, September 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000179.

Hauck:2006:FCS

[HK06a] Maria Hauck and Marcel Kroese. Fisheries compliance in South Africa: a decade of challenges and reform 1994– 2004. Marine Policy, 30(1):74–83, January 2006. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000473.

Hoel:2006:ASN

[HK06b] Alf Håkon Hoel and Ingrid Kvalvik. The allocation of scarce natural resources: the case of fisheries. Marine Policy, 30(4): 347–356, July 2006. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X05000370.

Haughton:2004:ECR

[HMM^{+}04] Milton O. Haughton, Robin Mahon, Patrick McConney, G. André Kong, and Anthony Mills. Establishment of the Caribbean Regional Fisheries Mechanism. Marine Policy, 28 (4):351–359, July 2004. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X03000897.

Herrick:2007:MAE

[HNMB07] Samuel. F. Herrick, Jerrold. G. Norton, Janet E. Mason, and Cindy Bessey. Management application of an empirical model of sardine–climate regime shifts. Marine Policy, 31(1):71–80, January 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000418.

Ho:2009:CPA

[Ho09a] Joshua Ho. Combating piracy and armed robbery in Asia: the ReCAAP Information Sharing Centre (ISC). Marine Policy, 33 (2):432–434, March 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08001449.

Ho:2009:RAM

[Ho09b] Joshua Ho. Recovering after a maritime terrorist attack: the APEC Trade Recovery Programme. Marine Policy, 33(4):733– 735, July 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000037.

Hollup:2000:SSC

[Hol00] Oddvar Hollup. Structural and sociocultural constraints for user-group participation in fisheries management in Mauritius. Marine Policy, 24(5):407–421, September 2000. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X00000166.

Honneland:2000:CBS

[Høn00] Geir Hønneland. Compliance in the Barents Sea fisheries. How fishermen account for conformity with rules. Marine Policy, 24 (1):11–19, January 2000. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X9800058X.

Howe:2001:LCT

[How01] Vicki Howe. Local community training and education in southern Tanzania — a case study. Marine Policy, 25(6):445–455, November 2001. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0100029X.

Howlett:2004:SRC

[HR04] Michael Howlett and Jeremy Rayner. (Not so) “Smart regulation”? Canadian shellfish aquaculture policy and the evolution of instrument choice for industrial development. Marine Policy, 28 (2):171–184, March 2004. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X03000861.

Hartley:2006:EMS

[HR06] Troy W. Hartley and Robert A. Robertson. Emergence of multistakeholder-driven cooperative research in the Northwest Atlantic: the case of the Northeast Consortium. Marine Policy, 30(5):580–592, September 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X05000709.

Holt:2008:ICA

[HRP08] Carrie A. Holt, Murray B. Rutherford, and Randall M. Peterman. International cooperation among nation-states of the North Pacific Ocean on the problem of competition among salmon for a common pool of prey resources. Marine Policy, 32(4):607–617, July 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001303.

Song:2000:CCS

[hS00] Yann huei Song. Codes of conduct in the South China Sea and Taiwan’s stand. Marine Policy, 24(6):449–459, November 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X00000233.

Hauck:2001:CFC

[HS01] Maria Hauck and Merle Sowman. Coastal and fisheries co-management in South Africa: an overview and analysis. Marine Policy, 25(3):173–185, May 2001. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X01000070.

Hentrich:2006:FMF

[HS06] Steffen Hentrich and Markus Salomon. Flexible management of fishing rights and a sustainable fisheries industry in Europe. Marine Policy, 30(6):712–720, November 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000813.

Homan:2008:OIR

[HS08] Anthony C. Homan and Todd Steiner. OPA 90’s impact at reducing oil spills. Marine Policy, 32(4):711–718, July 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001510.

Hanich:2009:MFC

[HT09] Quentin Hanich and Martin Tsamenyi. Managing fisheries and corruption in the Pacific Islands region. Marine Policy, 33(2): 386–392, March 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08001437.

Humphrey:2003:SDF

[Hum03] G. F. Humphrey. Scientific disciplines and fisheries management: a response. Marine Policy, 27(6):555–556, November 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000496.

Hunt:2003:EGI

[Hun03] Colin Hunt. Economic globalisation impacts on Pacific marine resources. Marine Policy, 27(1):79–85, January 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000556.

Hunt:2006:CSO

[Hun06] Colin Hunt. Concern for Southern Ocean fish stocks and bird populations. Marine Policy, 30(4):357–359, July 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000382.

Huntington:2009:PAT

[Hun09] Henry P. Huntington. A preliminary assessment of threats to Arctic marine mammals and their conservation in the coming decades. Marine Policy, 33(1):77–82, January 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X0800078X.

Haralambides:2003:FST

[HY03] Hercules E. Haralambides and Jiaqi Yang. A fuzzy set theory approach to flagging out: towards a new Chinese shipping policy. Marine Policy, 27(1):13–22, January 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000568.

Iliff:2008:CII

[Ili08a] Mike Iliff. Compromise in the IWC: Is it possible or desirable? Marine Policy, 32(6):997–1003, November 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000298.

Iliff:2008:IWR

[Ili08b] Mike Iliff. The international whaling regime post 2007. Marine Policy, 32(3):522–527, May 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X07001224.

Iliff:2008:MIC

[Ili08c] Mike Iliff. Modernisation of the International Convention for the Regulation of Whaling. Marine Policy, 32(3):402–407, May 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000978.

Iliff:2008:NIW

[Ili08d] Mike Iliff. Normalization of the International Whaling Commission. Marine Policy, 32(3):333–338, May 2008. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000796.

Iglesias-Malvido:2002:MSE

[IMGGVL02] Carlos Iglesias-Malvido, Dolores Garza-Gil, and Manuel Varela-Lafuente. Management systems in the EU fisheries. Marine Policy, 26(6):403–413, November 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X02000210.

Idda:2009:CEE

[IMP09] Lorenzo Idda, Fabio A. Madau, and Pietro Pulina. Capacity and economic efficiency in small-scale fisheries: Evidence from the Mediterranean Sea. Marine Policy, 33(5):860–867, September 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000499.

Isaacs:2006:SSF

[Isa06] Moenieba Isaacs. Small-scale fisheries reform: Expectations, hopes and dreams of “a better life for all”. Marine Policy, 30 (1):51–59, January 2006. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X05000497.

Jabour:2006:HLD

[Jab06] Julia Jabour. High latitude diplomacy: Australia’s Antarctic extended continental shelf. Marine Policy, 30(2):197–198, March 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000035.

Jabour:2009:ACS

[Jab09] Julia Jabour. The Australian continental shelf: Has Australia’s high-latitude diplomacy paid off?’. Marine Policy, 33(2):429– 431, March 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08001218.

Johnson:2007:CLH

[JBN07] D. E. Johnson, J. Bartlett, and L. A. Nash. Coastal lagoon habitat re-creation potential in Hampshire, England. Marine Policy, 31(5):599–606, September 2007. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X0700022X.

Jenkins:2009:HPP

[JBP09] Rhian E. Jenkins, Raymond D. H. Brown, and Michael R. Phillips. Harbour porpoise (Phocoena phocoena) conservation management: a dimensional approach. Marine Policy, 33(5):744– 749, September 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X09000220.

Jentoft:2009:FCG

[JC09a] Svein Jentoft and Ratana Chuenpagdee. Fisheries and coastal governance as a wicked problem. Marine Policy, 33(4):553–560, July 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001917.

Jones:2009:CDC

[JC09b] P. J. S. Jones and A. Carpenter. Crossing the divide: the challenges of designing an ecologically coherent and representative network of MPAs for the UK. Marine Policy, 33(5):737– 743, September 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X09000207.

Jentoft:2000:CML

[Jen00a] Svein Jentoft. The community: a missing link of fisheries management. Marine Policy, 24(1):53–60, January 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X99000093.

Jentoft:2000:LDF

[Jen00b] Svein Jentoft. Legitimacy and disappointment in fisheries management. Marine Policy, 24(2):141–148, March 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X99000251.

Jensen:2002:RFC

[Jen02] Carsten Lynge Jensen. Reduction of the fishing capacity in “common pool” fisheries. Marine Policy, 26(3):155–158, May 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X01000379.

Jentoft:2004:IFW

[Jen04] Svein Jentoft. Institutions in fisheries: what they are, what they do, and how they change. Marine Policy, 28(2):137–149, March 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0300085X.

Jentoft:2005:FCM

[Jen05] Svein Jentoft. Fisheries co-management as empowerment. Marine Policy, 29(1):1–7, January 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X04000132.

Jentoft:2006:BFM

[Jen06] Svein Jentoft. Beyond fisheries management: the Phronetic dimension. Marine Policy, 30(6):671–680, November 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0500076X.

Jentoft:2007:LGI

[Jen07] Svein Jentoft. Limits of governability: Institutional implications for fisheries and coastal governance. Marine Policy, 31(4):360– 370, July 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06001266.

Jensen:2008:UAI

[Jen08] Frank Jensen. Uncertainty and asymmetric information: an overview. Marine Policy, 32(1):89–103, January 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000541.

Jones:2009:IAF

[JGU09] Estelle Jones, Tim Gray, and Chanin Umponstira. The impact of artisanal fishing on coral reef fish health in Hat Thai Mueang, Phang-nga Province, Southern Thailand. Marine Policy, 33(4): 544–552, July 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08001905.

Johnston:2007:FHT

[JHMC07] Robert J. Johnston, Daniel S. Holland, Vishwanie Maharaj, and Tammy Warner Campson. Fish harvest tags: an alternative management approach for recreational fisheries in the US Gulf of Mexico. Marine Policy, 31(4):505–516, July 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06001370.

Jefferson:2009:PSC

[JHW09] Thomas A. Jefferson, Samuel K. Hung, and Bernd Würsig. Protecting small cetaceans from coastal development: Impact assessment and mitigation experience in Hong Kong. Marine Policy, 33 (2):305–311, March 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08001255.

Jarvinen:2000:PRP

[JM00] D. Jarvinen and G. Magnusson. Public resources for private mariculture: Northeastern United States, Atlantic Canada and Scotland after NAFTA and GATT. Marine Policy, 24(1):21–32, January 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X99000044.

Jentoft:2004:VCD

[JM04] Svein Jentoft and Knut H. Mikalsen. A vicious circle? The dynamics of rule-making in Norwegian fisheries. Marine Policy, 28 (2):127–135, March 2004. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X03001027.

Johnson:2002:ESC

[Joh02] David Johnson. Environmentally sustainable cruise tourism: a reality check. Marine Policy, 26(4):261–270, July 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000088.

Johnsen:2005:EHM

[Joh05] Jahn Petter Johnsen. The evolution of the “harvest machinery”: why capture capacity has continued to expand in Norwegian fisheries. Marine Policy, 29(6):481–493, November 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000892.

Johnson:2006:CNV

[Joh06] Derek Stephen Johnson. Category, narrative, and value in the governance of small-scale fisheries. Marine Policy, 30(6):747– 756, November 2006. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X06000030.

Jones:2006:CAP

[Jon06] Peter J. S. Jones. Collective action problems posed by no-take zones. Marine Policy, 30(2):143–156, March 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X04001034.

Jones:2008:FIR

[Jon08] P. J. S. Jones. Fishing industry and related perspectives on the issues raised by no-take marine protected area proposals. Marine Policy, 32(4):749–758, July 2008. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000055.

Jones:2009:EJP

[Jon09] P. J. S. Jones. Equity, justice and power issues raised by notake marine protected area proposals. Marine Policy, 33(5):759– 765, September 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X09000268.

Jacquet:2007:RSA

[JP07] Jennifer L. Jacquet and Daniel Pauly. The rise of seafood awareness campaigns in an era of collapsing fisheries. Marine Policy, 31 (3):308–313, May 2007. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X06000972.

Jacquet:2008:TSR

[JP08] Jennifer L. Jacquet and Daniel Pauly. Trade secrets: Renaming and mislabeling of seafood. Marine Policy, 32(3):309–318, May 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000760.

Kristofersson:2006:TRB

[KA06] Dadi Kristofersson and James L. Anderson. Is there a relationship between fisheries and farming? Interdependence of fisheries, animal production and aquaculture. Marine Policy, 30(6):721– 725, November 2006. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X05000825.

Kaiser:2006:LAR

[Kai06] Mark J. Kaiser. The Louisiana artificial reef program. Marine Policy, 30(6):605–623, November 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X06000467.

Kang:2003:UNC

[Kan03] Joon-Suk Kang. The United Nation convention on the law of the sea and fishery relations between Korea, Japan and China. Marine Policy, 27(2):111–124, March 2003. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000842.

Kaplan:2000:SAM

[Kap00] Ilene M. Kaplan. Seafood auctions, market equity and the buying and selling of fish: lessons on co-management from New England and the Spanish Mediterranean. Marine Policy, 24(2):165–177, March 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X99000287.

Karlsen:2001:CFH

[Kar01] Geir Runar Karlsen. Can formalisation help? The introduction of fisheries co-management in the inshore fisheries of Dingle, Co. Kerry, Ireland. Marine Policy, 25(1):83–89, January 2001. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X00000403.

Kaye:2004:IHS

[Kay04] Stuart Kaye. Implementing high seas biodiversity conservation: global geopolitical considerations. Marine Policy, 28(3):221–226, May 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0300109X.

Kitts:2007:ECM

[KdSR07] Andrew Kitts, Patricia Pinto da Silva, and Barbara Rountree. The evolution of collaborative management in the Northeast USA tilefish fishery. Marine Policy, 31(2):192–200, March 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000777.

Kitts:2003:CUF

[KE03] Andrew W. Kitts and Steven F. Edwards. Cooperatives in US fisheries: realizing the potential of the fishermen’s collective marketing act. Marine Policy, 27(5):357–366, September 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000502.

Kearney:2001:FPR

[Kea01] Robert E. Kearney. Fisheries property rights and recreational/ commercial conflict: implications of policy developments in Australia and New Zealand. Marine Policy, 25(1):49–59, January 2001. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0000035X.

Keyuan:2001:CEE

[Key01] Zou Keyuan. China’s exclusive economic zone and continental shelf: developments, problems, and prospects. Marine Policy, 25 (1):71–81, January 2001. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X00000373.

Keyuan:2003:SJJ

[Key03] Zou Keyuan. Sino-Japanese joint fishery management in the East China Sea. Marine Policy, 27(2):125–142, March 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000866.

Kaczynski:2002:EPW

[KF02] Vlad M. Kaczynski and David L. Fluharty. European policies in West Africa: who benefits from fisheries agreements? Marine Policy, 26(2):75–93, March 2002. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X01000392.

Knapp:2007:EAE

[KF07] Sabine Knapp and Philip Hans Franses. Econometric analysis on the effect of port state control inspections on the probability of casualty: Can targeting of substandard ships for inspections be improved? Marine Policy, 31(4):550–563, July 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06001230.

Knapp:2008:EAD

[KF08] Sabine Knapp and Philip Hans Franses. Econometric analysis to differentiate effects of various ship safety inspections. Marine Policy, 32(4):653–662, July 2008. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001352.

Knapp:2009:DRM

[KF09] Sabine Knapp and Philip Hans Franses. Does ratification matter and do major conventions improve safety and decrease pollution in shipping? Marine Policy, 33(5):826–846, September 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000475.

Korda:2008:FDU

[KHG08] R. C. Korda, J. M. Hills, and T. S. Gray. Fishery decline in Utila: Disentangling the web of governance. Marine Policy, 32(6):968– 979, November 2008. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08000262.

Kim:2002:TYA

[Kim02] Inho Kim. Ten years after the enactment of the Oil Pollution Act of 1990: a success or a failure. Marine Policy, 26(3):197–207, May 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000027.

Kim:2003:CBI

[Kim03a] Inho Kim. A comparison between the international and US regimes regulating oil pollution liability and compensation. Marine Policy, 27(3):265–279, May 2003. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000058.

Kim:2003:CLS

[Kim03b] Sun Pyo Kim. The UN convention on the law of the sea and new fisheries agreements in North East Asia. Marine Policy, 27 (2):97–109, March 2003. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X02000829.

Kim:2004:RKP

[Kim04] Duk-Ki Kim. A Republic of Korea perspective. Marine Policy, 28 (1):13–18, January 2004. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X03001222.

King:2005:SMS

[Kin05] John King. The security of merchant shipping. Marine Policy, 29 (3):235–245, May 2005. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X0400034X.

Kim:2005:KP

[kK05] Duk ki Kim. A Korean perspective. Marine Policy, 29(2): 157–161, March 2005. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X04000843.

Kaiser:2008:IHS

[KK08] Mark J. Kaiser and Richard A. Kasprzak. The impact of the 2005 hurricane season on the Louisiana Artificial Reef Program. Marine Policy, 32(6):956–967, November 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000250.

Kawabe:2009:ESD

[KKI^{+}09] Midori Kawabe, Hiroshi Kohno, Takashi Ishimaru, Osamu Baba, Naho Horimoto, Reiko Ikeda, Jota Kanda, Takafumi Kudo, Masaji Matsuyama, Masato Moteki, Yayoi Oshima, and Tsuyoshi Sasaki. Education for sustainable development for Tokyo Bay: Developing a practice framework of university-based coastal ESD. Marine Policy, 33(4):720–725, July 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000232.

Kaplan:2000:SSF

[KKP00] I. M. Kaplan and H. L. Kite-Powell. Safety at sea and fisheries management: fishermen’s attitudes and the need for co-management. Marine Policy, 24(6):493–497, November 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X00000269.

Knapp:2008:EAS

[KKR08] Sabine Knapp, Shashi N. Kumar, and Anna Bobo Remijn. Econometric analysis of the ship demolition market. Marine Policy, 32(6):1023–1036, November 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X08000328.

Kaplan:2004:CRC

[KM04] Ilene M. Kaplan and Bonnie J. McCay. Cooperative research, co-management and the social dimension of fisheries science and management. Marine Policy, 28(3):257–258, May 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000903.

Kaiser:2009:UCN

[KM09] Mark J. Kaiser and Jimmie Martin. An update on the cost of net-trawling operations in the Gulf of Mexico. Marine Policy, 33 (1):58–63, January 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08000766.

Kaiser:2004:BCS

[KP04a] Mark J. Kaiser and Allan G. Pulsipher. A binary choice severance selection model for the removal of offshore structures in the Gulf of Mexico. Marine Policy, 28(2):97–115, March 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000824.

Kaiser:2004:PVI

[KP04b] Mark J. Kaiser and Allan G. Pulsipher. The potential value of improved ocean observation systems in the Gulf of Mexico. Marine Policy, 28(6):469–489, November 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X03001520.

Knowler:2009:APS

[KPN^{+}09] D. Knowler, N. Philcox, S. Nathan, W. Delamare, W. Haider, and K. Gupta. Assessing prospects for shrimp culture in the Indian Sundarbans: a combined simulation modelling and choice experiment approach. Marine Policy, 33(4):613–623, July 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001978.

Kalikoski:2004:CFC

[KS04] Daniela C. Kalikoski and Theresa (Terre) Satterfield. On crafting a fisheries co-management arrangement in the estuary of Patos Lagoon (Brazil): opportunities and challenges faced through implementation. Marine Policy, 28(6):503–522, November 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000107.

Kamishiro:2009:PAO

[KS09] Norihiro Kamishiro and Toru Sato. Public acceptance of the oceanic carbon sequestration. Marine Policy, 33(3):466–471, May 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001528.

King:2007:SGL

[KTH^{+}07] Matthew Wilburn King, James Tobey, Jill Hepp, Stephen Olsen, and Barry Costa-Pierce. Sea Grant in Latin America?: Adapting the US Sea Grant model of linked applied research, extension, and education to a Latin American Context — Is there a fit? Marine Policy, 31(3):229–238, May 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06001047.

Kwak:2005:RMI

[KYC05] Seung-Jun Kwak, Seung-Hoon Yoo, and Jeong-In Chang. The role of the maritime industry in the Korean national economy: an input–output analysis. Marine Policy, 29(4):371–383, July 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000570.

Lane:2008:SCG

[Lan08] Marcus B. Lane. Strategic coastal governance issues in Fiji: the challenges of integration. Marine Policy, 32(6):856–866, November 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000043.

Laxe:2006:TFR

[Lax06] Fernando González Laxe. Transferability of fishing rights: the Spanish case. Marine Policy, 30(4):379–388, July 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000527.

Lobe:2004:PSC

[LB04] Kenton Lobe and Fikret Berkes. The padu system of communitybased fisheries management: change and local institutional innovation in south India. Marine Policy, 28(3):271–281, May 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000873.

LeGallic:2006:EAI

[LC06] Bertrand Le Gallic and Anthony Cox. An economic analysis of illegal, unreported and unregulated (IUU) fishing: Key drivers and possible solutions. Marine Policy, 30(6):689–695, November 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000783.

Lewsey:2004:ACC

[LCK04] Clement Lewsey, Gonzalo Cid, and Edward Kruse. Assessing climate change impacts on coastal infrastructure in the Eastern Caribbean. Marine Policy, 28(5):393–409, September 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03001313.

LeFloch:2008:CEP

[LDBB08] P. Le Floc’h, F. Daurès, C. Brigaudeau, and J. Bihel. A comparison of economic performance in the fisheries sector: a short- and long-term perspective. Marine Policy, 32(3):421–431, May 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001005.

Letts:2000:UFP

[Let00] David J. Letts. The use of force in patrolling Australia’s fishing zones. Marine Policy, 24(2):149–157, March 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X99000263.

Levesque:2008:IFA

[Lev08] Juan C. Levesque. International fisheries agreement: Review of the International Commission for the Conservation of Atlantic Tunas: Case study — Shark management. Marine Policy, 32 (3):528–533, May 2008. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X07001273.

Levesque:2009:CSU

[Lev09a] Juan C. Levesque. Characterization of the southeastern US black sea bass (Centropristis striata) pot commercial fishery and implications for western North Atlantic right whale (Eubalaena glacialis) management and policy. Marine Policy, 33(1):40–48, January 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000742.

Levesque:2009:ECD

[Lev09b] Juan C. Levesque. Establishment and closure of the directed Florida whiting (Menticirrhus americanus) fishery and implications for North Atlantic right whale (Eubalaena glacialis) conservation and management. Marine Policy, 33(2):233–247, March 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001048.

LeFloch:2001:ESF

[LFI01] Pascal Le Floc’h, Jacques Fuchs, and Ifremer. Economics of science in fishery sector — the European case. Marine Policy, 25 (2):133–142, March 2001. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X01000045.

Lindebo:2005:CSR

[LFL05a] Erik Lindebo, Hans Frost, and Jørgen Løkkegaard. Comments on: Suris-Reguiero, Varela-Lafuente and Iglesias-Malvido, “Effectiveness of the structural fisheries policy in the European Union”: [Marine Policy 27 (6) (2003) 535–544]. Marine Policy, 29(1):93–94, January 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X04000478.

Lindebo:2005:FCS

[LFL05b] Erik Lindebo, Hans Frost, and Jørgen Løkkegaard. Further comments on: Suris-Regueiro, Varela-Lafuente and Iglesias-Malvido, “Effectiveness of the structural fisheries policy in the European Union”. Marine Policy 2003;27(6):535–44. Marine Policy, 29(3): 289, May 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000491.

Li:2003:MCM

[Li03] Haiqing Li. Management of coastal mega-cities — a new challenge in the 21st century. Marine Policy, 27(4):333–337, July 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000459.

Lin:2006:CIE

[LL06] Bin Lin and Cherng-Yuan Lin. Compliance with international emission regulations: Reducing the air pollution from merchant vessels. Marine Policy, 30(3):220–225, May 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000138.

Llacer:2003:ORP

[Llá03] Francisco J. Montero Llácer. Open registers: past, present and future. Marine Policy, 27(6):513–523, November 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000770.

Llacer:2004:PMS

[Llá04] Francisco J. Montero Llácer. Panamanian maritime sector management. Marine Policy, 28(4):283–295, July 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X03001167.

Llacer:2005:PCM

[Llá05a] Francisco J. Montero Llácer. Panama Canal Management. Marine Policy, 29(1):25–37, January 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X04000156.

Llacer:2005:PCO

[Lla05b] Francisco J. Montero Llacer. The Panama Canal: operations and traffic. Marine Policy, 29(3):223–234, May 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000338.

Llacer:2006:PPP

[Lla06] Francisco J. Montero Llacer. Port privatization in Panama. Marine Policy, 30(5):483–495, September 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X05000540.

Lin:2007:ISI

[LLJ07] Bin Lin, Cherng-Yuan Lin, and Tain-Chyuan Jong. Investigation of strategies to improve the recycling effectiveness of waste oil from fishing vessels. Marine Policy, 31(4):415–420, July 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000140.

Lejano:2006:CDC

[LOS06] Raul P. Lejano and Alma Ocampo-Salvador. Context and differentiation: Comparative analysis of two community-based fishers’ organizations. Marine Policy, 30(6):726–736, November 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000928.

Liu:2005:SCF

[LOT05] Wen Hong Liu, Ching Hsiewn Ou, and Kuo Huan Ting. Sustainable coastal fishery development indicator system: a case of Gungliau, Taiwan. Marine Policy, 29(3):199–210, May 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000302.

Loucks:2007:PFI

[Lou07] Laura Loucks. Patterns of fisheries institutional failure and success: Experience from the Southern Gulf of St. Lawrence snow crab fishery, in Nova Scotia, Canada. Marine Policy, 31(3):320– 326, May 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000996.

Lane:2000:IAF

[LS00] Daniel E. Lane and Robert L. Stephenson. Institutional arrangements for fisheries: alternate structures and impediments to change. Marine Policy, 24(5):385–393, September 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X00000142.

Liu:2008:CFS

[LS08] Yajie Liu and Ussif Rashid Sumaila. Can farmed salmon production keep growing? Marine Policy, 32(3):497–501, May 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001182.

Lindkvist:2008:RCS

[LTX08] Knut Bjørn Lindkvist, Torbjørn Trondsen, and Jinghua Xie. Restructuring the Chinese seafood industry, global challenges and policy implications. Marine Policy, 32(3):432–441, May 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001017.

Leary:2009:MGR

[LVH^{+}09] David Leary, Marjo Vierros, Gwenaëlle Hamon, Salvatore Arico, and Catherine Monagle. Marine genetic resources: a review of scientific and commercial interest. Marine Policy, 33(2):183–194, March 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000985.

Leadbitter:2007:ESI

[LW07] Duncan Leadbitter and Trevor J. Ward. An evaluation of systems for the integrated assessment of capture fisheries. Marine Policy, 31(4):458–469, July 2007. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X0700005X.

Langley:2009:SST

[LWH^{+}09] Adam Langley, Andrew Wright, Glenn Hurry, John Hampton, Transform Aqorua, and Len Rodwell. Slow steps towards management of the world’s largest tuna fishery. Marine Policy, 33(2): 271–279, March 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08001206.

Maes:2008:ILF

[Mae08] Frank Maes. The international legal framework for marine spatial planning. Marine Policy, 32(5):797–810, September 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000687.

Mahon:2006:RCF

[Mah06] Robin Mahon. On the role of consulting in fisheries development. Marine Policy, 30(5):593–595, September 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000722.

Majanen:2007:RUC

[Maj07] Terhi Majanen. Resource use conflicts in Mabini and Tingloy, the Philippines. Marine Policy, 31(4):480–487, July 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000036.

Malone:2003:CMG

[Mal03] Thomas C. Malone. The coastal module of the Global Ocean Observing System (GOOS): an assessment of current capabilities to detect change. Marine Policy, 27(4):295–302, July 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000435.

Marshall:2001:LLS

[Mar01] Joan Marshall. Landlords, leaseholders & sweat equity: changing property regimes in aquaculture. Marine Policy, 25(5):335– 352, September 2001. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X01000203.

Marti:2004:TWE

[Mar04] Bruce E. Marti. Trends in world and extended-length cruising (1985–2002). Marine Policy, 28(3):199–211, May 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03001131.

Marti:2008:ANE

[Mar08] Bruce E. Marti. Assessing New England’s waterborne energy imports (1995–2004). Marine Policy, 32(4):740–748, July 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000031.

Mason:2003:CLO

[Mas03] Michael Mason. Civil liability for oil pollution damage: examining the evolving scope for environmental compensation in the international regime. Marine Policy, 27(1):1–12, January 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000519.

Masalu:2008:CDI

[Mas08] Desiderius C. P. Masalu. Coastal data and information management for integrated coastal management: the role of IODE. Marine Policy, 32(4):544–550, July 2008. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001236.

Matthiasson:2001:IDC

[Mat01] Thórólfur Matthíasson. The Icelandic debate on the case for a fishing fee: a non-technical introduction. Marine Policy, 25(4): 303–312, July 2001. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X01000082.

Maurstad:2002:FMW

[Mau02] Anita Maurstad. Fishing in murky waters — ethics and politics of research on fisher knowledge. Marine Policy, 26(3):159–166, May 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X01000458.

McFadden:2009:IEA

[MB09] Katherine W. McFadden and Cassandra Barnes. The implementation of an ecosystem approach to management within a federal government agency. Marine Policy, 33(1):156–163, January 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0800095X.

Mitchell:2008:CMR

[MBJI08] Ronald W. D. Mitchell, Osamu Baba, Gary Jackson, and Tatsuya Isshiki. Comparing management of recreational Pagrus fisheries in Shark Bay (Australia) and Sagami Bay (Japan): Conventional catch controls versus stock enhancement. Marine Policy, 32(1): 27–37, January 2008. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X07000462.

Myatt-Bell:2002:PPM

[MBSLP02] L. B. Myatt-Bell, M. D. Scrimshaw, J. N. Lester, and J. S. Potts. Public perception of managed realignment: Brancaster West Marsh, North Norfolk, UK. Marine Policy, 26(1):45–57, January 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X01000331.

McBeath:2004:MCB

[McB04] Jerry McBeath. Management of the commons for biodiversity: lessons from the North Pacific. Marine Policy, 28(6):523– 539, November 2004. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X04000119.

McGarvey:2003:DSF

[McG03a] Richard McGarvey. Demand-side fishery management: integrating two forms of input control. Marine Policy, 27(3):207–218, May 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000581.

McGaw:2003:AFP

[McG03b] Richard McGaw. Aboriginal fisheries policy in Atlantic Canada. Marine Policy, 27(5):417–424, September 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X0300023X.

McGoodwin:2007:ECV

[McG07] James R. McGoodwin. Effects of climatic variability on three fishing economies in high-latitude regions: Implications for fisheries policies. Marine Policy, 31(1):40–55, January 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000443.

Michler-Cieluch:2008:PCP

[MCK08] Tanja Michler-Cieluch and Gesche Krause. Perceived concerns and possible management strategies for governing ‘wind farm–mariculture integration’. Marine Policy, 32(6):1013–1022, November 2008. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08000316.

McKenna:2008:MPC

[MCO08] John McKenna, Andrew Cooper, and Anne Marie O’Hagan. Managing by principle: a critical analysis of the European principles of Integrated Coastal Zone Management (ICZM). Marine Policy, 32(6):941–955, November 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000249.

Minnegal:2008:MMB

[MD08] Monica Minnegal and Peter D. Dwyer. Mixed messages: Buying back Australia’s fishing industry. Marine Policy, 32(6):1063– 1071, November 2008. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08000377.

Mahon:2009:GPL

[MFM09] Robin Mahon, Lucia Fanning, and Patrick McConney. A governance perspective on the large marine ecosystem approach. Marine Policy, 33(2):317–321, March 2009. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001279.

Mason:2006:CPA

[MG06] Richard Mason and Warwick Gullett. Cancellation provisions in Australia’s Commonwealth-managed fisheries. Marine Policy, 30 (3):270–280, May 2006. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X05000060.

McLean:2009:LIH

[MG09] Craig McLean and Tim Gray. Liberal Intergovernmentalism, Historical Institutionalism, and British and German perceptions of the EU’s Common Fisheries Policy. Marine Policy, 33(3):458– 465, May 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001516.

Miguez:2009:IFE

[MGL09] Gonzalo Caballero Miguez, María Dolores Garza Gil, and Manuel M. Varela Lafuente. The institutional foundations of economic performance of mussel production: the Spanish case of the Galician floating raft culture. Marine Policy, 33(2):288–296, March 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001231.

Mikalsen:2007:LUG

[MHJ07] Knut H. Mikalsen, Hans-Kristian Hernes, and Svein Jentoft. Leaning on user-groups: the role of civil society in fisheries governance. Marine Policy, 31(2):201–209, March 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000789.

Miller:2007:CVT

[Mil07] Kathleen A. Miller. Climate variability and tropical tuna: Management challenges for highly migratory fish stocks. Marine Policy, 31(1):56–70, January 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X0600042X.

Mitroussi:2004:ROC

[Mit04] K. Mitroussi. The role of organisational characteristics of ship owning firms in the use of third party ship management. Marine Policy, 28(4):325–333, July 2004. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X0300112X.

Mitroussi:2008:ESE

[Mit08] Kyriaki Mitroussi. Employment of seafarers in the EU context: Challenges and opportunities. Marine Policy, 32(6):1043– 1049, November 2008. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08000341.

Mikalsen:2001:UGS

[MJ01] Knut H. Mikalsen and Svein Jentoft. From user-groups to stakeholders? The public interest in fisheries management. Marine Policy, 25(4):281–292, July 2001. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X0100015X.

Mikalsen:2003:LPH

[MJ03] Knut H. Mikalsen and Svein Jentoft. Limits to participation? On the history, structure and reform of Norwegian fisheries management. Marine Policy, 27(5):397–407, September 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000253.

Mikalsen:2008:PPF

[MJ08] Knut H. Mikalsen and Svein Jentoft. Participatory practices in fisheries across Europe: Making stakeholders more responsible. Marine Policy, 32(2):169–177, March 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001091.

Meek:2008:BRT

[MLRK08] Chanda L. Meek, Amy Lauren Lovecraft, Martin D. Robards, and Gary P. Kofinas. Building resilience through interlocal relations: Case studies of polar bear and walrus management in the Bering Strait. Marine Policy, 32(6):1080–1089, November 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000390.

MyongSop:2002:KFI

[MM02] Pak MyongSop and Joo MoonBae. Korea’s fisheries industry and government financial transfers. Marine Policy, 26(6):429– 435, November 2002. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X02000234.

Midgley:2004:PMP

[MM04] Stephen Midgley and Derek J. McGlashan. Planning and management of a proposed managed realignment project: Bothkennar, Forth Estuary, Scotland. Marine Policy, 28(5):429– 435, September 2004. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X03001349.

Makino:2005:CMJ

[MM05] Mitsutaku Makino and Hiroyuki Matsuda. Co-management in Japanese coastal fisheries: institutional features and transaction costs. Marine Policy, 29(5):441–450, September 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000636.

Mahon:2008:GFC

[MMR08] Robin Mahon, Patrick McConney, and Rathindra N. Roy. Governing fisheries as complex adaptive systems. Marine Policy, 32(1):104–112, January 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X07000553.

Makino:2009:EFC

[MMS09] Mitsutaku Makino, Hiroyuki Matsuda, and Yasunori Sakurai. Expanding fisheries co-management to ecosystem-based management: a case in the Shiretoko World Natural Heritage area, Japan. Marine Policy, 33(2):207–214, March 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001012.

Marlow:2006:LSI

[MN06] Peter Marlow and Rawindaran Nair. Liner shipping and information exchange — a European perspective. Marine Policy, 30(6): 681–688, November 2006. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X05000771.

Marlow:2008:SCI

[MN08] Peter Marlow and Rawindaran Nair. Service contracts — an instrument of international logistics supply chain: Under United States and European Union regulatory frameworks. Marine Policy, 32(3):489–496, May 2008. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001170.

Moreau:2009:OHS

[MN09] Darek T. R. Moreau and Barbara Neis. Occupational health and safety hazards in Atlantic Canadian aquaculture: Laying the groundwork for prevention. Marine Policy, 33(2):401–411, March 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001462.

Moen:2008:BBE

[Moe08] Amy E. Moen. Breaking Basel: the elements of the Basel Convention and its application to toxic ships. Marine Policy, 32(6):1053– 1062, November 2008. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08000365.

Morin:2000:FRE

[Mor00] Michel Morin. The fisheries resources in the European Union.: the distribution of TACs: principle of relative stability and quota-hopping. Marine Policy, 24(3):265–273, May 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0000004X.

Morishita:2006:MAW

[Mor06] Joji Morishita. Multiple analysis of the whaling issue: Understanding the dispute by a matrix. Marine Policy, 30(6):802– 808, November 2006. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X0600008X.

Morishita:2008:WEA

[Mor08] Joji Morishita. What is the ecosystem approach for fisheries management? Marine Policy, 32(1):19–26, January 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000450.

Mardle:2002:OFM

[MPB^{+}02] Simon Mardle, Sean Pascoe, Jean Boncoeur, Bertrand Le Gallic, Juan J. García-Hoyo, Inés Herrero, Ramon Jimenez-Toribio, Concepción Cortes, Nuria Padilla, Jesper Raakjaer Nielsen, and Christoph Mathiesen. Objectives of fisheries management: case studies from the UK, France, Spain and Denmark. Marine Policy, 26(6):415–428, November 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000222.

Maliao:2009:PCB

[MPT09] Ronald J. Maliao, Robert S. Pomeroy, and Ralph G. Turingan. Performance of community-based coastal resource management (CBCRM) programs in the Philippines: a metaanalysis. Marine Policy, 33(5):818–825, September 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000335.

Macinko:2001:FRP

[MR01] Seth Macinko and Leigh Raymond. Fish on the range: the perils of crossing conceptual boundaries in natural resource policy. Marine Policy, 25(2):123–131, March 2001. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X01000021.

Maunder:2002:IPS

[MS02] Mark N. Maunder and Paul J. Starr. Industry participation in stock assessment: the New Zealand SNA1 snapper (Pagrus auratus) fishery. Marine Policy, 26(6):481–492, November 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000465.

Murawski:2007:TMC

[Mur07] Steven A. Murawski. Ten myths concerning ecosystem approaches to marine resource management. Marine Policy, 31(6): 681–690, November 2007. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X07000292.

Moore:2009:RMM

[MWL^{+}09] Jeffrey E. Moore, Bryan P. Wallace, Rebecca L. Lewison, Ramúnas Žydelis, Tara M. Cox, and Larry B. Crowder. A review of marine mammal, sea turtle and seabird bycatch in USA fisheries and the role of policy in shaping management. Marine Policy, 33(3):435–451, May 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001486.

Mokashi:2002:SRC

[MWV02] A. J. Mokashi, J. Wang, and A. K. Vermar. A study of reliability-centred maintenance in maritime operations. Marine Policy, 26(5):325–335, September 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X02000143.

Nauen:2006:IWD

[Nau06] Cornelia E. Nauen. Implementing the WSSD decision of restoring marine ecosystems by 2015 — scientific information support in the public domain. Marine Policy, 30(5):455–461, September 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0400106X.

Nursey-Bray:2009:CMP

[NBR09] Melissa Nursey-Bray and Phillip Rist. Co-management and protected area management: Achieving effective management of a contested site, lessons from the Great Barrier Reef World Heritage Area (GBRWHA). Marine Policy, 33(1):118–127, January 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000912.

Nielsen:2006:SRD

[NC06] Jesper Raakjær Nielsen and Anne-Sofie Christensen. Sharing responsibilities in Danish fisheries management — experiences and future directions. Marine Policy, 30(2):181–188, March 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000047.

Nielsen:2004:FCM

[NDV^{+}04] Jesper Raakjær Nielsen, Poul Degnbol, K. Kuperan Viswanathan, Mahfuzuddin Ahmed, Mafaniso Hara, and Nik Mustapha Raja Abdullah. Fisheries co-management — an institutional innovation? Lessons from South East Asia and Southern Africa. Marine Policy, 28(2):151–160, March 2004. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000836.

Nielsen:2006:TSA

[NH06] Jesper Raakjær Nielsen and Mafaniso Hara. Transformation of South African industrial fisheries. Marine Policy, 30(1):43–50, January 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000503.

Nielsen:2007:BCF

[NH07] Kåre Nolde Nielsen and Petter Holm. A brief catalogue of failures: Framing evaluation and learning in fisheries resource management. Marine Policy, 31(6):669–680, November 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000425.

Nielsen:2003:AFS

[Nie03] Jesper Raakjær Nielsen. An analytical framework for studying: compliance and legitimacy in fisheries management. Marine Policy, 27(5):425–432, September 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X03000228.

Nielsen:2006:SRF

[Nie06] Max Nielsen. Supply regimes in fisheries. Marine Policy, 30(5):596–603, September 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X05000734.

Nielsen:2009:MFT

[Nie09] Max Nielsen. Modelling fish trade liberalisation: Does fish trade liberalisation result in welfare gains or losses? Marine Policy, 33 (1):1–7, January 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08000602.

Nielsen:2003:IFI

[NM03] Jesper Raakjær Nielsen and Christoph Mathiesen. Important factors influencing rule compliance in fisheries lessons from Denmark. Marine Policy, 27(5):409–416, September 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000241.

Noble:2000:ICC

[Nob00] Bram F. Noble. Institutional criteria for co-management. Marine Policy, 24(1):69–77, January 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X99000147.

Noble:2003:COF

[Nob03] Tim Noble. Co-operating in fisheries management: trials and tribulations in Scotland. Marine Policy, 27(5):433–439, September 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000216.

Nostbakken:2008:FLE

[Nøs08] Linda Nøstbakken. Fisheries law enforcement — a survey of the economic literature. Marine Policy, 32(3):293–300, May 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000747.

Novaczek:2001:FMC

[NSH01] Irene Novaczek, Juliaty Sopacua, and Ingvild Harkes. Fisheries management in Central Maluku, Indonesia, 1997–98. Marine Policy, 25(3):239–249, May 2001. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X01000161.

Neiland:2001:SAE

[NSVW01] Arthur E. Neiland, Neill Soley, Joan Baron Varley, and David J. Whitmarsh. Shrimp aquaculture: economic perspectives for policy development. Marine Policy, 25(4):265–279, July 2001. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X01000173.

Negro:2009:EEI

[NVPR09] M. C. García Negro, S. Villasante, A. Carballo Penela, and G. Rodríguez Rodríguez. Estimating the economic impact of the Prestige oil spill on the Death Coast (NW Spain) fisheries. Marine Policy, 33(1):8–23, January 2009. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000614.

Ning:2009:QEP

[NZF09] Feng Tao Ning, Chao Zhang, and Rod Fujita. Quantitative evaluation of the performance of a permit auction system in reducing bycatch of sea turtles in the Hawaii Swordfish longline fishery. Marine Policy, 33(1):101–105, January 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000894.

OHagan:2009:CGN

[OB09] A. M. O’Hagan and R. Ballinger. Coastal governance in North West Europe: an assessment of approaches to the European stocktake. Marine Policy, 33(6):912–922, November 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000621.

OMahony:2009:ARA

[OGC^{+}09] Cathal O’Mahony, Jeremy Gault, Val Cummins, Kathrin Köpke, and Darragh O’Suilleabhain. Assessment of recreation activity and its application to integrated management and spatial planning for Cork Harbour, Ireland. Marine Policy, 33(6):930– 937, November 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X09000645.

Ota:2008:FSF

[OJ08] Yoshitaka Ota and Roger Just. Fleet sizes, fishing effort and the ‘hidden’ factors behind statistics: an anthropological study of small-scale fisheries in UK. Marine Policy, 32(3):301–308, May 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000759.

Okey:2003:MER

[Oke03] Thomas A. Okey. Membership of the eight Regional Fishery Management Councils in the United States: are special interests over-represented? Marine Policy, 27(3):193–206, May 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000022.

OConnor:2009:PVP

[OLC^{+}09] M. C. O’Connor, G. Lymbery, J. A. G. Cooper, J. Gault, and J. McKenna. Practice versus policy-led coastal defence management. Marine Policy, 33(6):923–929, November 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000608.

Ovetz:2006:BLI

[Ove06] Robert Ovetz. The bottom line: an investigation of the economic, cultural and social costs of industrial longline fishing in the Pacific and the benefits of sustainable use marine protected areas. Marine Policy, 30(6):809–820, November 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000200.

Ovetz:2007:BLI

[Ove07] Robert Ovetz. The bottom line: an investigation of the economic, cultural and social costs of high seas industrial longline fishing in the Pacific and the benefits of conservation. Marine Policy, 31 (2):217–228, March 2007. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X06000960.

Parkes:2000:PFM

[Par00] Graeme Parkes. Precautionary fisheries management: the CCAMLR approach. Marine Policy, 24(2):83–91, March 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X99000159.

Pollack:2008:FSR

[PBB08] Gudrun Pollack, Augustin Berghöfer, and Uta Berghöfer. Fishing for social realities — Challenges to sustainable fisheries management in the Cape Horn Biosphere Reserve. Marine Policy, 32(2): 233–242, March 2008. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X07001157.

Patrick:2006:HPS

[PBSS06] Wesley S. Patrick, Okmyung Bin, Kurt A. Schwabe, and Peter W. Schuhmann. Hatchery programs, stock enhancement, and cost effectiveness: a case study of the Albemarle Sound/ Roanoke River stocking program 1981–1996. Marine Policy, 30 (4):299–307, July 2006. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X05000278.

Pascoe:2000:IDT

[PC00] Sean Pascoe and Louisa Coglan. Implications of differences in technical efficiency of fishing boats for capacity measurement and reduction. Marine Policy, 24(4):301–307, July 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X00000063.

Pomeroy:2008:ESM

[PD08] Robert Pomeroy and Fanny Douvere. The engagement of stakeholders in the marine spatial planning process. Marine Policy, 32(5):816–822, September 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X08000705.

Pinkerton:2009:ERH

[PE09] Evelyn Pinkerton and Danielle N. Edwards. The elephant in the room: the hidden costs of leasing individual transferable fishing quotas. Marine Policy, 33(4):707–713, July 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000190.

Pedersen:2008:CPS

[Ped08] Torbjørn Pedersen. The constrained politics of the Svalbard offshore area. Marine Policy, 32(6):913–919, November 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000171.

Perez:2009:FMT

[Per09] Arlenie Perez. Fisheries management at the tri-national border between Belize, Guatemala and Honduras. Marine Policy, 33(2): 195–200, March 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08000997.

Petersen:2002:EPI

[Pet02] Elizabeth H. Petersen. Economic policy, institutions and fisheries development in the Pacific. Marine Policy, 26(5):315– 324, September 2002. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X02000118.

Petersen:2003:CTF

[Pet03] Elizabeth Petersen. The catch in trading fishing access for foreign aid. Marine Policy, 27(3):219–228, May 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000593.

Pettit:2008:UKP

[Pet08] S. J. Pettit. United Kingdom ports policy: Changing government attitudes. Marine Policy, 32(4):719–727, July 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001534.

Pelc:2002:REO

[PF02] Robin Pelc and Rod M. Fujita. Renewable energy from the ocean. Marine Policy, 26(6):471–479, November 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000453.

Pedersen:2009:MFG

[PFS09] Søren Anker Pedersen, Heino O. Fock, and Anne F. Sell. Mapping fisheries in the German exclusive economic zone with special reference to offshore Natura 2000 sites. Marine Policy, 33 (4):571–590, July 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08001930.

Pascoe:2009:DBL

[PG09] Sean Pascoe and Tony Gibson. Do boat licences have a role in fisheries managed through individual quotas? Experience in Australian fisheries. Marine Policy, 33(2):297–304, March 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001243.

Pettit:2005:ESS

[PGM^{+}05] S. J. Pettit, B. M. Gardner, P. B. Marlow, M. M. Naim, and R. Nair. Ex-seafarers shore-based employment: the current UK situation. Marine Policy, 29(6):521–531, November 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04001010.

Pawson:2008:DMR

[PGP08] M. G. Pawson, H. Glenn, and G. Padda. The definition of marine recreational fishing in Europe. Marine Policy, 32(3):339–350, May 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000802.

Pinkerton:2008:CLM

[PJ08] Evelyn Pinkerton and Leonard John. Creating local management legitimacy. Marine Policy, 32(4):680–691, July 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001480.

Pomeroy:2001:CAS

[PKH01] Robert S. Pomeroy, Brenda M. Katon, and Ingvild Harkes. Conditions affecting the success of fisheries co-management: lessons from Asia. Marine Policy, 25(3):197–208, May 2001. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X01000100.

Phillips:2002:PPP

[PKH02] Gregory Phillips, Lorne Kriwoken, and Peter Hay. Private property and public interest in fisheries management: the Tasmanian rock lobster fishery. Marine Policy, 26(6):459–469, November 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0200026X.

Pitcher:2009:EPI

[PKS^{+}09] Tony J. Pitcher, Daniela Kalikoski, Katherine Short, Divya Varkey, and Ganapathiraju Pramod. An evaluation of progress in implementing ecosystem-based management of fisheries in 33 countries. Marine Policy, 33(2):223–232, March 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001036.

Perez-Labajos:2001:SPR

[PL01] Carlos A. Pérez-Labajos. Spending pattern of the recreational maritime sector and its impact on employment: the case of Cantabria, Spain. Marine Policy, 25(3):187–196, May 2001. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X01000094.

Pradhan:2004:MES

[PL04] Naresh C. Pradhan and PingSun Leung. Modeling entry, stay, and exit decisions of the longline fishers in Hawaii. Marine Policy, 28(4):311–324, July 2004. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X03001143.

Perez-Labajos:2008:FSP

[PL08] Carlos Perez-Labajos. Fishing safety policy and research. Marine Policy, 32(1):40–45, January 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X07000486.

Plant:2000:TST

[Pla00] G. Plant. The Turkish Straits and tanker traffic: an update. Marine Policy, 24(3):193–214, May 2000. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X99000329.

Plasman:2008:IMS

[Pla08] Ir. Cathy Plasman. Implementing marine spatial planning: a policy perspective. Marine Policy, 32(5):811–815, September 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000699.

Perez-Labajos:2004:CPC

[PLB04] Carlos Perez-Labajos and Beatriz Blanco. Competitive policies for commercial sea ports in the EU. Marine Policy, 28(6):553– 556, November 2004. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X04000326.

Paixao:2002:SWS

[PM02] A. C. Paixão and P. B. Marlow. Strengths and weaknesses of short sea shipping. Marine Policy, 26(3):167–178, May 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X01000471.

Piniella:2008:AFA

[PNN08] F. Piniella, J. P. Novalbos, and P. J. Nogueroles. Artisanal fishing in Andalusia (II): Safety and working conditions policy. Marine Policy, 32(4):551–558, July 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001248.

Pomeroy:2009:SSM

[PNT09] Robert Pomeroy, Kim Anh Thi Nguyen, and Ha Xuan Thong. Small-scale marine fisheries policy in Vietnam. Marine Policy, 33 (2):419–428, March 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08001498.

Polacheck:2002:ECP

[Pol02a] Tom Polacheck. Experimental catches and the precautionary approach: the Southern Bluefin Tuna dispute. Marine Policy, 26 (4):283–294, July 2002. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X02000106.

Polacheck:2002:WSC

[Pol02b] Tom Polacheck. Will “small core” fisheries solve the fishery management dilemma? Marine Policy, 26(5):369–371, September 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000180.

Polacheck:2006:TLC

[Pol06] Tom Polacheck. Tuna longline catch rates in the Indian Ocean: Did industrial fishing result in a 90% rapid decline in the abundance of large predatory species? Marine Policy, 30(5):470– 482, September 2006. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X05000539.

Pontecorvo:2001:SSU

[Pon01] Giulio Pontecorvo. Supply side uncertainty and the management of commercial fisheries: Peruvian Anchovetta, an illustration. Marine Policy, 25(2):169–172, March 2001. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X00000397.

Pontecorvo:2003:ISD

[Pon03] Giulio Pontecorvo. Insularity of scientific disciplines and uncertainty about supply: the two keys to the failure of fisheries management. Marine Policy, 27(1):69–73, January 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000532.

Pontecorvo:2008:NO

[Pon08] Giulio Pontecorvo. A note on “overfishing”. Marine Policy, 32(6):1050–1052, November 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X08000353.

Portman:2009:IPI

[Por09] Michelle Portman. Involving the public in the impact assessment of offshore renewable energy facilities. Marine Policy, 33 (2):332–338, March 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08001292.

Potts:2000:RDN

[Pot00] Jonathan S. Potts. Recent developments at the National Maritime Museum, UK, and their potential for capacity building. Marine Policy, 24(1):79–81, January 2000. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X99000226.

Pretes:2004:RFP

[PP04] Michael Pretes and Elizabeth Petersen. Rethinking fisheries policy in the Pacific. Marine Policy, 28(4):297–309, July 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03001155.

Pomeroy:2006:FRA

[PPB06] Robert S. Pomeroy, John E. Parks, and Cristina M. Balboa. Farming the reef: is aquaculture a solution for reducing fishing pressure on coral reefs? Marine Policy, 30(2):111–130, March 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000983.

Pomeroy:2007:FWC

[PPP^{+}07] Robert Pomeroy, John Parks, Richard Pollnac, Tammy Campson, Emmanuel Genio, Cliff Marlessy, Elizabeth Holle, Michael Pido, Ayut Nissapa, Somsak Boromthanarat, and Nguyen Thu Hue. Fish wars: Conflict and collaboration in fisheries management in Southeast Asia. Marine Policy, 31(6):645–656, November 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000413.

Pomeroy:2008:EPO

[PPP^{+}08] Robert S. Pomeroy, Michael D. Pido, John Francisco A. Pontillas, Benjamin S. Francisco, Alan T. White, Eva Marie C. Ponce De Leon, and Geronimo T. Silvestre. Evaluation of policy options for the live reef food fish trade in the province of Palawan, Western Philippines. Marine Policy, 32(1):55–65, January 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000504.

Pascoe:2009:SOP

[PPW^{+}09] Sean Pascoe, Wendy Proctor, Chris Wilcox, James Innes, Wayne Rochester, and Natalie Dowling. Stakeholder objective preferences in Australian Commonwealth managed fisheries. Marine Policy, 33(5):750–758, September 2009. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000244.

Pomeroy:2006:CDR

[PRH^{+}06] Robert S. Pomeroy, Blake D. Ratner, Stephen J. Hall, Jate Pimoljinda, and V. Vivekanandan. Coping with disaster: Rehabilitating coastal livelihoods and communities. Marine Policy, 30(6): 786–793, November 2006. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X06000066.

Pontecorvo:2001:SCF

[PS01] Giulio Pontecorvo and William E. Schrank. A small core fishery: a new approach to fisheries management. Marine Policy, 25(1): 43–48, January 2001. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X00000348.

Perry:2007:MEV

[PS07] R. Ian Perry and U. Rashid Sumaila. Marine ecosystem variability and human community responses: the example of Ghana, West Africa. Marine Policy, 31(2):125–134, March 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000492.

Piniella:2007:AFA

[PSFE07] F. Piniella, M. C. Soriguer, and M. A. Fernández-Engo. Artisanal fishing in Andalusia: a statistical study of the fleet. Marine Policy, 31(4):573–581, July 2007. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X06001059.

Queffelec:2009:IMM

[QCB09] Betty Queffelec, Valerie Cummins, and Denis Bailly. Integrated management of marine biodiversity in Europe: Perspectives from ICZM and the evolving EU Maritime Policy framework. Marine Policy, 33(6):871–877, November 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000694.

Qiu:2009:CDC

[QWJA09] Wanfei Qiu, Bin Wang, Peter J. S. Jones, and Jan C. Axmacher. Challenges in developing China’s marine protected area system. Marine Policy, 33(4):599–605, July 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001954.

Ram-Bidesi:2004:ITM

[RBT04] Vina Ram-Bidesi and Martin Tsamenyi. Implications of the tuna management regime for domestic industry development in the Pacific Island States. Marine Policy, 28(5):383–392, September 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03001301.

Rova:2001:WRF

[RC01] Carl Rova and Lars Carlsson. When regulation fails: vendace fishery in the Gulf of Bothnia. Marine Policy, 25(5):323– 333, September 2001. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X01000197.

Repetto:2001:NEF

[Rep01] Robert Repetto. A natural experiment in fisheries management. Marine Policy, 25(4):251–264, July 2001. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X01000185.

Reid-Grant:2009:FMP

[RGB09] Kimesha Reid-Grant and Mahadev G. Bhat. Financing marine protected areas in Jamaica: an exploratory study. Marine Policy, 33(1):128–136, January 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X08000924.

Rutherford:2005:IOM

[RHCS05] R. J. Rutherford, G. J. Herbert, and S. S. Coffen-Smout. Integrated ocean management and the collaborative planning process: the Eastern Scotian Shelf Integrated Management (ESSIM) Initiative. Marine Policy, 29(1):75–83, January 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0400020X.

Rausser:2009:UCS

[RHKS09] Gordon Rausser, Stephen Hamilton, Marty Kovach, and Ryan Stifter. Unintended consequences: the spillover effects of common property regulations. Marine Policy, 33(1):24–39, January 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000626.

Robinson:2000:RTW

[RM00] Cathy Robinson and David Mercer. Reconciliation in troubled waters? Australian oceans policy and offshore native title rights. Marine Policy, 24(4):349–360, July 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X00000117.

Roberts:2002:CDB

[RM02] Stephen E. Roberts and Peter B. Marlow. Casualties in dry bulk shipping (1963–1996). Marine Policy, 26(6):437–450, November 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000246.

Ryu:2006:LKC

[RNG06] Jeong-Gon Ryu, Jongoh Nam, and John M. Gates. Limitations of the Korean conventional fisheries management regime and expanding Korean TAC system toward output control systems. Marine Policy, 30(5):510–522, September 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000564.

Roach:2004:IEM

[Roa04] J. Ashley Roach. Initiatives to enhance maritime security at sea. Marine Policy, 28(1):41–66, January 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X0300126X.

Rossiter:2003:DSF

[RS03] Tom Rossiter and Selina Stead. Days at sea: from the fishers’ mouths. Marine Policy, 27(3):281–288, May 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X0300006X.

Reid:2003:AFC

[RSJ^{+}03] Christopher Reid, Dale Squires, Yongil Jeon, Len Rodwell, and Raymond Clarke. An analysis of fishing capacity in the western and central Pacific Ocean tuna fishery and management implications. Marine Policy, 27(6):449–469, November 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000654.

Raakjaer:2007:AFM

[RSS^{+}07] Jesper Raakjær, Dao Manh Son, Karl-Johan Stæhr, Holger Hovgård, Nguyen Thi Dieu Thuy, Kristine Ellegaard, Frank Riget, Dang Van Thi, and Phung Giang Hai. Adaptive fisheries management in Vietnam: the use of indicators and the introduction of a multi-disciplinary Marine Fisheries Specialist Team to support implementation. Marine Policy, 31(2):143–152, March 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000522.

Roberts:2008:ILO

[RT08] Julian Roberts and Martin Tsamenyi. International legal options for the control of biofouling on international vessels. Marine Policy, 32(4):559–569, July 2008. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X0700125X.

Roberts:2005:WEP

[RTWJ05] Julian Roberts, Martin Tsamenyi, Tim Workman, and Lindy Johnson. The Western European PSSA proposal: a “politically sensitive sea area”. Marine Policy, 29(5):431–440, September 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000624.

Rauschmayer:2008:ICR

[RWB08] Felix Rauschmayer, Heidi Wittmer, and Augustin Berghöfer. Institutional challenges for resolving conflicts between fisheries and endangered species conservation. Marine Policy, 32(2):178–188, March 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001108.

Rennie:2009:DCM

[RWB09] Hamish G. Rennie, Roger White, and Lars Brabyn. Developing a conceptual model of marine farming in New Zealand. Marine Policy, 33(1):106–117, January 2009. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000900.

Russ:2003:MLM

[RZ03] Garry R. Russ and Dirk C. Zeller. From Mare Liberum to Mare Reservarum. Marine Policy, 27(1):75–78, January 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000544.

Standal:2002:TSC

[SA02] Dag Standal and Bernt Aarset. The tragedy of soft choices: capacity accumulation and lopsided allocation in the Norwegian coastal cod fishery. Marine Policy, 26(3):221–230, May 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000040.

Standal:2008:IRN

[SA08] Dag Standal and Bernt Aarset. The IVQ regime in Norway: a stable alternative to an ITQ regime? Marine Policy, 32(4):

663–668, July 2008. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X07001364.

Sage:2005:IHR

[Sag05] Bénédicte Sage. Identification of ‘High Risk Vessels’ in coastal waters. Marine Policy, 29(4):349–355, July 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000545.

Saharuddin:2001:NOP

[Sah01] Abdul Hamid Saharuddin. National ocean policy — new opportunities for Malaysian ocean development. Marine Policy, 25(6): 427–436, November 2001. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X01000276.

Sumaila:2006:GSE

[SAK06] U. R. Sumaila, J. Alder, and H. Keith. Global scope and economics of illegal fishing. Marine Policy, 30(6):696–703, November 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000795.

Sanchirico:2005:SSP

[San05] James N. Sanchirico. A social scientist’s perspective on the Census of Marine Life. Marine Policy, 29(6):507–520, November 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04001009.

Sissener:2005:CCM

[SB05] Elin H. Sissener and Trond Bjørndal. Climate change and the migratory pattern for Norwegian spring-spawning herring — implications for management. Marine Policy, 29(4):299–309, July 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000399.

Stojanovic:2009:RRN

[SBB^{+}09] T. A. Stojanovic, I. Ball, R. C. Ballinger, G. Lymbery, and W. Dodds. The role of research networks for science-policy collaboration in coastal areas. Marine Policy, 33(6):901–911, November 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000530.

Scholz:2004:PSA

[SBF^{+}04] Astrid Scholz, Kate Bonzon, Rod Fujita, Natasha Benjamin, Nicole Woodling, Peter Black, and Charles Steinback. Participatory socioeconomic analysis: drawing on fishermen’s knowledge for marine protected area planning in California. Marine Policy, 28(4):335–349, July 2004. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X03001118.

Stokke:2004:PIC

[SC04] Olav Schram Stokke and Clare Coffey. Precaution, ICES and the common fisheries policy: a study of regime interplay. Marine Policy, 28(2):117–126, March 2004. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000812.

Schrank:2005:NFT

[Sch05] William E. Schrank. The Newfoundland fishery: ten years after the moratorium. Marine Policy, 29(5):407–420, September 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000600.

Schrank:2007:ACC

[Sch07a] William E. Schrank. The ACIA, climate change and fisheries. Marine Policy, 31(1):5–18, January 2007. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X0600039X.

Schrank:2007:THF

[Sch07b] William E. Schrank. Is there any hope for fisheries management? Marine Policy, 31(3):299–307, May 2007. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000959.

Schumann:2007:CMC

[Sch07c] Sarah Schumann. Co-management and “consciousness”: Fishers’ assimilation of management principles in Chile. Marine Policy, 31(2):101–111, March 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X06000455.

Schug:2008:IIE

[Sch08] Donald M. Schug. The institutional implications of environmental ethics for fishery management in the US exclusive economic zone. Marine Policy, 32(3):514–521, May 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001200.

Scott:2000:AFT

[Sco00] Shirley V. Scott. Australia’s first tuna negotiations with Japan. Marine Policy, 24(4):309–318, July 2000. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X00000075.

Shields:2009:SPA

[SDWF09] Mark A. Shields, Lora Jane Dillon, David K. Woolf, and Alex T. Ford. Strategic priorities for assessing ecological impacts of marine renewable energy devices in the Pentland Firth (Scotland, UK). Marine Policy, 33(4):635–642, July 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001991.

Serdy:2004:OFT

[Ser04] Andrew Serdy. One fin, two fins, red fins, bluefins: some problems of nomenclature and taxonomy affecting legal instruments governing tuna and other highly migratory species. Marine Policy, 28(3):235–247, May 2004. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X03001015.

Salayo:2008:MEC

[SGP^{+}08] Nerissa Salayo, Len Garces, Michael Pido, Kuperan Viswanathan, Robert Pomeroy, Mahfuzuddin Ahmed, Ida Siason, Keang Seng, and Awae Masae. Managing excess capacity in smallscale fisheries: Perspectives from stakeholders in three Southeast Asian countries. Marine Policy, 32(4):692–700, July 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001492.

Sibert:2003:MTT

[SH03] John Sibert and John Hampton. Mobility of tropical tunas and the implications for fisheries management. Marine Policy, 27(1): 87–95, January 2003. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X0200057X.

Sharma:2005:IP

[Sha05] O. P. Sharma. An Indian perspective. Marine Policy, 29(2): 147–151, March 2005. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X0400082X.

StMartin:2008:MLG

[SHA08] Kevin St. Martin and Madeleine Hall-Arber. The missing layer: Geo-technologies, communities, and implications for marine spatial planning. Marine Policy, 32(5):779–786, September 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000663.

Sanchirico:2006:CQB

[SHQF06] James N. Sanchirico, Daniel Holland, Kathryn Quigley, and Mark Fina. Catch-quota balancing in multispecies individual fishing quotas. Marine Policy, 30(6):767–785, November 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000054.

Side:2002:TTI

[SJ02] Jonathan Side and Paul Jowitt. Technologies and their influence on future UK marine resource development and management. Marine Policy, 26(4):231–241, July 2002. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000052.

Sutinen:2003:AMO

[SJ03] Jon G. Sutinen and Robert J. Johnston. Angling management organizations: integrating the recreational sector into fishery management. Marine Policy, 27(6):471–487, November 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000794.

Skaridov:2004:RAR

[Ska04] Alexander Skaridov. The Russian approach to ROE. Marine Policy, 28(1):19–23, January 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X03001234.

Skaridov:2005:NAF

[Ska05] Alexander S. Skaridov. Naval activity in the foreign EEZ — the role of terminology in law regime. Marine Policy, 29(2):153–155, March 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000831.

Skjaerseth:2006:PNE

[Skj06] Jon Birger Skjærseth. Protecting the North-East Atlantic: enhancing synergies by institutional interplay. Marine Policy, 30 (2):157–166, March 2006. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X04001046.

Stoneham:2005:RRR

[SLCS05] Gary Stoneham, Nicola Lansdell, Anne Cole, and Loris Strappazzon. Reforming resource rent policy: an information economics perspective. Marine Policy, 29(4):331–338, July 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000521.

Sheriff:2008:APC

[SLT08] Natasja Sheriff, David C. Little, and Kwanta Tantikamton. Aquaculture and the poor — Is the culture of high-value fish a viable livelihood option for the poor? Marine Policy, 32(6):1094– 1102, November 2008. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08000535.

Satria:2004:DFM

[SM04] Arif Satria and Yoshiaki Matsuda. Decentralization of fisheries management in Indonesia. Marine Policy, 28(5):437–450, September 2004. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X03001362.

Storrier:2006:DMC

[SM06] K. L. Storrier and D. J. McGlashan. Development and management of a coastal litter campaign: the voluntary coastal partnership approach. Marine Policy, 30(2):189–196, March 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000102.

Schumann:2007:SCF

[SM07] Sarah Schumann and Seth Macinko. Subsistence in coastal fisheries policy: What’s in a word? Marine Policy, 31(6):706– 718, November 2007. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X07000176.

Singh:2008:EPM

[SM08] Asha Singh and Laurence Mee. Examination of policies and MEAs commitment by SIDS for sustainable management of the Caribbean Sea. Marine Policy, 32(3):274–282, May 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000723.

Smith:2003:EIO

[Smi03] Hance D. Smith. Emerging issues in oceans, coasts and islands. Marine Policy, 27(4):289–290, July 2003. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000484.

Satria:2006:CST

[SMS06] Arif Satria, Yoshiaki Matsuda, and Masaaki Sano. Contractual solution to the tragedy of property right in coastal fisheries. Marine Policy, 30(3):226–236, May 2006. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000126.

Somers:2001:ASM

[Som01] Eddy Somers. The arrest of ships in maritime zones beyond internal waters in Belgian maritime law. Marine Policy, 25(1): 61–69, January 2001. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X00000361.

Soma:2003:HIS

[Som03] Katrine Soma. How to involve stakeholders in fisheries management — a country case study in Trinidad and Tobago. Marine Policy, 27(1):47–58, January 2003. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000507.

Song:2005:CSI

[Son05] Yann-Huei Song. Cross–strait interactions on the South China Sea issues: a need for CBMs. Marine Policy, 29(3):265–280, May 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000375.

Sowman:2006:SSS

[Sow06] Merle Sowman. Subsistence and small-scale fisheries in South Africa: a ten-year review. Marine Policy, 30(1):60–73, January 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000515.

Smith:2001:SIO

[SP01] Hance D. Smith and Jonathan Potts. Special issue on ocean governance and capacity building. Marine Policy, 25(6):389– 390, November 2001. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X01000318.

Suris-Regueiro:2005:RCE

[SR05] J. C. Suris-Regueiro. Reply to the comments by: Erik Lindebo, Hans Frost and Jørgen Løkkergaard on: Surís-Regueiro, Varela-Lafuente and Iglesias-Malvido, “Effectiveness of the structural fisheries policy in the European Union”. Marine Policy 2003; 27(6), 535–544. Marine Policy, 29(3):285–287, May 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0400048X.

Schafer:2008:AFA

[SR08] Adalberto Gularte Schafer and Enir Girondi Reis. Artisanal fishing areas and traditional ecological knowledge: the case study of the artisanal fisheries of the Patos Lagoon estuary (Brazil). Marine Policy, 32(3):283–292, May 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000735.

Singh-Renton:2003:SCC

[SRMM03] Susan Singh-Renton, Robin Mahon, and Patrick McConney. Small Caribbean (CARICOM) states get involved in management of shared large pelagic species. Marine Policy, 27(1):39–46, January 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000490.

Suris-Regueiro:2002:PFF

[SRVLGG02] Juan C. Surís-Regueiro, Manuel M. Varela-Lafuente, and M. Dolores Garza-Gil. Profitability of the fishing fleet and structural aid in the European Union. Marine Policy, 26(2):107–119, March 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X01000422.

Suris-Regueiro:2003:ESF

[SRVLIM03] Juan C. Surís-Regueiro, Manuel M. Varela-Lafuente, and Carlos Iglesias-Malvido. Effectiveness of the structural fisheries policy in the European Union. Marine Policy, 27(6):535–544, November 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000800.

Strain:2006:MAS

[SRW06] Lisa Strain, Abbas Rajabifard, and Ian Williamson. Marine administration and spatial data infrastructure. Marine Policy, 30 (4):431–441, July 2006. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X05000345.

Stenevik:2007:ICC

[SS07] Erling Kåre Stenevik and Svein Sundby. Impacts of climate change on commercial fish stocks in Norwegian waters. Marine Policy, 31(1):19–31, January 2007. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000406.

Standal:2003:FLF

[Sta03] Dag Standal. Fishing the last frontier — controversies in the regulations of shrimp trawling in the high Arctic. Marine Policy, 27(5):375–388, September 2003. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000460.

Standal:2005:NBF

[Sta05] Dag Standal. Nuts and bolts in fisheries management — a technological approach to sustainable fisheries? Marine Policy, 29 (3):255–263, May 2005. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X04000363.

Standal:2006:RDB

[Sta06] Dag Standal. The rise and decline of blue whiting fisheries — capacity expansion and future regulations. Marine Policy, 30 (4):315–327, July 2006. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X0500031X.

Standal:2007:ICF

[Sta07] Dag Standal. Institutional changes and fleet structure: Towards the final solution? Marine Policy, 31(2):94–100, March 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000364.

Standal:2008:RFF

[Sta08] Dag Standal. The rise and fall of factory trawlers: an eclectic approach. Marine Policy, 32(3):326–332, May 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000784.

Stokke:2007:LRA

[Sto07] Olav Schram Stokke. A legal regime for the Arctic?: Interplay with the Law of the Sea Convention. Marine Policy, 31(4):402– 408, July 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06001035.

Stokke:2009:TMC

[Sto09] Olav Schram Stokke. Trade measures and the combat of IUU fishing: Institutional interplay and effective governance in the Northeast Atlantic. Marine Policy, 33(2):339–349, March 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001309.

Standal:2007:CCF

[SU07] Dag Standal and Ingrid Bouwer Utne. Can cod farming affect cod fishing? A system evaluation of sustainability. Marine Policy, 31 (4):527–534, July 2007. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X06001357.

Subade:2007:MCE

[Sub07] Rodelio F. Subade. Mechanisms to capture economic values of marine biodiversity: the case of Tubbataha Reefs UNESCO World Heritage Site, Philippines. Marine Policy, 31(2):135–142, March 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000509.

Sumaila:2007:PEB

[SVL07] Ussif Rashid Sumaila, John Volpe, and Yajie Liu. Potential economic benefits from sablefish farming in British Columbia. Marine Policy, 31(2):81–84, March 2007. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000236.

Stewart:2008:CCS

[SW08] James Stewart and Kim Walshe. Compliance costs and the small fisher: a study of exiters from the New Zealand fishery. Marine Policy, 32(1):120–131, January 2008. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000589.

Steinback:2009:SSF

[SWC09] Scott Steinback, Kristy Wallmo, and Patricia Clay. Saltwater sport fishing for food or income in the Northeastern US: Statistical estimates and policy implications. Marine Policy, 33(1): 49–57, January 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08000754.

Stewart:2006:DSF

[SWM06] James Stewart, Kim Walshe, and Beverley Moodie. The demise of the small fisher? A profile of exiters from the New Zealand fishery. Marine Policy, 30(4):328–340, July 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000321.

Sydnes:2001:NRF

[Syd01] Are K. Sydnes. New regional fisheries management regimes: establishing the South East Atlantic fisheries organisation. Marine Policy, 25(5):353–364, September 2001. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X01000215.

Sydnes:2002:RFO

[Syd02] Are K. Sydnes. Regional fishery organisations in developing regions: adapting to changes in international fisheries law. Marine Policy, 26(5):373–381, September 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000192.

Taussik:2007:OSP

[Tau07] Jane Taussik. The opportunities of spatial planning for integrated coastal management. Marine Policy, 31(5):611–618, September 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000243.

Thomson:2009:CBC

[TG09] Kaleekal Thomson and Tim Gray. From community-based to co-management: Improvement or deterioration in fisheries governance in the Cherai Poyil fishery in the Cochin Estuary, Kerala, India? Marine Policy, 33(4):537–543, July 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001899.

Thompson:2008:FSB

[Tho08] Myles H. Thompson. Fostering sustainable behaviours in community-based co-managed fisheries. Marine Policy, 32(3): 413–420, May 2008. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X07000991.

Tiller:2008:NSD

[Til08] Rachel Gjelsvik Tiller. The Norwegian system and the distribution of claims to redfeed. Marine Policy, 32(6):928–940, November 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000237.

Tzannatos:2009:AAG

[TK09] Ernestos Tzannatos and Dimitris Kokotos. Analysis of accidents in Greek shipping during the pre- and post-ISM period. Marine Policy, 33(4):679–684, July 2009. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X09000074.

Thatje:2007:UNN

[TLHN07] Sven Thatje, Jürgen Laudien, Olaf Heilmayer, and Cornelia E. Nauen. Understanding El Niño — The importance of Grey Literature in Coastal Ecosystem Research and Management. Marine Policy, 31(2):85–93, March 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000261.

Townsend:2006:DPI

[TMY06a] Ralph E. Townsend, James McColl, and Michael D. Young. Design principles for individual transferable quotas. Marine Policy, 30(2):131–141, March 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X04001022.

Trondsen:2006:TMO

[TMY06b] Torbjorn Trondsen, Thorolfur Matthiasson, and James A. Young. Towards a market-oriented management model for straddling fish stocks. Marine Policy, 30(3):199–206, May 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000242.

Townend:2002:MSS

[Tow02] Ian Townend. Marine science for strategic planning and management: the requirement for estuaries. Marine Policy, 26(3): 209–219, May 2002. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X02000039.

Tillin:2008:ACB

[TRF08] H. M. Tillin, S. I. Rogers, and C. L. J. Frid. Approaches to classifying benthic habitat quality. Marine Policy, 32(3):455–464, May 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001030.

Trondsen:2004:TMO

[Tro04] Torbjørn Trondsen. Toward market orientation: the role of auctioning individual seasonal quotas (ISQ). Marine Policy, 28(5):375–382, September 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X03001180.

Thorpe:2005:WFI

[TRvAB05] Andy Thorpe, Chris Reid, Raymon van Anrooy, and Cecile Brugere. When fisheries influence national policy-making: an analysis of the national development strategies of major fishproducing nations in the developing world. Marine Policy, 29 (3):211–222, May 2005. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X04000314.

Tsai:2005:PRA

[TS05] Ming-Chih Tsai and Chin-Hui Su. Political risk assessment of five East Asian ports — the viewpoints of global carriers. Marine Policy, 29(4):291–298, July 2005. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000387.

Teh:2007:MOP

[TS07] Louise Teh and U. Rashid Sumaila. Malthusian overfishing in Pulau Banggi? Marine Policy, 31(4):451–457, July 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07000061.

Tsuru:2004:RPA

[Tsu04] Yasuko Tsuru. Rethinking the principle of abstention: the North Pacific and beyond. Marine Policy, 28(6):541–552, November 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000120.

Teh:2009:OSE

[TTSS09] Lydia C. L. Teh, Louise S. L. Teh, Ben Starkhouse, and U. Rashid Sumaila. An overview of socio-economic and ecological perspectives of Fiji’s inshore reef fisheries. Marine Policy, 33(5):807– 817, September 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X09000323.

Thorpe:2009:FFS

[TWN^{+}09] Andy Thorpe, David Whitmarsh, Ernest Ndomahina, Andrew Baio, Miatta Kemokai, and Thomas Lebbie. Fisheries and failing states: the case of Sierra Leone. Marine Policy, 33(2):393–400, March 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001450.

Tzannatos:2005:TRG

[Tza05] Ernestos S. Tzannatos. Technical reliability of the Greek coastal passenger fleet. Marine Policy, 29(1):85–92, January 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000296.

Uggla:2007:EPF

[Ugg07] Ylva Uggla. Environmental protection and the freedom of the high seas: the Baltic Sea as a PSSA from a Swedish perspective. Marine Policy, 31(3):251–257, May 2007. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000819.

Utne:2006:SEP

[Utn06] Ingrid Bouwer Utne. Systems engineering principles in fisheries management. Marine Policy, 30(6):624–634, November 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000248.

Utne:2007:SES

[Utn07] Ingrid Bouwer Utne. System evaluation of sustainability in the Norwegian cod-fisheries. Marine Policy, 31(4):390–401, July 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06001084.

Utne:2008:ASF

[Utn08a] Ingrid Bouwer Utne. Acceptable sustainability in the fishing fleet. Marine Policy, 32(3):475–482, May 2008. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001054.

Utne:2008:SFV

[Utn08b] Ingrid Bouwer Utne. Are the smallest fishing vessels the most sustainable? — trade-off analysis of sustainability attributes. Marine Policy, 32(3):465–474, May 2008. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001042.

Valencia:2004:FA

[VA04] Mark J. Valencia and Kazumine Akimoto. Foreword and Acknowledgments. Marine Policy, 28(1):1, January 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X03001192.

Valencia:2005:FA

[VA05a] Mark J. Valencia and Kazumine Akimoto. Foreword and acknowledgements. Marine Policy, 29(2):95, March 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000752.

Valencia:2005:RTM

[VA05b] Mark J. Valencia and Kazumine Akimoto. Report of the Tokyo meeting and progress to date. Marine Policy, 29(2):101–106, March 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000776.

Valencia:2006:GNO

[VA06] Mark J. Valencia and Kazumine Akimoto. Guidelines for navigation and overflight in the exclusive economic zone. Marine Policy, 30(6):704–711, November 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X05000801.

Valencia:2000:CPE

[Val00] Mark J. Valencia. Comment on “Prospects for effective marine governance in the NW Pacific region”. Marine Policy, 24(5):439– 440, September 2000. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X00000191.

Vallega:2001:OGP

[Val01] Adalberto Vallega. Ocean governance in post-modern society — a geographical perspective. Marine Policy, 25(6):399–414, November 2001. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X01000240.

Valencia:2003:VMB

[Val03] M. J. Valencia. Validity of Malaysia’s baselines and territorial sea claim in the northern Malacca Strait. Marine Policy, 27(5):367– 373, September 2003. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X03000472.

Valencia:2004:CRB

[Val04a] Mark J. Valencia. Conclusions, regime building and the way forward. Marine Policy, 28(1):89–96, January 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X03001295.

Valencia:2004:I

[Val04b] Mark J. Valencia. Introduction. Marine Policy, 28(1):3–5, January 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03001209.

Valencia:2004:SBD

[Val04c] Mark J. Valencia. Summary of the Bali dialogue. Marine Policy, 28(1):7–12, January 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X03001210.

Valencia:2005:IMI

[Val05a] M. J. Valencia. Introduction: military and intelligence gathering activities in the exclusive economic zones: consensus and disagreement II. Marine Policy, 29(2):97–99, March 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000764.

Valencia:2005:CWF

[Val05b] Mark J. Valencia. Conclusions and the way forward. Marine Policy, 29(2):185–187, March 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X04000879.

Valcic:2009:SPB

[Val09] Branka Valcic. Spatial policy and the behavior of fishermen. Marine Policy, 33(2):215–222, March 2009. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X08001024.

VanDyke:2004:MSP

[Van04] Jon M. Van Dyke. Military ships and planes operating in the exclusive economic zone of another country. Marine Policy, 28 (1):29–39, January 2004. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X03001258.

VanDyke:2005:DRN

[Van05] Jon M. Van Dyke. The disappearing right to navigational freedom in the exclusive economic zone. Marine Policy, 29(2):107– 121, March 2005. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X04000788.

Varjopuro:2008:AMG

[Var08] Riku Varjopuro. Addressing multiple goals of ecosystem approach through deliberation and technology development. Marine Policy, 32(2):212–222, March 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X07001121.

vanderBurg:2000:NCE

[vdB00] Tsjalle van der Burg. Neo-classical economics, institutional economics and improved fisheries management. Marine Policy, 24 (1):45–51, January 2000. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X99000081.

vandenBerg:2007:WSK

[vdBL07] Marrit van den Berg and Bart Jan Lensing. Work sharing in Kerala’s fisheries. Marine Policy, 31(4):535–539, July 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06001345.

vandenBergh:2002:EHA

[vdBND^{+}02] Jeroen C. J. M. van den Bergh, Paulo A. L. D. Nunes, Harm M. Dotinga, Wiebe H. C. F. Kooistra, Engel G. Vrieling, and Louis Peperzak. Exotic harmful algae in marine ecosystems: an integrated biological–economic–legal analysis of impacts and policies. Marine Policy, 26(1):59–74, January 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0100032X.

Veenstra:2002:NEC

[Vee02] Albert W. Veenstra. Nautical education in a changing world: the case of the Netherlands. Marine Policy, 26(2):133–141, March 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0100046X.

Vetemaa:2002:EFS

[VEH02] Markus Vetemaa, Margit Eero, and Rögnvaldur Hannesson. The Estonian fisheries: from the Soviet system to ITQs and quota auctions. Marine Policy, 26(2):95–102, March 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X01000409.

Verlaan:2007:EAI

[Ver07] Philomène A. Verlaan. Experimental activities that intentionally perturb the marine environment: Implications for the marine environmental protection and marine scientific research provisions of the 1982 United Nations Convention on the Law of the Sea. Marine Policy, 31(2):210–216, March 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000790.

Vetemaa:2006:TSS

[VES06] Markus Vetemaa, Redik Eschbaum, and Toomas Saat. The transition from the Soviet system to a market economy as a cause of instability in the Estonian coastal fisheries sector. Marine Policy, 30(6):635–640, November 2006. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000680.

Viladrich-Grau:2003:MPP

[VG03] Montserrat Viladrich-Grau. Monitoring policies to prevent oil spills: lessons from the theoretical literature. Marine Policy, 27 (3):249–263, May 2003. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X03000046.

Varjopuro:2008:IIB

[VGH^{+}08] Riku Varjopuro, Tim Gray, Jenny Hatchard, Felix Rauschmayer, and Heidi Wittmer. Introduction: Interaction between environment and fisheries — The role of stakeholder participation. Marine Policy, 32(2):147–157, March 2008. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001078.

Vince:2009:NZO

[VH09] Joanna Vince and Marcus Haward. New Zealand oceans governance: Calming turbulent waters? Marine Policy, 33(2): 412–418, March 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08001474.

vanHoof:2009:EMP

[vHvT09] L. van Hoof and J. van Tatenhove. EU marine policy on the move: the tension between fisheries and maritime policy. Marine Policy, 33(4):726–732, July 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X09000256.

Vince:2006:SER

[Vin06] Joanna Vince. The South East Regional Marine Plan: Implementing, Australia’s Oceans Policy. Marine Policy, 30(4):420– 430, July 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000333.

Valencia:2002:SKR

[VL02] Mark J. Valencia and Yong Hee Lee. The South Korea–Russia– Japan fisheries imbroglio. Marine Policy, 26(5):337–343, September 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000155.

Veitayaki:2001:CBM

[VS01] Joeli Veitayaki and G. Robin South. Capacity building in the marine sector in the Pacific Islands: the role of the University of the South Pacific’s Marine Studies Programme. Marine Policy, 25(6):437–444, November 2001. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X01000288.

vanSittert:2002:WCR

[vS02] Lance van Sittert. “Those who cannot remember the past are condemned to repeat it”: comparing fisheries reforms in South Africa. Marine Policy, 26(4):295–305, July 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X0200012X.

vanSittert:2006:BFD

[vSBHS06] L. van Sittert, G. Branch, M. Hauck, and M. Sowman. Benchmarking the first decade of post-apartheid fisheries reform in South Africa. Marine Policy, 30(1):96–110, January 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0500045X.

vanSittert:2006:IPA

[vSH06] L. van Sittert and M. Hauck. Introduction: Post-apartheid marine fisheries in South Africa: Through the ten-year looking glass. Marine Policy, 30(1):1–2, January 2006. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000485.

VanDyke:2003:NSK

[VVG03] Jon M. Van Dyke, Mark J. Valencia, and Jenny Miller Garmendia. The North/South Korea Boundary Dispute in the Yellow (West) Sea. Marine Policy, 27(2):143–158, March 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0200088X.

vanZwieten:2002:IUF

[vZvDT02] Paul A. M. van Zwieten, Wim L. T. van Densen, and Dang Van Thi. Improving the usage of fisheries statistics in Vietnam for production planning, fisheries management and nature conservation. Marine Policy, 26(1):13–34, January 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X01000367.

Watson:2005:CSN

[WAKP05] Reg Watson, Jackie Alder, Adrian Kitchingman, and Daniel Pauly. Catching some needed attention. Marine Policy, 29(3): 281–284, May 2005. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X04000557.

Wilson:2006:CSL

[WASK06] Douglas Clyde Wilson, Mahfuzuddin Ahmed, Susanna V. Siar, and Usha Kanagaratnam. Cross-scale linkages and adaptive management: Fisheries co-management in Asia. Marine Policy, 30(5):523–533, September 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X05000576.

Wiber:2004:PRS

[WBCK04] Melanie Wiber, Fikret Berkes, Anthony Charles, and John Kearney. Participatory research supporting community-based fishery management. Marine Policy, 28(6):459–468, November 2004. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03001374.

Waugh:2008:CPR

[WBGC08] S. M. Waugh, G. B. Baker, R. Gales, and J. P. Croxall. CCAMLR process of risk assessment to minimise the effects of longline fishing mortality on seabirds. Marine Policy, 32(3):442–454, May 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001029.

Whitmarsh:2006:SSA

[WCB06] David J. Whitmarsh, Elizabeth J. Cook, and Kenneth D. Black. Searching for sustainability in aquaculture: an investigation into the economic prospects for an integrated salmon– mussel production system. Marine Policy, 30(3):293–298, May 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000114.

Wiber:2009:ECE

[WCKB09] Melanie Wiber, Anthony Charles, John Kearney, and Fikret Berkes. Enhancing community empowerment through participatory fisheries research. Marine Policy, 33(1):172–179, January 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X08000973.

Wagenaar:2007:DSS

[WD07] Arnout Wagenaar and Marijke D’Haese. Development of smallscale fisheries in Yemen: an exploration. Marine Policy, 31(3): 266–275, May 2007. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X06000832.

Wonham:2000:MTI

[WDAT00] J. Wonham, C. M. Davies, V. G. Asimakopoulos, and B. S. Tselentis. Marine transportation of irradiated nuclear fuel, plutonium and radioactive wastes: the continuing debate on regulatory measures. Marine Policy, 24(4):287–299, July 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X00000051.

Wind:2002:IMG

[WdK02] H. G. Wind and J.-L. de Kok. Interdisciplinary modeling: gaps and bridges. Marine Policy, 26(4):243–251, July 2002. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000064.

West:2003:ISA

[Wes03] Mary Beth West. Improving science applications to coastal management. Marine Policy, 27(4):291–293, July 2003. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000381.

Wang:2001:FSA

[WF01] J. Wang and P. Foinikis. Formal safety assessment of containerships. Marine Policy, 25(2):143–157, March 2001. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X01000057.

Witbooi:2006:LFR

[Wit06] Emma Witbooi. Law and fisheries reform: Legislative and policy developments in South African fisheries over the decade 1994–2004. Marine Policy, 30(1):30–42, January 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X05000424.

Witbooi:2008:ISB

[Wit08] Emma Witbooi. The infusion of sustainability into bilateral fisheries agreements with developing countries: the European Union example. Marine Policy, 32(4):669–679, July 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001376.

Whitmarsh:2000:PMC

[WJPN00] David Whitmarsh, Carl James, Helen Pickering, and Arthur Neiland. The profitability of marine commercial fisheries: a review of economic information needs with particular reference to the UK. Marine Policy, 24(3):257–263, May 2000. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X00000026.

Ward:2009:MMS

[WK09] John M. Ward and Michael Kelly. Measuring management success: Experience with United States fisheries. Marine Policy, 33(1):164–171, January 2009. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X08000961.

Ward:2000:SAR

[WKT00] Peter Ward, Bob Kearney, and Nektarios Tsirbas. Science arrangements for the regional management of tuna fisheries. Marine Policy, 24(2):93–108, March 2000. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X99000160.

Windle:2008:FOH

[WNB^{+}08] M. J. S. Windle, B. Neis, S. Bornstein, M. Binkley, and P. Navarro. Fishing occupational health and safety: a comparison of regulatory regimes and safety outcomes in six countries. Marine Policy, 32(4):701–710, July 2008. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X07001509.

Whitmarsh:2009:SAM

[WP09] David Whitmarsh and Maria Giovanna Palmieri. Social acceptability of marine aquaculture: the use of survey-based methods for eliciting public and stakeholder preferences. Marine Policy, 33(3):452–457, May 2009. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X08001504.

Whitmarsh:2003:ESA

[WPBD03] David Whitmarsh, Carlo Pipitone, Fabio Badalamenti, and Giovanni D’Anna. The economic sustainability of artisanal fisheries: the case of the trawl ban in the Gulf of Castellammare, NW Sicily. Marine Policy, 27(6):489–497, November 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X03000629.

Whitfield:2007:UPW

[WR07] Rebecca Whitfield and Ronan Roche. UK Personal Watercraft Management: a user perspective. Marine Policy, 31(4):564–572, July 2007. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06001072.

Wilson:2006:LEK

[WRD06] Douglas Clyde Wilson, Jesper Raakjær, and Poul Degnbol. Local ecological knowledge and practical fisheries management in the tropics: a policy brief. Marine Policy, 30(6):794–801, November 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X06000078.

Wu:2005:CSS

[WW05] Bin Wu and Nik Winchester. Crew study of seafarers: a methodological approach to the global labour market for seafarers. Marine Policy, 29(4):323–330, July 2005. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X0400051X.

Wallace:2006:RMI

[WW06] Jude Wallace and Ian Williamson. Registration of marine interests in Asia–Pacific region. Marine Policy, 30(3):207–219, May 2006. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X0500014X.

Xizhong:2004:CPO

[Xiz04] Cheng Xizhong. A Chinese perspective on ‘Operational Modalities’. Marine Policy, 28(1):25–27, January 2004. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X03001246.

Xue:2006:CDW

[Xue06] Guifang (Julia) Xue. China’s distant water fisheries and its response to flag state responsibilities. Marine Policy, 30(6):651– 658, November 2006. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X05000746.

Xiaofeng:2005:CP

[XX05] Ren Xiaofeng and Senior Colonel Cheng Xizhong. A Chinese Perspective. Marine Policy, 29(2):139–146, March 2005. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X04000818.

Yandle:2003:CBS

[Yan03] Tracy Yandle. The challenge of building successful stakeholder organizations: New Zealand’s experience in developing a fisheries co-management regime. Marine Policy, 27(2):179–192, March 2003. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X02000714.

Yandle:2008:PPB

[Yan08] Tracy Yandle. The promise and perils of building a comanagement regime: an institutional assessment of New Zealand fisheries management between 1999 and 2005. Marine Policy, 32(1):132–141, January 2008. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X07000590.

Yu:2006:NIA

[YM06] Yunjun Yu and Yongtong Mu. The new institutional arrangements for fisheries management in Beibu Gulf. Marine Policy, 30(3):249–260, May 2006. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X05000084.

Yu:2008:FCM

[YY08] Huiguo Yu and Yunjun Yu. Fishing capacity management in China: Theoretic and practical perspectives. Marine Policy, 32 (3):351–359, May 2008. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X07000917.

Zeller:2005:LRF

[ZFP05] Dirk Zeller, Rainer Froese, and Daniel Pauly. On losing and recovering fisheries and marine science data. Marine Policy, 29 (1):69–73, January 2005. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X04000193.

Zhang:2009:PTA

[ZYN09] Dao Hua Zhang, Tsz Leung Yip, and Chiu-On Ng. Predicting tsunami arrivals: Estimates and policy implications. Marine Policy, 33(4):643–650, July 2009. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0308597X08002005.